A Complete Bibliography of Publications in Concurrency: Practice and Experience

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

14 October 2017 Version 1.19

Title word cross-reference

3 [BDOP95, Fah96]. ^{T} [Cun97]. ^{T M} [IKY^{+}00]. _{H}P F [OP99]. E^{3} [CLMP95].

1980s [Mes91]. 1999 [Gue00].

2 [BFFL97, BR89, GS97a, Hel93, LF91]. 2-D [LF91]. 2D [CYL95, DT98].

3 [AAFG94, RPB91]. 3-D [RPB91]. 365 [Han00, Leó00]. 3D [LCLL97, WB95b].

510 [ES92, WCB91].

6000 [BRSR93]. 64-bit [KG97].

77 [SF93].

860 [BH92, HHGW95].

= [Cun97].

abstract [NHZ96]. abstractions [KAS98]. Accelerated [GH95]. accelerator [HBKR96]. access [CDD97, HAFY00, PH00a]. Accurate [JDZW99]. Achievements [Fox91a]. ACM [Fox00a, Fox00b, Fox00c]. across [GS98, vL99]. Action [MK93, DD94]. Ada [BT93]. adaptation [SB99]. Adaptive [BSS91, HM98, KB98, Lap92, RGSB94, CYL95, CD99, CH99, DW97, HBG00, LZM97, LL95, LL99, WB95a, SBS93]. address [BDL00, KSN97]. Advanced [MBF^{+}90]. advection [Ber94]. advection-diffusion [Ber94]. advisory [CQ95]. aerial [ND94]. Aeroelastic [SW91].

Aerospace [HWPG92]. Age [Fox89]. Agent [BMR^{+}00, XW00]. agents [KB98]. Aircraft [TV90, KWB^{+}99]. Ajents [ICB00]. Algebra [Dem91, CDW96, CWB94, WKE98]. Algebraic [SS98, MPR95]. Algorithm [ALV^{+}91, AH91, CSTG92, KK90, SBBG89, ST92, TV90, Whe90, AD95, AMVA95, CHB98, CE95, Cho98, CYL95, CB98, DV94, HBE98, JT96, KSK00, LRN97, LSF97, LL95, LL99, MRR^{+}96, WB95b, Wu95, vdGW97]. Algorithms [BC91, BSS91, BF92b, GRS^{+}97, LF91, MF92, MF93, Wil91, AGMM94, AW97, CH99, DM98, GM00, GHSJ98, JPK98, LZM96, MR96, MF94, MW98, MP96, MGG00, NA96, SB99, XMLL95, CDW94]. Allocating [MF92, MF94]. allocation [KSN97, MMN95, NR94]. alternating [Lin97]. AMT [CWB94, WCB91]. Analysis [Hel93, JR92, KMT93, MT91, SSB92, SF93, TDD93, AFH98, ARS00, BGZ97, Cla97, DV94, FSDC99, GLV96, JT96, LP00, LF98, PH00b, SS94, Tog97, Tog98, VNL98, ZP99]. analytical [WDT^{+}99]. animation [LCLL97]. Annealing [KK90, PdOG99]. Annotating [HAKN97, JK00]. annotation [ANH00]. annotation-aware [ANH00]. announce [CRS98]. announce-listen [CRS98]. anti [LLL98]. anti-tasks [LLL98]. applets [BKKK98, GPO98]. Application [GWP93, RCZM91, SCL97, VBFW91, ABC^{+}00, CWB94, CPM00, DM96, RR96, WRMB98, HWPG92]. Application-level [SCL97]. Applications [BKTJ92, BGL92, Rin98, SW91, SSB92, vR89, Bak99b, BGZ97, Ben00, BCB00, CFLH99, DAB97, EAK00, ICB00, Kar99, Kea99, KSAC95, LK96, MLF97, XLD97, Bak99a]. Apply [WDW91]. approach [Dav97, RG00, TBBK94]. Approaches [ALV^{+}91, DCGC92, CC97a, PS00b]. Approximate [Lap92, PRA00]. Architecture [MBF^{+}90, AB98, ARS00, CvE00, EE95, JH00, JCS98, LZM96, WLR^{+}00]. Architectures [GOJ89, MPS93, PDH91, VBFW91, ACF99, BN00, KSAC95, Wad99, YTPS95]. area [BFFL97, vNMB^{+}00]. Argus [Bal92]. Ariadne [MR98b]. ARMI [RWB97]. Array [ES92, WDW91, Fid97, LP00]. Arrays [Woo91, Zom92, ZM92]. ARS [CSS98]. Aspects [Bak99c, GGS00]. Assembled [SBS93]. Assignment [DCGC92, AD95]. Asynchronous [RWB97]. atmospheric [KSS^{+}96, LF98]. Atomic [MK93]. authentication [Sch99]. Author [Ano95, Ano96, Ano97, Ano98, Leó00]. Automata [Han93b]. automatable [MJC99]. automated [Wei99]. Automatic [SS97, CCCL97, CP98, vRvGS97]. Automatically [BG97a]. Autopilot [TV90]. Averaging [MRCH^{+}93]. avoiding [HBG00]. aware [ARS97, ANH00].

BACI [BCB00]. backpropagation [AB98]. balance [RG00, AH91]. balanced [BAGGB99, EE95]. Balancing [Boi90, Wil91, AGMM94, CYL95, HBE98, MO95, MGG00, Nak99, RR97, SS97, Tär94, Tog98, VO98b, WB95a, Wu97, XMLL95]. bandwidth [WRMB98]. Based [FFH^{+}93, Ada91, AFH98, AISS97, AW97, AR94, Bak00, BBC^{+}97, BMR^{+}00, CMM^{+}97, CF96, CCI^{+}97, CSS98, CPM00, Cun97, DM96, DF97, DAB97, DTY91, GS92, GS98, HAFY00, HG99, JA98, KABG^{+}91, Lin97, LCLL97, PSS97, RM91, SRW^{+}00, SHGD98, TBBK94, VO98a, WJC93, Wu97, XW00, YC98]. Basic [BH92, CDW96]. Bayanihan [SHW98]. BBN [SMC92]. beam [VSKN96]. behaviour [CP95]. Benchmark [AAB^{+}93, AH91, MR98a, BSW^{+}00, Cun97]. Benchmarking [BR89, DBRF96, MBF^{+}90]. Benchmarks [AAB^{+}93, HHGW95]. benefits [BRTK94]. best [MJC99]. between [AW97, Kar98, Lil95]. Beyond [ALCG95, Phi00a]. binary [PH00a]. Biological [GWP93]. bisection [BV97, BS94a]. bit [KSK00, KG97]. bit-rate [KSK00]. ‘Blade [CFMP91]. BLAS [AAFG94, BG97b, CDW96, CGM^{+}97]. block [CDW96, DD94, GHvdGV97, GHSJ98, HWPG92, WO96]. block-cyclic [WO96]. Boltzmann [BRSR93]. Bound [SBBG89]. boundaries [GS98]. bounds [CDMW94]. Branch [SBBG89]. Branch-and-Bound [SBBG89]. Bridging [Kar98, RCZM91]. browser [CS97]. BSP [DHS99, RPL96]. buffer [CHB98]. Building [JJ96, KAS98, SHW98]. bursty [LL99]. bytecode [BG98, Cla97]. bytecodes [CL97b, GTV98, HAKN97].

C [Han00, Leó00, Ada91, BB00, Dem98, KPO00]. C/C [KPO00]. CACAO [KG97]. caches [BRTK94]. Caching [PFS^{+}93]. CAD [KSAC95]. calculate [LP00]. Calculations [Wil91, VSKN96]. calculus [MJ96]. Caltech [Fox89]. capabilities [ZK99]. CAPE [NHMW91]. care [BCF^{+}97]. Carlo [MR96, SK00, VSKN96]. case [AGD97, Bri00, GFS99, HHKP97, UTS00]. CCp [BB00]. Cedar [EHJ^{+}93]. cell [RIST00, SMC92, Wal90]. Cellular [Han93b]. CEV [RBF98]. chain [MR96]. chains [PS96]. Challenges [HV93, Ahm95b]. Character [JMM90]. characteristics [BN00]. Characterizing [Wal90]. Cholesky [CH99, DD00, GS97a]. circulation [LF98]. class [CHMZ94, JK00]. classes [SS98]. Claus [BA98]. client [Kar99]. client-server [Kar99]. Clock [Dun92, DV94, JJ96]. closely [ACF99]. Cluster [BC91, BRSR93, DTY91, KB98, SS94]. clustering [EKR97, SCM95]. clusters [AM00b, BGKR99, Phi00a, Tog97, Tog98, ZHS99]. CM [DBRF96, Hel93]. CM-2 [Hel93]. CM-5 [DBRF96]. co [Kea99, OP99, PD99, vL99]. co-operating [vL99]. co-ordination [Kea99, OP99, PD99]. coarse [BBC^{+}97, HHKP97]. Code [SMC92, Wal90, HM98, LF98, SSS98, SK00]. codes [HASnP00]. coding [KSK00]. collaboration [CC97b, DAB97]. Collaborative [KK97, BCF^{+}97, LCLL97, RBF98]. Collective [BI98]. collectors [Phi00a]. coloring [GM00]. COLT} [OP99]. Combat [WHF^{+}89]. Comes [Fox89]. commercial [Bak99b]. COMMS1 [HHGW95]. Communicating [Ada91, JL95]. Communicating-sequential-processes-based [Ada91]. Communication [CDL^{+}93, HPPF94, HR92, YC98, ACF99, BI98, BS94b, CDR99, DBRF96, DHS99, EAK00, FTT97, GHSJ98, Hin97, PS00b, WC00]. communication/computation [CDR99]. Communications [DHN93, ZK99]. Comparative [SCM95, ZS93, VNL98]. Comparison [AM91, CWB94, PFS^{+}93, CMF97, hSM97]. compendium [MMN95]. Competitive [CPV93]. Compilation [SBW91, CCCL97, vRvGS97]. Compile [CS95, SS97]. Compile-time [CS95, SS97]. compiled [NM00]. Compiler [JJ92, ADG96, BVG97, HM98, Hin97, IKY^{+}00, KG97, WCC99]. compiler-generated [Hin97]. compiler-parallelized [HM98]. Compilers [KLW93]. compiling [HNS97a]. complete [SU96]. Complex [PG00]. complexity [VNL98]. component [SHGD98, WKE98]. component-based [SHGD98]. Composing [GFS99]. composition [EAK00, PD99, SRW^{+}00]. compositional [AGW97]. compressible [LL96]. compression [JT96]. Computation [BF92a, Fox91b, GOJ89, Lap92, WCB91, Zom92, ZM92, CDR99, CCCL97, Cun97, DBRF96, GHSJ98, LGFK99, MJ96, WF97]. Computational [Han93a, Han93b, Hoc95, AGD97, Fox97a, Fox97b]. Computations [Fox89, Ger90, MF92, AFH98, CSS98, CHMZ94, HHKP97, JJ96, WG95]. Computer [AAB^{+}93, BH92, CP92, Hel93, ZS93, BV00, BN00, DM98, GRS^{+}97, GLV96, HKSK97, Kon99, WKE98]. Computers [Bru92, CSTG92, CPV93, DHN93, Dem91, JJ92, MBF^{+}90, Bat91, BGKR99, CDW94, Cho98, DD97, HHGW95, JJ96, LRN97, PdOG99, Pir97, hSM97, XMLL95]. Computing [Bak99a, Bak99c, Cor98, Fox89, Fox91a, FFH^{+}93, HV93, HWPG92, Mes91, Sun90, AISS97, ASLV95, BKKK98, BCF^{+}97, BBC^{+}97, BAGGB99, CKV98, CMM^{+}97, CFLH99, CCI^{+}97, DR99, Fer98, FS98, FF97, Fox98, GS92, GS97b, GS98, HLA95, Han94b, HPPF94, HBKR96, HM95, KB98, MMV98, MKD99, NBK^{+}00, RPL96, SHW98, SBD^{+}96, SMS00, TDB00, TSS98, WJH^{+}97, XW00, YC97, ZTLV98, ZHS99, RSH97]. conceptual [UTS00]. Concurrency [Han00, Leó00, KRS96]. Concurrent [AM00a, ALV^{+}91, CSTG92, PR92, RIST00, CDW94, Cho98, GS92, HG99, MR96, Nas99, Phi00b, RR96]. Conference [Gue00, Fox00a, Fox00b, Fox00c]. Configuration [TKKN94]. Conjugate [Han98]. Connected [DHN93, MMN95]. Connection [Hin97, AM91, Ber94, MT91, SW91, ZS93]. connections [TSV94]. conservative [AR94]. constraints [YTPS95]. Constructing [FDC99]. Construction [JMM90, Dav97, GJGC96]. constructs [BN00]. Contention [ZW93]. context [FHC98]. Contextual [BGZ97]. continuous [CCCL97]. Contour [HHKP97]. Control [SBBG89, CRS98, KWB^{+}99, PH00a, SHPT00]. Controller [PDH91]. Controlling [MF93]. Convection [ES92, WF97]. Convergence [MR96]. conversion [YTY00]. converter [FLQZ97]. Converting [MJC99, Dem98]. convex [DT98]. Cooperating [Phi00a]. cooperation [AW97]. coordination [LHMZ00]. CORBA [BSP00, CPM00, ZTLV98]. CORBA-based [CPM00, SRW^{+}00]. core [DD00, HN00]. corner [MW98]. correction [Fid97]. correlation [Cor98]. cost [The96]. Coupled [Whe90, Whe91, ACF99, LK96, vL99]. COWs [HG99]. Cracking [GPO98]. CRAFT [WCC99]. Cray [BMT96, CRM94]. creation [HG99]. curved [Nak99]. cutting [NASW98]. cyclic [WO96].

D [LF91, RPB91]. DACAPO [Gra91]. DACAPO-III [Gra91]. DAEs [ZP99]. DAP [WCB91, ES92]. DAP510 [CWB94]. DARP [AFH98]. Data [BKU91, BF92a, CP92, CHMZ94, Dag92, MF92, MPS93, NQ93, SBBG89, SSB92, TDD93, Wol93, AFH98, BR00, CZF^{+}98, CFKL00, CS95, CMF97, CQ95, CP98, FDC99, GLV96, HW96, JDA97, KK97, MF94, MJ96, MB99, SU96, TSH98, WO96, XLD97]. data-farm [FDC99]. data-flow [HW96]. Data-parallel [NQ93, CQ95, GLV96, JDA97]. database [CHB98]. Databases [Wil92, BMAT97, WDT^{+}99]. dataflow [CF96]. dataflow-based [CF96]. deadlines [SS95]. Deadlock [CSTG92]. Deadlock-Free [CSTG92]. debugging [BGZ97, SSG97]. Decentralized [XLD97]. Deck [BP92]. Declarative [DRW90]. Decomposition [CP92, DMM98, PdOG99, van90]. definition [LP00]. degree [PRA00]. Delaunay [CLMP95]. Delta [HLJTZ94, TFB94]. demonstrating [BCE^{+}99]. dense [GS97a, LSF97]. Density [MRCH^{+}93]. Dependence [SF93, Wol93]. Dependency [HS90]. depth [AW97]. Design [CSTG92, Chr91, Gor93, IKY^{+}00, JS00, NASW98, PDH91, AC98, CHB98, DD00, Kar99, KWB^{+}99, WJH^{+}97].

Designing [BF92b, CPM00]. designs [UTS00]. Determining [CDMW94, DT98]. develop [CRS98]. Developing [BDP^{+}98]. Development [SK00, ABC^{+}00]. developments [Per92]. dialect [YSP^{+}98]. different [BN00]. differential [MPR95, TFB94]. differential-algebraic [MPR95]. diffusion [Ber94]. Digital [Woo91]. dilemma [TSV94]. Dimension [Wu97, Cor98]. Dimension-exchange-based [Wu97]. dimensional [ES92, LL96, LGFK99, LSF97, MMN95]. directories [HW00]. discrete [BSS98, Cow98, VO98a]. discrete-event [BSS98, Cow98]. Discussion [ST92]. Displacements [SDT91]. Display [VBFW91]. dissection [PRA00]. Distributed [AAB^{+}93, AMB^{+}93, Bak99a, BKTJ92, BGL92, BD93, BSS91, BCB00, DAB97, FFH^{+}93, Ger90, GOJ89, HV93, JJ92, JL95, KWB^{+}99, Lap92, LDP89, MR91, MRCH^{+}93, MPS93, MB99, Sun90, TV90, WHF^{+}89, WGC93, dMP92, Ahm95b, Ahm95a, AFH98, AM00b, ASLV95, ACF99, BDL00, BBC^{+}97, BRTK94, BAGGB99, BGKR99, CRS98, CKH91, CC97b, CMM^{+}97, CT00, CDW94, Cho98, CHMZ94, CP98, DD94, EE95, FS98, GGS00, GS97b, GS98, HKSK97, HPPF94, HW00, HHGW95, HYI98, HHA96, ICB00, JT96, JA98, JS00, JJ96, JCS98, Kar98, Kea99, KAS98, LRN97, LCLL97, Man98, MF94, MMV98, NR94, PdOG99, PA95, Phi00a, Pir97, SHGD98, TBBK94, Tär94, TKKN94, VO98a, VO98b, WLR^{+}00, XLD97, ZHS99]. Distributed-Memory [LDP89, AAB^{+}93, Lap92, Cho98, CHMZ94, MF94]. Distributing [Woo91]. distribution [CP98, Fah96, HS94, LL95, LLL98, LL99]. distributions [WO96]. division [Lin97]. Do [Han94a, NHZ96]. Do-loop-surface [NHZ96]. DOGMA [JCS98]. domain [PdOG99]. dose [VSKN96]. DOVE [JA98]. downhill [CB98]. Driving [Zim91]. DSM [YC97]. DSP [LZM96]. duplication [SCM95]. Dusty [BP92]. Dynamic [EAK00, HLA95, LLL98, PFS^{+}93, SDT91, WB95a, WMC98, Wil91, Ahm95a, BSD94, CP95, CYL95, GARR00, HBE98, Pir97, Tär94]. Dynamically [HTM91, NM00, SBS93]. Dynamics [PS96, Zom92, AGD97]. Dynamo [Tär94].

Echidna [JJ92]. EcliPSe} [KRS96].

editing [PH00a]. Editor [Han00]. Editorial [Bak96, Bak99a, Bak99b, Bak99c, Fox97a, Fox98, Fox00a, Fox00b, Fox00c, KZ93, Pri91, Wad99, vD96]. education [BCF^{+}97, WCL97]. effect [Cla97]. Effective [MJ96]. Effects [MR91]. Efficiency [MMV98]. Efficient [AC98, CW91, HR92, LFA98, MCB^{+}99, Zom92, Han00, KSK00, Leó00, PRA00, PHN00, RDLG99, WC00]. efficiently [KP98]. eigenproblem [BV00]. Eigenvalues [WCB91]. Eigenvectors [WCB91]. Electron [MRCH^{+}93, VSKN96]. element [CYL95]. elimination [Fur96]. embedded [RR99, YTPS95]. EMMA2 [ABMR91]. Empirical [SS94]. emulation [ZS94]. enabling [BCF^{+}97, WC00]. encryption [Sch99]. encryption/authentication [Sch99]. end [SHGD98]. Engineering [KABG^{+}91, LTH^{+}99, Wol93, BSP00, DAB97, Fox97a, Fox97b, FF97]. Environment [NHMW91, AFH98, AC98, CF96, DF97, DAB97, ICB00, JA98, LCLL97, RBF98, SRW^{+}00, WLR^{+}00, SBS93]. Environments [Bak99a, CWB94, HSvdL^{+}96, SS94, TSS98, WG95]. Equation [Boi90]. equations [CH99, Han98, HW96, TFB94]. ES-Kit [CKH91]. ESA [DTY91]. ESA/NASA [DTY91]. ESL [HNS97a]. ESM [BKU91]. ESP [WGC93]. Estelle [JJ92]. estimating [Fah96]. estimation [KSK00, Kon99]. estimator [Fah96]. European [BC00, BCE^{+}99, Gue00, Per92]. Evaluating [BH92, ZK99]. Evaluation [CSTG92, CLMP95, PR92, CHB98, HYI98, HASnP00, IKY^{+}00, MR98a, VO98b, SBS93]. event [BSS98, Cow98, VO98a]. events [CRS98]. EWOMP’99 [BC00]. Exception [Ada91]. exceptions [KP98]. exchange [Wu97]. Execution [BSS91, GGS00, Ols99, RR99, RRMA99, YQ95]. executions [NM00]. Exemplar [JR92]. Existing [SF93]. Experience [BT93, JM89, SSS98, Han00, Leó00, Rin98]. Experiences [BFFL97, BMT96, GWP93, HM95, SBD^{+}96, WGC93]. Experiment [PR92, FGN^{+}98, SS96]. Experimental [VO98b, BI98, DM96, LZM96, MGG00]. Experimentation [SBBG89]. Experiments [CCF^{+}97, CRM94, MBS91, RM91, van90, BDV^{+}98]. exploitation [GAM^{+}00]. Exploiting [DM96, BG97a, GHM98]. extend [Lin97]. Extended [HDF00]. Extending [YTY00]. extensible [SHW98]. Extension [Ada91]. extensions [CZF^{+}98]. extrapolation [RR97].

F [Han00, Leó00]. F90 [WCC99]. F90/HPF [WCC99]. Factor [ALV^{+}91, AMVA95, CH99]. factorization [DD00, GS97a, MP96]. Fail [KRS96]. Fail-safe [KRS96]. Falcon [GESV98]. family [MP96]. Farm [MP93, FDC99]. farms [MR98b]. Fast [BS94a, dMP92, CE95]. faster [Han94a]. fatally [Pug00]. Fault [Bal92, YQ95, The96]. fault-tolerant [The96]. feasibility [JPK98]. feedforward [AB98]. FEM [DMM98]. FEM-meshes [DMM98]. FFT [AG98]. FFT1 [HHGW95]. FFTs [GTV98]. file [BRTK94]. files [JK00, MLF97]. filesystem [GS98]. financial [Kea96, Kea99]. fine [LFA98, PH00a, The96]. fine-grain [LFA98, The96]. fine-grained [PH00a]. finite [CYL95]. First [BC00]. Flagship [MK93]. flawed [Pug00]. Flexible [SHPT00, GAM^{+}00]. FLITE3D} [BMT96]. Flow [MO91, Wil89, HW96, RIST00]. Flows [AMP^{+}91, LL96]. Fluid [AMP^{+}91, Wil89, AGD97, Pir97]. footprinting [LA95]. formal [BSP00, PDH91]. Fornax [VO98a]. Fortran [ADG96, BP92, CWB94, EHJ^{+}93, FLQZ97, KLW93, KP91, SF93, SSG97]. Fortran-77 [SF93]. Fortran-to-Java [FLQZ97]. Fourier [Fur96]. Fourier-Motzkin [Fur96]. fractal [JT96]. Fracturing [NHMW91]. Framework [Sun90, CMM^{+}97, CFLH99, DW97, Kea99, LP00, NR94, SHW98, SU96, TMN^{+}98, TDB00, WG95, WCC99]. Free [CSTG92, Nas99]. front [SHGD98]. front-end [SHGD98]. fully [EE95]. functional [HHLM95]. fuzzy [EKR97].

G [LTH^{+}99]. Gain [GWP93]. Galactica [CDMW94]. gap [Kar98]. garbage [Phi00a]. García [Han00, Leó00]. Gate [Woo91]. Gate-level [Woo91]. Gateway [HAFY00]. GEC [RIST00]. General [AAB^{+}93, CGM^{+}97, DM98, FF97, LF98]. generalized [ZK99]. generated [Hin97]. Generation [HWPG92, LZM97]. generator [CP98]. generic [FDC99, HSvdL^{+}96, MJC99]. Genesis [AAB^{+}93, HHGW95]. genetic [AD95, MF94]. geographically [BGKR99]. geophysical [JPK98]. Global [BWA94, DHN93, Lap92, SBBG89, AISS97, CRS98, JJ96, NBK^{+}00, SS95, Wu97, XW00]. GLU [JDA97]. gradient [Han98]. grain [BBC^{+}97, LFA98, The96]. grained [HHKP97, PH00a]. Grande [Fox00a, Fox00b, Fox00c]. granular [JDA97]. Graph [Boi90, RM91, GM00]. Graphs [HS90]. Grid [HWPG92, DW97, LSF97, LK96]. Grids [HWPG92, HBG00]. group [JCS98]. groups [WMC98].

H [AM00b, LTH^{+}99]. H-SWEB [AM00b]. Habanero [DAB97]. halo [PRA00]. halos [Ben00]. Handling [Ada91]. Handwritten [JMM90]. Hardware [Gra91]. harmonic [RG00]. Harness [CW91]. HDF [MLF97]. health [BCF^{+}97]. Heat [MP93]. Hector [RRG^{+}99]. Hermitian [WCB91]. heterogeneity [HS94]. Heterogeneous [BD93, NQ93, SMS00, HLA95, LAKL00, Lil95, MMV98, WB95b, WG95, YC97]. heuristics [CMF97]. hiding [Tog97]. hierarchical [AM00b, ZS94]. High [Dem91, FFH^{+}93, GHM98, Gor93, HV93, HBKR96, ZS94, AFH98, BDOP95, BS94b, BGKR99, BSW^{+}00, CL97a, Fox98, GFS99, HPPF94, JDA97, OP99, RGSB94, SBD^{+}96, SSG97, YSP^{+}98]. High-level [Gor93, BDOP95, JDA97, OP99]. High-performance [Dem91, FFH^{+}93, GHM98, HV93, HBKR96, ZS94, BS94b, BGKR99, CL97a, Fox98, GFS99, HPPF94, YSP^{+}98]. Highly [Zom92]. Histogram [GOJ89]. Hopping [Wu95]. ‘HP [CCF^{+}97]. HPC [BDV^{+}98, BCE^{+}99]. HPCC [CFLH99]. HPF [Ben00, Bri00, OP99, WCC99]. HPJava} [CZF^{+}98]. hull [DT98]. hybrid [ARS00, LZM96, NR00]. Hybridizing [PRA00]. Hypercube [ALV^{+}91, BR89, Dun92, ZS94, ZK99]. Hypercubes [HR92, ZS93, Han94a, Wu97]. hypothesize [Lin97]. hypothesize-verify-extend [Lin97].

I-WAY [FGN^{+}98]. I/O [WC00]. IBGF [BMAT97]. IBM [BFFL97, BRSR93]. IceT [GS97b]. Identification [BC91, LF91, Lin97]. idling [YQ95]. II [Fox97b, WWD92]. III [Gra91]. Image [CP92, FHC98, JT96, LZM97, MCB^{+}99]. image-processing [FHC98]. Images [LF91, MT91, ND94]. Imbalance [MR91]. implement [KP98]. Implementation [ALV^{+}91, AH91, CSTG92, CP92, EE95, MF93, ST92, WDW91, AHS93, AAFG94, ANH00, BS94a, BDL00, BS94b, BB00, CGM^{+}97, DD00, IKY^{+}00, LHMZ00, LRN97, Lin97, MCB^{+}99, NASW98, SB99, SCL97, SU96, WJH^{+}97]. implementations [DT98, JT96, JL95, VO98b]. implemented [DAB97]. implementing [ACF99, DHS99, HW00]. implicit [BG97a]. Incompressible [AMP^{+}91]. incremental [CLMP95]. Index [Ano95, Ano96, Ano97, Ano98, SS95]. Indigo [KAS98]. Industrial [vR89, Bak99b, BCE^{+}99]. information [Kea96, SS97]. infrastructure [BKKK98, FGN^{+}98]. Initial [ST92]. injured [Wu97]. Inmos [CW91]. Inspection [Mir91]. inspired [TDB00]. instruction [AGW97]. Instructional [LA95]. Instrument [DTY91]. Instrumentation [MO91]. integer [MRR^{+}96, YTY00]. integer-reference [YTY00]. Integrals [Lap92]. integrated [CSS98]. Integrating [BR00, CF96, CFLH99]. integration [AA97, SBS93]. Intel [BH92, HLJTZ94, TFB94]. intensity [AW97]. Interactive [HNS97a, KMT93, CFLH99, Dav97]. interactive-oriented [Dav97]. interconnected [Lil95]. Interconnection [ZW93, The96, ZK99]. Interface [TDB00, Bak00, CFKL00, CvE00, FTT97]. interface-based [Bak00]. interfaces [WJH^{+}97]. Internet [CCI^{+}97, ZTLV98]. Internet-based [CCI^{+}97]. interoperability [BDV^{+}98]. interpretation [AGW97, ND94]. Inverse [BV00]. Invocation [RWB97, TTK98, vNMB^{+}00]. IP [DHS99]. IPSC [HHGW95, BH92, BR89]. iPSC/2 [BR89]. IPSC/860 [HHGW95, BH92]. irregular [Ben00, HASnP00, Wu95].

Irregularly [Yan00, EAK00]. Issue [Bak99c, Bak00, BC00, Fox00a, Fox00b, Fox00c, Yan00, Gue00]. issues [AISS97, NBK^{+}00, PA95]. Iterated [HS90, RR99]. Iterative [ST92, CHMZ94].

Jacobi [GHvdGV97]. Jacobians [ZM92]. Jade [Rin98]. Jaguar [WC00]. Japanese [JMM90]. Java [Fox00a, Fox00b, Fox00c, CCF^{+}97, AFH98, AISS97, AGD97, ANH00, Bak00, BKKK98, BCF^{+}97, BBC^{+}97, BG97a, BVG97, BG97b, BSS98, BDP^{+}98, BDV^{+}98, BK97, BSW^{+}00, CKV98, CZF^{+}98, CGJ^{+}00, CFKL00, CDD97, CCCL97, CvE00, CC97b, CCI^{+}97, CL97a, CL97b, Cla97, Cow98, Dav97, Dem98, DF97, DAB97, EKR97, Fer98, FTT97, Fox97a, Fox97b, FF97, FLQZ97, Fox98, GHM98, GPO98, GTV98, GS97b, HNS97a, HW00, HN00, HYI98, HAKN97, HNS97b, IKY^{+}00, ICB00, JPK98, JK00, KB98, KP98, MLF97, NM00, PH00a, PS00a, PZ97, PG00, PHN00, Pug00, RWB97, RBF98, RSH97, SHW98, SS98, SRW^{+}00, SMS00, SGB97, SHGD98, Sta00, TMN^{+}98, TDB00, VO98a, WCL97, WC00, YC98, YSP^{+}98, YC97, YTY00, ZTLV98]. Java-based [AFH98, AISS97, DAB97, YC98]. Java/CORBA [SRW^{+}00]. Java/CORBA-based [SRW^{+}00]. Java/DSM [YC97]. Java/HPC [BDV^{+}98]. JavaNOW [TDB00]. JavaParty [PZ97, PH00b]. javar [BVG97]. JavaScript [DF97]. JavaVM [KG97]. Javelin [CCI^{+}97, NBK^{+}00]. Javia [CvE00]. JET [PSS97]. JNI [GGS00]. JNI-wrapped [GGS00]. Job [HS90, vL99]. jobs [YQ95]. join [ARS99, BMAT97]. JPVM [Fer98]. JSPICE [SHGD98]. JTED [Cow98]. Just [CL97a, KG97, IKY^{+}00]. Just-in-time [CL97a, KG97, IKY^{+}00]. JWarp [BSS98].

Kan [JS00]. Kanerva [HKSK97]. Kernel [CDL^{+}93, ACF99, Dav97]. Kernel-Kernel [CDL^{+}93]. kernels [MR98a]. Kit [CKH91]. knapsack [MRR^{+}96]. knowledge [DM96]. knowledge-based [DM96]. KSR1 [CB96]. Kutta [RR99].

L [Han00, Leó00, BDOP95]. Label [Mir91]. Laboratory [DF97]. language [BDOP95, FF97, GS98, Han94b, LAKL00, LHMZ00, Per92, RR96, VNL98, vRvGS97]. Languages [DRW90, BR00, Phi00b, SU96]. LAPACK [Dem91]. Large [JPK98, JMM90, LL96, ST92, Wal90, WCB91, ASLV95, DD94, KSN97, MPR95]. Large-scale [JPK98, JMM90, LL96, ST92, Wal90, MPR95]. latency [CDMW94, Tog97, WRMB98]. Lattice [BRSR93]. leaks [PS00a]. León [Han00]. Letter [Han00]. Level [Bru92, Fox89, BDOP95, BG97b, Gor93, HHKP97, JDA97, KAS98, OP99, SCL97, Woo91, CGM^{+}97]. Levels [MK93]. libraries [BDP^{+}98, CDD97, GHM98, GGS00, Han00, HN00, Leó00, RDLG99]. Library [Dem91, BSS98, PD99, PS00b, SGB97]. like [CGJ^{+}00]. limitations [HN00]. Linda [BWA94, LA95, TDB00]. line [GESV98, ZM92]. Linear [Dem91, Fid97, Whe91, CDW96, CWB94, DD94, Han98, HW96, Kon99, TFB94]. Link [JM89]. listen [CRS98]. Load [Boi90, MR91, RR97, Wil91, AGMM94, BAGGB99, CYL95, CD99, HBE98, LL95, LLL98, LL99, MO95, MGG00, Nak99, SS95, SS97, SCL97, Tär94, Tog98, VO98b, WB95a, Wu97, XMLL95]. load-balancing [AGMM94, CYL95, WB95a, XMLL95]. Local [Ger90]. localities [CS95]. Locality [PH00b, TSH98]. Locally [DHN93]. locking [Nas99]. logic [SU96]. Longitudinal [TV90]. Loop [SF93, NHZ96]. Loosely [MF92, vL99]. low [HHKP97, KSK00, The96]. low-cost [The96]. low-level [HHKP97]. LPVM [ZG98]. LU [DD00, van90].

LU-decomposition [van90]. Lyapunov [CH99].

Machine [JM89, RM91, ANH00, WB95b, YTY00, AM91, Ber94, MT91, SW91, ZS93]. Machines [BSS91, ST92, TDD93, AGW97, CHB98, CP98, HHKP97, Han94a, LFA98, YC98]. Macromolecular [MRCH^{+}93]. magnetostatic [LGFK99]. management [Ahm95b, Ahm95a, BFFL97, BSD94, CHMZ94, Hin97, TKKN94]. Manager [NR94]. Managing [BGKR99]. Manifold [AHS93]. Manipulators [MF93]. manual [MJC99]. manufacturing [AA97]. many [ZP99]. mapping [CE95, CMF97, HD96, PdOG99, UTS00, YTPS95]. Mappings [ZS93]. Markov [MR96]. marshaling [CFKL00]. MARVIN [RPB91]. MasPar [GS97a]. Massive [TSV94]. Massively [Hel93, dR93, GLV96, Kon99]. matching [AW97]. math [SGB97]. Matrices [WCB91, TSH98]. Matrix [CDW94, HLJTZ94, Cho98, LSF97, PRA00, SGB97, TSH98, vdGW97]. Means [Zim91, PH00b]. measure [SS96]. Measurement [SDT91, NM00]. measurements [MMV98]. mechanism [Nas99, RWB97]. Mechatronics [Epp91]. Memory [BP92, BSS91, GOJ89, LDP89, MR91, MRCH^{+}93, MPS93, dMP92, AAB^{+}93, CDL^{+}93, CDW94, Cho98, CHMZ94, CP98, DD94, EE95, GOM00, HKSK97, HHGW95, JJ96, Kar98, KSN97, Lap92, LRN97, LZM96, LFA98, MF94, PA95, PS00a, Pir97, PS00b, Pug00, Tär94, XLD97, Tog98]. Mentat [GWP93]. Mesh [Wil89, Wil91, CT00, GHvdGV97, MJC99, MMN95, SB99]. meshes [CYL95, DMM98, LK96, MO95, WB95a]. Message [AW97, CPV93, DHN93, DD97, KK90, BWM98, CGJ^{+}00, CSS98, CH99, GGS00, Han00, Kar98, KSAC95, Leó00, LZM96, PS00b, RRMA99, RDLG99, WRMB98, Bak00, TDB00]. Message-passing [CPV93, DHN93, DD97, KK90, BWM98, CSS98, KSAC95, LZM96, PS00b, WRMB98]. message-passing/shared-memory [LZM96]. metacomputer [vL99]. metacomputing [ARS97, BFFL97, CKV98, FGN^{+}98, JCS98, PSS97]. Method [BRSR93, RWB97, WWD92, WD92, DD94, DM96, GHvdGV97, KSN97, TTK98, VSKN96, vNMB^{+}00]. Methodology [FSDC99, Han93a, ST89, Han00, Leó00, RDLG99, XW00, AAB^{+}93]. Methods [PDH91, PdOG99, RR97, RR99]. Metrobridge [RCZM91]. micro [ACF99]. micro-kernel [ACF99]. microkernel [AC98]. Migrant [MR98b]. migratable [TMN^{+}98]. Migration [AMB^{+}93, CP95, HBE98, SCL97]. MIMD [AB98, AGW97, AG98, CHMZ94, MRCH^{+}93, ST92, TSV94]. minimum [PRA00]. Mission [DTY91]. mixed [SK00]. mobile [ICB00, XW00]. Model [FFH^{+}93, Han93a, Han93b, BDV^{+}98, Dav97, KSS^{+}96, Kon99, LF98, MCB^{+}99, Pug00, vNMB^{+}00]. modeling [AGMM94, BMR^{+}00, BWM98, Fox97a, Fox97b, FF97, GS97a]. modelling [NA96]. Models [BC91, RRMA99, Wal00]. Monitor [WJC93]. Monitoring [ZW93, GESV98]. Monitors [KP98]. Monte [MR96, SK00, VSKN96]. Monte-Carlo [MR96]. motion [KSK00]. Motzkin [Fur96]. Movement [BF92a]. MOVIE [FFH^{+}93]. MP [CRM94, GS97a, RR96]. MP-2 [GS97a]. MPE [BGL92]. MPI [TDB00, CGJ^{+}00, CFKL00, Man98, PS00b, RRG^{+}99, SK00, WO96, Wal00]. MPI-like [CGJ^{+}00]. MPJ [CGJ^{+}00]. MPOOL [PD99]. MPP [MR98a]. Multi [Bat91, BT93, HWPG92, JM89, MR91, MBS91, ARS99]. Multi-block [HWPG92]. Multi-computers [Bat91]. multi-join [ARS99]. Multi-processor [MR91, MBS91]. Multi-scale [Bat91]. Multi-Transputer [JM89, BT93]. multibody [PS96]. multicast [AC98]. Multicomputer [MF92, van90, DV94, GH95]. Multicomputers [Han93a, SBBG89, WJC93, MGG00, Tär94]. Multidimensional [Lap92, MT91, BV97]. multidisciplinary [KWB^{+}99]. Multilevel [Gra91, BS94a, GAM^{+}00]. Multiparadigm [FS98]. multiple [CS95, CF96, vL99]. Multiplication [CDW94, Cho98, HLJTZ94, LSF97, vdGW97]. multiply [TSH98]. multiprocessing [The96]. Multiprocessor [ABMR91, BP92, BR89, CDL^{+}93, KK90, Lap92, LDP89, RPB91, ZW93, ARS99, ARS00, GOM00, YTPS95]. Multiprocessors [SBW91, Whe90, Whe91, dMP92, AG98, CH99, DD94, KSN97, MF94, XLD97]. Multiprogramming [LDP89]. Multiresolution [Nak99, LZM97]. multistage [The96]. multithread [CS95, ZG98]. multithreaded [BGZ97]. MUSCL [LL96].

N [HNS97a]. N3S [LL96]. N3S-MUSCL [LL96]. NanosCompiler [GAM^{+}00]. NASA [DTY91]. Native [GS98, HSvdL^{+}96, BG97b, GHM98]. Native-language-based [GS98]. natural [WF97]. NCSA [MLF97]. NCUBE [JMM90]. nCUBE2 [GH95]. Near [Whe91]. Nearest [XMLL95]. Nearest-neighbor [XMLL95]. neighbor [XMLL95]. nested [PRA00]. net [CB96, CDMW94]. Nets [Gor93]. Network [GS92, HTM91, JMM90, MBS91, NQ93, The96, ZW93, Zim91, AB98, ARS00, BFFL97, BKKK98, Cor98, Fer98, Fox98, GS98, NA96, TSV94, TSS98, ZK99]. Network-based [GS92]. networked [CQ95, HM95, RRG^{+}99, WJH^{+}97]. Networks [CPV93, AGMM94, BV00, CD99, DHS99, LAKL00, RGSB94, TDB00, Wei99].

Neural

[CPV93, JMM90, AGMM94, CB96, TSV94]. neutral [RIST00]. Nexus [FTT97, TKKN94]. Ninflet [TMN^{+}98]. No [GWP93]. Nodes [MF92]. non [BRTK94, SCM95, hSM97]. non-clustering [SCM95]. non-periodic [hSM97]. non-volatile [BRTK94]. nonlinear [LGFK99]. notation [BSP00]. note [BG97b]. Notebook [WJH^{+}97]. NOWs [BR00]. numbers [PG00]. Numerical [AMP^{+}91, ES92, Pir97, BDP^{+}98, CDD97, SS98, HS94].

O [WC00]. Object

[CFKL00, Epp91, Gue00, TBBK94, ACF99, BDV^{+}98, CKH91, CSS98, HYI98, JS00, JCS98, Lin97, Phi00b, PD99, UTS00]. Object-based [TBBK94, CSS98]. Object-oriented [Epp91, Gue00, ACF99, CKH91, Phi00b, PD99]. object-relational [UTS00]. Objects [LF91, CE00, HW00, JA98, JL95, PZ97, TMN^{+}98]. observing [JJ96]. Occam [Hoa91]. Occam2 [ST89]. ocean [Wal00]. OdinMP [BB00]. OdinMP/CCp [BB00]. On-line [ZM92, GESV98]. One [Mes91]. onto [ZS93]. Open [FFH^{+}93, Sta00]. OpenMP [ABC^{+}00, BN00, Bri00, BC00, BB00, GAM^{+}00, HASnP00, KPO00, SHPT00, SK00, Wal00]. OpenMP/MPI [SK00]. Operating [ABMR91, SBBG89, WHF^{+}89, TKKN94, vL99]. operation [Cun97]. operational [SBD^{+}96]. operations [BI98, HDF00]. Optimal [JH00, TV90, HBE98]. optimisation [CB98, GRS^{+}97]. Optimization [KLW93, MF92, CMF97, HAKN97, LF98, PH00b, SS98]. optimizations [AMVA95, CL97a, IKY^{+}00, WCC99]. optimizer [Cla97]. Optimizing [Ben00, BK97, CL97b]. Opus [LHMZ00]. ordering [PRA00]. Ordinary [Kon99]. ordination [Kea99, OP99, PD99]. organization [HD96]. oriented [ACF99, CKH91, Dav97, Epp91, Gue00, LRGS97, Phi00b, PD99]. Origin2000 [Bri00]. out-of-core [DD00]. Outer [SF93, Cun97]. outer-product [Cun97]. overheads [SS94]. overlap [CDR99]. overlapped [KK90]. overlapping [GHSJ98]. Overrelaxation [dR93]. overview [AHS93, Per92].

P [LTH^{+}99, BDOP95, Fah96]. Package [HS90, CSS98]. Pain [GWP93]. pair [Lil95]. Paradigm [Wil92, BBC^{+}97]. paradigms [CF96]. PARADOX [WB95b]. Paragon [hSM97]. Parallel [AMB^{+}93, AH91, AR94, BC91, Bal92, BKTJ92, BSD94, BV97, BH92, BMAT97, BS94b, Bru92, CPV93, CP92, CGM^{+}97, CH99, CB96, Dag92, DRW90, DCGC92, DT98, DHN93, DMM98, FHC98, Fox89, Fox91a, GOM00, Gor93, HKSK97, Han93b, HS90, HWPG92, Hel93, Kea96, KMT93, KK90, KP91, LRN97, LGFK99, LZM96, LZM97, MPR95, MK93, MF93, MW98, Mes91, MPS93, MGG00, ND94, Per92, RM91, RR99, SBBG89, SB99, SSB92, Sun90, TDD93, Tog97, Tog98, TSH98, VBFW91, Wal90, WF97, WKE98, WCB91, Whe90, Wil89, ZP99, dR93, Ahm95b, Ahm95a, AW97, AAFG94, AG98, BDOP95, Bak00, BFFL97, BN00, BR00, BSS98, BCE^{+}99, CP95, CC97a, CZF^{+}98, CCCL97, CT00, CF96, CDW96, Cho98, CCI^{+}97, CYL95, CD99, CWB94, CB98, Cow98, CQ95, Cun97, DD94, DD00]. parallel [DF97, DR99, DM98, DW97, EE95, Fah96, Fer98, FSDC99, FF97, GM00, GRS^{+}97, GLV96, GHM98, GARR00, GJGC96, GESV98, Han94b, HHLM95, HD96, HG99, HSvdL^{+}96, HM95, HHA96, ICB00, JPK98, JDA97, JJ96, KB98, KSS^{+}96, KSK00, Kon99, LL96, LAKL00, LSF97, Lin97, LTH^{+}99, LF98, MR98b, MP96, MRR^{+}96, MCB^{+}99, MJ96, NHZ96, NA96, NQ93, NASW98, PdOG99, Pir97, RPL96, RG00, SMS00, SSS98, SU96, hSM97, SS96, TMN^{+}98, TDB00, TFB94, VNL98, Wad99, WG95, WDT^{+}99, XMLL95, XLD97, XW00, YC98, vNMB^{+}00, vRvGS97, CDW94, Gra91, JR92, ZHS99]. Parallel-Processing [SSB92]. parallelisation [MJC99]. parallelising [BMT96]. Parallelism [MBS91, RDLG99, Wil92, BG97a, BR00, DM96, GAM^{+}00, HM98, KWB^{+}99, LFA98, SHPT00, Han00, Leó00]. Parallelization [LK96, SF93, BG98, CLMP95]. parallelized [HM98]. Parallelizing [VSKN96, WWD92, WD92, ADG96]. parameterized [Fur96]. parameters [SS97, ZP99]. ParaPART [CT00]. PARLOG [ST89]. PARSE [GJGC96]. Part [WWD92, Bak99b, AAB^{+}93, HWPG92, HHGW95]. partial [TFB94]. Particle [BD93, Dag92, MO91, SMC92, Wal90]. Particle-in-cell [SMC92, Wal90]. partitioned [BMAT97]. Partitioning [HD96, Lil95, MO95, YTPS95, BS94a, CT00, CQ95, PA95, WG95]. Passing [TDB00, Bak00, BWM98, CGJ^{+}00, CPV93, CSS98, CH99, DHN93, DD97, GGS00, Han00, Kar98, KSAC95, KK90, Leó00, PS00b, RRMA99, RDLG99, WRMB98]. passing/shared [LZM96]. patterns [EAK00, PS00a]. PB [CDW96]. PB-BLAS [CDW96]. PDE [WB95a]. PDEs [HHA96]. Performance [AB98, AGMM94, AH91, ARS00, BH92, BN00, BRTK94, BWM98, GS97a, Hel93, HN00, HYI98, JT96, JR92, LF98, NA96, NM00, PFS^{+}93, SMC92, hSM97, WJC93, Wal90, WHF^{+}89, Wil91, ALCG95, AFH98, AR94, BS94b, BDV^{+}98, BGKR99, BSW^{+}00, CL97a, CS97, DV94, Dem91, DBRF96, DD97, DM98, Fah96, Fox98, FFH^{+}93, GHM98, GFS99, HV93, HPPF94, HBKR96, HASnP00, HHA96, JK00, JDZW99, MR98a, NHZ96, RPL96, SBD^{+}96, SK00, SSG97, VNL98, WDT^{+}99, YC98, YSP^{+}98, ZS94]. periodic [EE95, hSM97]. Person [Mes91]. Personal [Hoa91, BV00, ZTLV98]. Petri [Gor93]. Physical [Fox91b, MF92]. physics [WCL97]. PIC [CC97a]. Pipeline [AMVA95]. pipelined [AGW97, KSAC95, NA96]. pipelines [FSDC99]. placement [KB98]. Plasma [SMC92, Wal90]. platform [HSvdL^{+}96, HHA96, YC97]. Pleiades [WGC93]. Pleiades/ESP [WGC93]. PNPACK [RSH97]. pointers [Dem98]. Poisson [Boi90, dMP92]. Policies [PFS^{+}93]. policy [SS95]. poly [LSF97]. poly-algorithm [LSF97]. polyhedron [Fur96]. popular [HYI98]. Portability [RPL96, GGS00]. Portable [AAFG94, ADG96, AG98, Dem91, KP91, Man98, BB00, CS97, MCB^{+}99, Tär94, Wad99]. Porting [BP92, BCE^{+}99, HASnP00]. POST [ABC^{+}00]. powerful [VSKN96]. PowerPC [HSvdL^{+}96]. Practical [JM89]. Practice [Han00, Leó00, BK97, LSS00, MMV98, MJC99, RPL96]. PRAM [LSS00, LRGS97]. Preconditionings [Bru92]. predictability [RPL96]. Predictable [DHS99]. Predicting [RRMA99, HHA96, WDT^{+}99]. prediction [JDZW99, SBD^{+}96]. presence [YTPS95]. preservation [LA95]. Price [PNRC90]. Prime [ALV^{+}91, AMVA95]. Primitives [HR92, SBBG89]. priority [Man98]. probabilities [RSH97]. Problem [DCGC92, AD95, BA98, Ber94, CMF97, GRS^{+}97, MRR^{+}96, NASW98, WLR^{+}00]. problem-solving [WLR^{+}00]. problem-space [AD95]. Problems [KABG^{+}91, Yan00, Ahm95b, BS94a, LGFK99, Pir97, Wu95]. process [AA97, CP95, EAK00, HG99, LSF97, MR98b, WMC98]. processes [Ada91, KSS^{+}96]. Processing [CP92, SBH95, SSB92, VBFW91, CD99, CQ95, FHC98, GTV98, HG99, Kea96, LZM97, MCB^{+}99, MB99]. Processor [BKU91, ES92, LL95, MR91, MBS91, MMN95, NR94, PdOG99]. processors [CS95, GHvdGV97]. Product [Mir91, Cun97]. profiling [Tog97]. Program [BP92, CRM94, DRW90, Gor93, Han93b, NHZ96, PA95, SRW^{+}00, Tog97, Tog98]. Programming [Bal92, DRW90, DF97, Gue00, Han93a, HS90, KMT93, KSAC95, KP91, NQ93, AM00a, ACF99, Bak00, BSD94, BBC^{+}97, BR00, CF96, CS97, GHM98, GARR00, HHLM95, HSvdL^{+}96, HNS97b, JDA97, LAKL00, LSS00, LRGS97, MR98b, MCB^{+}99, SSS98, Sta00, SS96, TBBK94, VNL98, vNMB^{+}00, vRvGS97]. Programs [AH91, EHJ^{+}93, Han93a, SF93, dR93, BWM98, CP95, CL97a, Fah96, GESV98, HD96, LTH^{+}99, Ols99, PH00a, RRG^{+}99, SSG97, Wei99, vRvGS97]. project [ABC^{+}00, HS94, BCE^{+}99]. Promises [HV93]. Prophet [Wei99]. Propulsion [HS94]. Prospects [Fox91a]. Prospero [NR94]. Protocol [SBS93]. Protocols [JJ92]. Prototype [JJ92, AGD97, BVG97, BG98, FLQZ97]. prototypes [JDZW99]. prototyping [AFH98]. Providing [PH00a]. PSEs [SRW^{+}00]. Psi [MJ96]. publication [Han94b]. PUMMA [CDW94]. purpose [DM98]. push [XW00]. PVaniM [TSS98]. PVM [GS92, SCL97, Sun90, ZG98]. Pyramid [ZS93].

QCD2 [HHGW95]. QR [DD00, MP96]. Quadratic [dR93]. Quadrature [Lap92]. quantum [SK00]. queries [ARS99, SBH95]. Queues [HS90, Man98].

Radar [BKU91]. Radiation [KABG^{+}91]. Radiative [MP93]. Radon [HTM91]. RAID [JH00]. ranking [HHKP97]. rapid [AFH98]. rate [KSK00]. Ray [JR92, GH95, LRN97]. RC5 [GPO98]. reaching [LP00]. Real [Bat91, CFMP91, TV90, VBFW91, GFS99].

Real-time [Bat91, CFMP91, TV90, VBFW91, GFS99]. realizations [HKSK97]. Recognition [JMM90]. Recognizer [CFMP91]. reconfigurable [HBKR96]. Reconfiguration [JM89, FHC98]. recursion [Han00, Leó00, RDLG99]. recursive [BS94a, CE95, GHSJ98, SBH95]. Recycling [HS90]. Redistribution [WO96]. Reduction [RM91]. reference [PS00a, RIST00, YTY00]. references [Dem98, LP00]. Reflective [TTK98]. Region [TDD93]. Region-to-Region [TDD93]. registers [JK00]. Regularity [PNRC90]. relational [BMAT97, UTS00, WDT^{+}99]. Relativity [AAB^{+}93]. relevance [BCE^{+}99]. Reliable [GJGC96]. remapping [XLD97]. Remote [PFS^{+}93, RWB97, HAFY00, PZ97, TTK98, YQ95, vNMB^{+}00]. Rendering [MPS93]. Replication [BKTJ92]. representation [NHZ96, Nak99]. Reproducible [Ols99]. research [AISS97]. resolving [LL95, LL99]. Resource [ARS97, Ahm95b, Ahm95a, NR94, Hin97]. Resource-aware [ARS97]. resources [HAFY00]. Response [Leó00]. Restructuring [EHJ^{+}93, BVG97]. Results [AAB^{+}93, BI98]. RISC [BRSR93]. RMI [BDV^{+}98, PHN00, RBF98]. Robot [MF93, PDH91, Zom92, ZM92]. Rock [NHMW91]. Rodríguez [Han00]. role [CC97b]. routines [DD00]. Routing [CSTG92, CW91]. row [DD94]. row-action [DD94]. RP3 [BF92a, BF92b]. Run [CP92, JM89, PA95, SBW91, OP99, RRG^{+}99, SS97]. Run-Time [JM89, CP92, PA95, SBW91, OP99, SS97]. Runge [RR99]. Runner’ [CFMP91].

safe [KRS96]. Sande [Han00, Leó00]. Santa [BA98]. scalability [NBK^{+}00, SB99]. Scalable [GM00, KB98, Nas99, NR94, NR00, Wu95, vdGW97]. ScaLAPACK [DD00]. scale [ASLV95, Bat91, JPK98, JMM90, LL96, MPR95, SBD^{+}96, ST92, Wal90]. Scanning [Fur96]. SCHEDULE [HS90]. schedulers [GFS99]. Scheduling [ARS99, AGW97, Ahm95b, Ahm95a, AM00b, CS95, HLA95, RG00, SCM95, SS95, WMC98, Wei99, Wu95]. schemas [UTS00]. schemes [RR97]. Schooner [HS94]. SCI [ZHS99]. SCIDDLE [ASLV95]. Science [Han93a, Han93b, Fox97a, Fox97b]. Sciences [SSB92]. Scientific [BSS91, BMR^{+}00, DAB97, FF97, Han94b, JA98, KK97, WJH^{+}97]. SCIPVM [ZHS99]. Scivis [KK98]. SDSC [Sch99]. SEA [Sch99]. Seamless [Bak99c, MKD99, CKV98]. search [ND94]. Seismic [AMB^{+}93, MB99, MB99]. SEL [SU96]. selection [DM98, SS97]. Self [HBG00, HD96]. Self-avoiding [HBG00]. self-organization [HD96]. semantic [LA95]. semaphore [HDF00]. semi [vRvGS97]. semi-automatic [vRvGS97]. Sensing [PDH91]. sensitivity [WRMB98, ZP99]. Sequence [AM91]. Sequent [AH91]. Sequent/Balance [AH91]. Sequential [BC91, Ada91]. serialization [CFKL00, PHN00]. server [AM00b, Kar99]. service [HG99, RGSB94]. Set [MPS93, CDW96, hSM97]. sets [AGW97]. shaped [HKSK97]. Shared [BP92, CDL^{+}93, PS00b, CSS98, GOM00, HW00, Kar98, KAS98, KSN97, LFA98]. Shared-Memory [BP92, CDL^{+}93, PS00b, GOM00, LZM96, LFA98]. shared-variable [CSS98]. sharing [CD99]. ship [BSD94]. side [Cla97]. side-effect [Cla97]. signal [GTV98]. SIMD [AGW97, Hel93, KLW93, Kon99]. Similarity [Hoc95]. simple [LP00]. simplex [CB98]. Simulated [PdOG99]. Simulation [AMP^{+}91, BD93, Dag92, ES92, Gra91, MP93, NHMW91, SMC92, Wal90, WHF^{+}89, Woo91, Zim91, AR94, BMR^{+}00, Cow98, EKR97, Fox97a, Fox97b, FF97, GFS99, Pir97, PS96, RG00, RIST00, VO98a, HS94, JMM90]. Simulations [Epp91, AA97, BSS98, DW97, HNS97a, TSV94, WCL97]. Simulator [Zim91]. Single [BDL00]. sites [vL99]. SLALOM [WGC93]. Slkye [AH91]. small [LK96]. Social [SSB92]. Software [FGN^{+}98, PR92, ST92, ALCG95, CF96, GJGC96, Kar99, TSH98, Wad99, WLR^{+}00, WKE98]. Solution [DCGC92, KABG^{+}91, dR93, CYL95, Han98, HW96, MPR95, NASW98, TFB94]. solution-adaptive [CYL95]. solutions [Ahm95b, LL96]. solve [BA98]. Solver [dMP92, EE95, LL96, NR00]. solvers [hSM97, WB95a]. Solving [Ber94, Yan00, CH99, DD94, GRS^{+}97, HHA96, PS00a, WLR^{+}00]. sort [Han94a]. Sorting [Dag92, Whe90, Whe91, DM98, NA96]. SP [BFFL97]. SP-2 [BFFL97]. SP2 [hSM97]. space [AD95, BDL00, KSN97, Nak99]. Spar [vRvGS97]. SPARCstation [MP93]. Sparse [ST92, CCCL97, DD94, HKSK97, NR00, PRA00, TSH98]. Spatial [Wil92, GRS^{+}97]. Special [Bak99c, Bak00, BC00, Fox00a, Fox00b, Fox00c, Gue00, Yan00]. specific [RR96]. Spectral [WWD92, WD92, BS94a, KSS^{+}96]. Speech [CFMP91]. Speed [Whe91, RGSB94, HHA96]. Speed-up [Whe91]. Speeding [BKTJ92]. speedup [Tog97]. spelling [Fid97]. SPICE [SHGD98]. Spin [BC91]. SPMD [BWM98, HG99, HNS97b, Sta00, Wal00, Wei99]. SR [Ols99]. stage [AH91]. starvation [Nas99]. starvation-free [Nas99]. state [LLL98]. static [Ahm95b, Fah96, PH00b, SCM95]. Stations [Chr91]. steering [GESV98]. Step [KK90, ZG98]. Step-wise-overlapped [KK90]. stitching [MW98]. Stochastic [AH91]. stock [NASW98]. storm [SBD^{+}96]. Story [Hoa91]. strategies [CLMP95, MJC99, MMN95]. strategy [Lin97]. striping [JH00]. Structural [SDT91]. Structure [Wil92, CC97a]. Structured [Yan00, BDOP95, EAK00]. structures [LZM97, SHPT00, ZS94]. Study [ZS93, AGD97, AR94, Bri00, Cun97, GFS99, HHKP97, JPK98, MGG00, SCM95, UTS00, WRMB98]. submissions [vL99]. subprograms [CDW96]. subscript [LP00]. subsystem [BS94b]. Successive [dR93]. suite [BSW^{+}00]. SUMMA [vdGW97]. Supercomputer [Fox89, TFB94]. Supercomputers [BD93, ZS94]. supercomputing [FTT97, vL99]. SuperPascal [Han94b]. Supersonic [Wil89]. SuperWeb [AISS97]. Support [BSS91, BDOP95, CCCL97, HG99, HS94, HAKN97, KAS98, LK96, LFA98, OP99]. Supporting [BGL92, ZTLV98, GAM^{+}00]. SUPRENUM [AAB^{+}93, HHGW95]. surface [NHZ96]. survey [Phi00b]. SWEB [AM00b]. Switched [HTM91]. symbolic [LTH^{+}99]. Symmetrical [Wu95]. synchronisation [BWA94]. Synchronization [DR99, Dun92, MR91, DV94]. Synchronous [MF92]. Synthesizing [CE00]. System [ABMR91, HS94, KABG^{+}91, KK90, MT91, MK93, RPB91, SBBG89, SDT91, WHF^{+}89, WGC93, AM00b, BAGGB99, CKH91, CQ95, FSDC99, FTT97, GS92, HPPF94, HAFY00, JDA97, JS00, KRS96, LAKL00, LRGS97, LZM96, Lin97, PS96, Sch99, Sta00, TKKN94, BRSR93]. System/6000 [BRSR93]. Systems [BKTJ92, DRW90, FFH^{+}93, MR91, MRCH^{+}93, Ahm95b, Ahm95a, BFFL97, BDL00, BSP00, BRTK94, CDR99, CRS98, CT00, CSS98, CHMZ94, DD94, DM96, DV94, HLA95, HW96, HHA96, Kea96, KSK00, MPR95, MMV98, MMN95, NR94, PA95, SS96, TBBK94].

T [Fah96]. T3D [BMT96]. tale [HW00]. Task [AD95, Lin97, BR00, PdOG99, PD99, SCM95]. Tasks [CP92, ARS99, Lil95, LLL98, OP99, SS95]. TC2000 [SMC92]. TCP [DHS99]. TCP/IP [DHS99]. technique [GHSJ98, Kar99]. Techniques [BKTJ92, KLW93, CGM^{+}97, HLA95, SCM95]. Technologies [FTT97, HYI98]. templates [FDC99, SSS98]. Test [Wol93, GOM00]. tetrahedral [SB99]. theory [BI98, BK97, GARR00, LSS00, MMV98]. Thermal [ES92]. thinning [MGG00]. thrashing [LL95]. threading [KPO00]. threads [MR98b]. Three [ES92, LL96, LGFK99, NA96]. Three-dimensional [ES92, LL96, LGFK99]. Tightly [Whe90, Whe91]. Tiling [CDR99]. Time [BSS91, JM89, TFB94, AR94, Bat91, CS95, CFMP91, CP92, CL97a, GFS99, KG97, OP99, PA95, RRMA99, SBW91, SS97, TV90, VBFW91, IKY^{+}00, WHF^{+}89]. Time-parallel [TFB94]. Timing [LF91, Woo91]. TINA [RPB91]. Tiny [CW91]. Titanium [YSP^{+}98]. Tmon [WJC93]. Toeplitz [BV00]. tolerance [YQ95]. Tolerant [Bal92, The96]. Tool [KMT93, ASLV95, BFFL97, BSP00, BG98, CT00, EKR97, MR98a, MB99, SSG97, Tär94, TSS98, WDT^{+}99]. Toolkit [BGL92, BCB00, HHLM95]. Tools [ABC^{+}00, KP91, FSDC99, GARR00, Wad99]. top [SCL97]. topologies [LSF97]. tori [CYL95]. Touchstone [HLJTZ94, TFB94]. TPVM [FS98]. Tracing [JR92, GH95, LRN97]. Tracking [SDT91]. trained [AB98]. training [CB96]. TRANS1 [HHGW95]. Transfer [MP93, SS95]. Transform [HTM91, WWD92, WD92]. Transformation [DRW90, KMT93, SBS93]. transformations [DR99]. Translating [ST89]. Transparent [RCZM91, PZ97]. transport [BS94b, KSS^{+}96, RGSB94]. Transputer [BKU91, CW91, DTY91, HTM91, Hoa91, HW96, JM89, KABG^{+}91, PDH91, RM91, WJC93, WDW91, Zim91, Zom92, ZM92, BT93, HSvdL^{+}96, NA96, PS96]. Transputer-based [DTY91, KABG^{+}91, RM91, WJC93]. Transputers [Chr91, Gra91, Mir91, RCZM91, Woo91, vR89, SBH95]. tree [HKSK97, Han94a]. tree-shaped [HKSK97]. Triangulator [CLMP95]. tridiagonal [EE95, hSM97]. TRIO [CPM00]. TurboNet [LZM96]. Twisted [Bru92]. Two [ALV^{+}91, AH91, Bru92, HW00, LSF97, MMN95]. two-dimensional [LSF97, MMN95]. Two-Level [Bru92]. Two-stage [AH91]. type [PH00b].

ubiquitous [FTT97]. UDP [DHS99].

UDP/IP [DHS99]. unconstrained [CB98]. Unification [CRM94]. uniform [HAFY00]. Universal [CDW94, ZTLV98, vdGW97]. Unix [MB99]. unpredictable [DHS99]. Unstructured [Wil89, Wil91, BS94a, DMM98, HBG00, MO95, MJC99, SB99, WB95a]. update [CDMW94]. Updating [Ger90]. usability [SS96]. Use [NHMW91, LP00]. used [VBFW91]. useful [Fah96]. Usefulness [CD99]. user [KAS98]. user-level [KAS98]. Using [CRS98, Chr91, CPM00, DF97, HWPG92, HS94, Mir91, RRG^{+}99, TDD93, AGMM94, AD95, BFFL97, Ben00, CCI^{+}97, Cla97, Fur96, Gor93, GJGC96, GH95, HBKR96, HASnP00, JDZW99, LL96, LLL98, MLF97, MJ96, RBF98, SHW98, SBD^{+}96, SS97, SMS00, SSS98, TMN^{+}98, WO96, WCB91, vNMB^{+}00]. utilisation [Tog98]. uv [Cun97].

Variable [AGW97, CSS98]. Variables [Ger90]. variation [WRMB98]. various [DD97]. vector [CS95, Cun97, TSH98]. vectors [LLL98]. verify [Lin97]. versatility [CC97a]. versus [KPO00, SCM95, Wal00]. very [KSK00]. via [PH00a, dR93]. video [KSK00]. View [Mes91]. virtual [ANH00, CvE00, JK00, KSN97, YC98, YTY00, DF97]. Visibility [GLV96, TDD93]. Vision [Bat91, CP92, RPB91, ZS93, HHKP97, WB95b]. VISO [AM00a]. Visual [SDT91, Zim91, BBC^{+}97, JDZW99, SRW^{+}00]. visualization [ALCG95, CFLH99, JA98, KK97, RBF98, TSS98]. Visualizing [PS00a, ZW93]. VLSI [Chr91, GOM00]. volatile [BRTK94]. Volumetric [MPS93]. volunteer [SHW98]. Voxel [Wil92]. voyage [BSD94]. vs [BI98].

W [The96]. W-Network [The96]. walks [HBG00]. Warp [WHF^{+}89]. wavefront [CH99]. wavelet [Nak99]. WAY [FGN^{+}98]. weather [SBD^{+}96]. Web [AA97, BBC^{+}97, CMM^{+}97, DF97, HNS97a, HAFY00, LCLL97, PSS97, VO98a, WKE98]. Web-based [CMM^{+}97, DF97, LCLL97, PSS97, VO98a]. Web/Java [BBC^{+}97]. WebFlow [BBC^{+}97]. WebSubmit [MKD99]. Wets [AH91]. Wide [AA97, vNMB^{+}00, BFFL97]. Wide-area [vNMB^{+}00]. windows [AR94]. wise [KK90]. WMPI [SMS00]. work [Fah96]. workload [LL99]. Workshop [BC00]. workstation [Tog97, Tog98, Wei99, ZHS99]. Workstations [BRSR93, NQ93, CD99, Cor98, RRG^{+}99, TDB00, YQ95]. World [AA97]. wrapped [GGS00]. WWW [AM00b, WJH^{+}97].

Y-MP [CRM94].

References

Adapalli:1997:WWW

[AA97] Sridhar Adapalli and Krishna Addepalli. World Wide Web integration of manufacturing process simulations. Concurrency: Practice and Experience, 9(11):1341–1350, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13844;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13844&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Addison:1993:GDB

[AAB^{+}93] Cliff Addison, James Allwright, Norman Binsted, Nigel Bishop, Bryan Carpenter, Peter Dalloz, David Gee, Vladimir Getov, Tony Hey, Roger Hockney, Max Lemke, John Merlin, Mark Pinches, Chris Scott, and Ivan Wolton. The Genesis distributed-memory benchmarks. Part 1: Methodology and general relativity benchmark with results for the SUPRENUM computer. Concurrency: Practice and Experience, 5(1):1–22, February 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Averbuch:1994:PPI

[AAFG94] A. Averbuch, D. Amitai, R. Friedman, and E. Gabber. Portable parallel implementation of BLAS 3. Concurrency: Practice and Experience, 6(5): 411–459, August 1994. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Abbas:1998:PBT

[AB98] Hazem M. Abbas and Mohamed M. Bayoumi. Performance of a backpropagation trained feedforward network on an MIMD architecture. Concurrency: Practice and Experience, 10(14):1223–1234, December 10, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=40005933;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=40005933&PLACEBO=IE.pdf.

Adhianto:2000:TOA

[ABC^{+}00] L. Adhianto, F. Bodin, B. Chapman, L. Hascoet, A. Kneer, D. Lancaster, I. Wolton, and M. Wirtz. Tools for OpenMP application development: the POST project. Concurrency: Practice and Experience, 12 (12):1177–1191, October 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76500357/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500357&PLACEBO=IE.pdf.

Appiani:1991:EMO

[ABMR91] E. Appiani, D. Bianco, L. Merlo, and L. Roncarolo. The EMMA2 multiprocessor operating system. Concurrency: Practice and Experience, 3(6):541– 557, December 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Averbuch:1998:EMD

[AC98] A. Averbuch and A. Cohen. Efficient multicast design in a microkernel environment. Concurrency: Practice and Experience, 10(7):497–548, June 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10008691;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10008691&PLACEBO=IE.pdf.

Armano:1999:CMK

[ACF99] Giuliano Armano, Antonio Corradi, and Valeria Filippi. A communication micro-kernel for implementing object-oriented programming on closely coupled distributed architectures. Concurrency: Practice and Experience, 11(1):43–68, January 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=40006535;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=40006535&PLACEBO=IE.pdf.

Ahmad:1995:TAU

[AD95] I. Ahmad and M. K. Dhodhi. Task assignment using a problem-space genetic algorithm. Concurrency: Practice and Experience, 7(5):411– 428, August 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Adamo:1991:EHC

[Ada91] J. M. Adamo. Exception handling for a communicatingsequential-processes-based extension of C++. Concurrency: Practice and Experience, 3(1): 15–42, February 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Averbuch:1996:PPF

[ADG96] A. Averbuch, R. Dekel, and E. Gabber. Portable parallelizing Fortran compiler. Concurrency: Practice and Experience, 8(2):91–123, March 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=23268.

Akarsu:1998:DJB

[AFH98] Erol Akarsu, Geoffrey Fox, and Tomasz Haupt. DARP: Java-based data analysis and rapid prototyping environment for distributed high performance computations. Concurrency: Practice and Experience, 10(11–13):1097–1105, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050424;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050424&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Averbuch:1998:PPF

[AG98] Amir Averbuch and Eran Gabber. Portable parallel FFT for MIMD multiprocessors. Concurrency: Practice and Experience, 10(8):583–605, July 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10008695;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10008695&PLACEBO=IE.pdf.

Atwood:1997:PCF

[AGD97] Christopher A. Atwood, Rajat P. Garg, and Dennis Deryke. A prototype computational fluid dynamics case study in Java. Concurrency: Practice and Experience, 9(11):1311–1318, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13842;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13842&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Ahmad:1994:PML

[AGMM94] I. Ahmad, A. Ghafoor, K. Mehrotra, and C. K. Mohan. Performance modeling of loadbalancing algorithms using neural networks. Concurrency: Practice and Experience, 6(5): 393–409, August 1994. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Abu-Ghazaleh:1997:VIS

[AGW97] Nael B. Abu-Ghazaleh and Philip A. Wilsey. Variable instruction scheduling for MIMD interpretation on pipelined SIMD machines and for compositional instruction sets. Concurrency: Practice and Experience, 9(1):21–39, January 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13804;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13804&PLACEBO=IE.pdf.

Ariyawansa:1991:PBP

[AH91] K. A. Ariyawansa and D. D.

Hudson. Performance of a benchmark parallel implementation of the Van Slkye and Wets algorithm for two-stage stochastic programs on the Sequent/Balance. Concurrency: Practice and Experience, 3(2): 109–128, April 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Ahmad:1995:RMPb

[Ahm95a] I. Ahmad. Resource management in parallel and distributed systems with dynamic scheduling: dynamic scheduling. Concurrency: Practice and Experience, 7(7):587–590, October 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Ahmad:1995:RMPa

[Ahm95b] I. Ahmad. Resource management of parallel and distributed systems with static scheduling: challenges, solutions and new problems. Concurrency: Practice and Experience, 7(5): 339–347, August 1995. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Arbab:1993:OMI

[AHS93] F. Arbab, I. Herman, and P. Spilling. An overview of Manifold and its implementation. Concurrency: Practice and Experience, 5(1):23–70, February 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Alexandrov:1997:SRI

[AISS97] Albert D. Alexandrov, Maximilian Ibel, Klaus E. Schauser, and Chris J. Scheiman. Super-Web: research issues in Javabased global computing. Concurrency: Practice and Experience, 9(6):535–553, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13876;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13876&PLACEBO=IE.pdf.

Abrams:1995:BSP

[ALCG95] M. Abrams, T. J. Lee, H. T.

Cadiz, and K. Ganugapati. Beyond software performance visualization. Concurrency: Practice and Experience, 7(8): 737–764, December 1995. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Aloisio:1991:TAC

[ALV^{+}91] G. Aloisio, E. Lopinto, N. Veneziani, G. C. Fox, and J. S. Kim. Two approaches to the concurrent implementation of the prime factor algorithm on a hypercube. Concurrency: Practice and Experience, 3(5):483– 495, October 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Atallah:1991:SCC

[AM91] M. J. Atallah and S. McFaddin.

Sequence comparison on the Connection Machine. Concurrency: Practice and Experience, 3(2):89–107, April 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Al-Mulhem:2000:CPV

[AM00a] Muhammed S. Al-Mulhem.

Concurrent programming in VISO. Concurrency: Practice and Experience, 12(5): 281–288, April 25, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72509233/

START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72509233&PLACEBO=IE.pdf.

Andresen:2000:HSH

[AM00b] Daniel Andresen and Timothy McCune. H-SWEB: a hierarchical scheduling system for distributed WWW server clusters. Concurrency: Practice and Experience, 12(2-3): 189–210, February–March 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72504941/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72504941&PLACEBO=IE.pdf.

Almasi:1993:PDS

[AMB^{+}93] G. S. Almasi, T. McLucki, J. Bell, A. Gordon, and D. Hale. Parallel distributed seismic migration. Concurrency: Practice and Experience, 5(2):105–131, April 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Anagnostou:1991:NSI

[AMP^{+}91] G. Anagnostou, Y. Maday, A. T. Patera, P. F. Fischer, and E. M. Renquist. Numerical simulation of incompressible fluid flows. Concurrency: Practice and Experience, 3(6): 667–685, December 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Albrizio:1995:POP

[AMVA95] R. Albrizio, A. Mazzone, N. Veneziani, and G. Aloisio. Pipeline optimizations for the prime factor algorithm. Concurrency: Practice and Experience, 7(1):29–41, February 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Azevedo:2000:AAJ

[ANH00] Ana Azevedo, Alex Nicolau, and Joe Hummel. An annotation-aware Java virtual machine implementation. Concurrency: Practice and Experience, 12(6):423–444, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72515731/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72515731&PLACEBO=IE.pdf.

Anonymous:1995:AI

[Ano95] Anonymous. Author index.

Concurrency: Practice and Experience, 7(8):823, December 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Anonymous:1996:AI

[Ano96] Anonymous. Author index.

Concurrency: Practice and Experience, 8(10):813, December 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Anonymous:1997:AI

[Ano97] Anonymous. Author index. Concurrency: Practice and Experience, 9(12):1407–1408, December 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Anonymous:1998:AI

[Ano98] Anonymous. Author index. Concurrency: Practice and Experience, 10(15):1343–1344, December 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Ayani:1994:PSB

[AR94] R. Ayani and H. Rajaei. Parallel simulation based on conservative time windows: a performance study. Concurrency: Practice and Experience, 6(2): 119–142, April 1994. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Acharya:1997:RAM

[ARS97] Anurag Acharya, M. Ranganathan, and Joel Saltz. Resource-aware metacomputing. Concurrency: Practice and Experience, 9(6):649–674, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13881;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13881&PLACEBO=IE.pdf.

Averbuch:1999:STM

[ARS99] A. Averbuch, Y. Roditty, and B. Shoham. Scheduling tasks of multi-join queries in a multiprocessor. Concurrency: Practice and Experience, 11(5):247–279, April 25, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=61004076;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=61004076&PLACEBO=IE.pdf.

Averbuch:2000:PAH

[ARS00] A. Averbuch, Y. Roditty, and B. Shoham. Performance analysis of hybrid network multiprocessor architecture. Concurrency: Practice and Experience, 12(9):821–844, August 10, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72516483/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516483&PLACEBO=IE.pdf.

Arbenz:1995:STL

[ASLV95] P. Arbenz, C. Sprenger, H. P.

Luethi, and S. Vogel. SCID-DLE: A tool for large scale distributed computing. Concurrency: Practice and Experience, 7(2):121–146, April 1995. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Austin:1997:MBC

[AW97] W. J. Austin and A. M. Wallace. Message based cooperation between parallel depth and intensity matching algorithms. Concurrency: Practice and Experience, 9(2):141–162, February 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13853;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13853&PLACEBO=IE.pdf.

Ben-Ari:1998:HSS

[BA98] Mordechai Ben-Ari. How to solve the Santa Claus problem. Concurrency: Practice and Experience, 10(6):485–496, May 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=5390;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5390&PLACEBO=IE.pdf.

Bozyigit:1999:LBD

[BAGGB99] M. Bozyigit, J. Al-Ghamdi, M. Ghouseuddin, and H. Barada. A load balanced distributed computing system. Concurrency: Practice and Experience, 11(12):753–771, October 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/66002641/
START;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=66002641&PLACEBO=IE.pdf.

Baker:1996:E

[Bak96] Mark Baker. Editorial. Concurrency: Practice and Experience, 8(10):729–730, December 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23266.

Baker:1999:EAD

[Bak99a] Mark Baker. Editorial: Applications of distributed computing environments. Concurrency: Practice and Experience, 11(4):167–168, April 10, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=61004079;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=61004079&PLACEBO=IE.pdf.

Special Issue: Applications of Distributed Computing Environments.

Baker:1999:ECI

[Bak99b] Mark Baker. Editorial: commercial and industrial applications: part 2. Concurrency: Practice and Experience, 11(2):69, February 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=61004055;
http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=61004055&PLACEBO=IE.pdf.

Baker:1999:ESI

[Bak99c] Mark Baker. Editorial: Special issue: Aspects of seamless computing. Concurrency: Practice and Experience, 11(15):885, December 25, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/71005729/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=71005729&PLACEBO=IE.pdf.

Baker:2000:SIM

[Bak00]  Mark Baker.  Special is-

sue: Message passing interfacebased parallel programming with Java. Concurrency: Practice and Experience, 12(11): 1017–1018, September 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76000190/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76000190&PLACEBO=IE.pdf.

Bal:1992:FTP

[Bal92] H. E. Bal. Fault tolerant parallel programming in Argus. Concurrency: Practice and Experience, 4(1):37–55, February 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Battiti:1991:RMV

[Bat91] R. Battiti. Real-time multiscale vision on multi-computers. Concurrency: Practice and Experience, 3(2):55–87, April 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Brunschen:2000:OCP

[BB00] Christian Brunschen and Mats Brorsson. OdinMP/CCp — a portable implementation of OpenMP for C. Concurrency: Practice and Experience, 12(12):1193–1203, October 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/76500347/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500347&PLACEBO=IE.pdf.

Bhatia:1997:WVP

[BBC^{+}97] Dimple Bhatia, Vanco Burzevski, Maja Camuseva, Geoffrey Fox, Wojtek Furmanski, and Girish Premchandran. WebFlow — a visual programming paradigm for Web/Java based coarse grain distributed computing. Concurrency: Practice and Experience, 9(6):555–577, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13877;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13877&PLACEBO=IE.pdf.

Baillie:1991:CIA

[BC91] C. F. Baillie and P. D. Coddington. Cluster identification algorithms for spin models — sequential and parallel. Concurrency: Practice and Experience, 3(2):129–144, April 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Brorsson:2000:SIE

[BC00] Mats Brorsson and Barbara Chapman. Special issue: EWOMP’99 — First European Workshop on OpenMP. Concurrency: Practice and Experience, 12(12):1117–1119, October 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/76500352/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76500352&PLACEBO=IE.pdf.

Burdette:2000:DBT

[BCB00] Steven Burdette, Tracy Camp, and Bill Bynum. Distributed BACI: a toolkit for distributed applications. Concurrency: Practice and Experience, 12(1):35–52, January 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/72504927/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72504927&PLACEBO=IE.pdf.

Brindle:1999:EPP

[BCE^{+}99] Stephen Brindle, Adrian Colbrook, John Elliott, Daron Green, and Francis Wray. The European Porting Project — demonstrating industrial relevance of parallel HPC. Concurrency: Practice and Experience, 11(2):93–108, February 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=61004053;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=61004053&PLACEBO=IE.pdf.

Beca:1997:JEC

[BCF^{+}97] Lukasz Beca, Gang Cheng, Geoffrey C. Fox, Tomasz Jurga, Konrad Olszewski, Marek Podgorny, Piotr Sokolowski, and Krzysztof Walczak. Java enabling collaborative education, health care, and computing. Concurrency: Practice and Experience, 9(6):521–533, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13875;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13875&PLACEBO=IE.pdf.

Becker:1993:PSH

[BD93] J. C. Becker and L. Dagum.

Particle simulation on heterogeneous distributed supercomputers. Concurrency: Practice and Experience, 5(4):367–377, June 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Bartoli:2000:SAS

[BDL00] Alberto Bartoli, Gianluca Dini, and Lanfranco Lopriore. Single address space implementation in distributed systems. Concurrency: Practice and Experience, 12(4):251–280, April 10, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72504932/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72504932&PLACEBO=IE.pdf.

Bacci:1995:PSH

[BDOP95] B. Bacci, M. Danelutto, S. Orlando, and S. Pelagatti. P^{3}L: a structured high-level parallel language, and its structured support. Concurrency: Practice and Experience, 7(3):225–255, May 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Boisvert:1998:DNL

[BDP^{+}98] Ronald F. Boisvert, Jack J. Dongarra, Roldan Pozo, Karin A. Remington, and G. W. Stewart. Developing numerical libraries in Java. Concurrency: Practice and Experience, 10(11–13):1117–1129, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.

com/cgi-bin/abstract?ID=10050395;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050395&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Breg:1998:JRP

[BDV^{+}98] Fabian Breg, Shridhar Diwan, Juan Villacis, Jayashree Balasubramanian, Esra Akman, and Dennis Gannon. Java RMI performance and object model interoperability: experiments with Java/HPC++. Concurrency: Practice and Experience, 10(11–13):941–955, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050400;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050400&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Benkner:2000:OIH

[Ben00] Siegfried Benkner. Optimizing irregular HPF applications using halos. Concurrency: Practice and Experience, 12 (2-3):137–155, February–March 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72504938/
START;   http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?
                                        ID=72504938&PLACEBO=IE.pdf.

Berggren:1994:SAP

[Ber94]  M. Berggren.  Solving an

advection-diffusion problem on the Connection Machine. Concurrency: Practice and Experience, 6(1):55–68, February 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Brochard:1992:CDM

[BF92a] L. Brochard and A. Freau. Computation and data movement on RP3. Concurrency: Practice and Experience, 4(1): 57–78, February 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Brochard:1992:DAR

[BF92b] L. Brochard and A. Freau. Designing algorithms on RP3. Concurrency: Practice and Experience, 4(1):79–106, February 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Baraglia:1997:EWA

[BFFL97] R. Baraglia, G. Faieta, M. Formica, and D. Laforenza. Experiences with a wide area network metacomputing management tool using IBM SP-2 parallel systems. Concurrency: Practice and Experience, 9(3):223–239, March 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.

com/cgi-bin/abstract?ID=13858;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13858&PLACEBO=IE.pdf.

Bik:1997:AEI

[BG97a] Aart J. C. Bik and Dennis B.

Gannon. Automatically exploiting implicit parallelism in Java. Concurrency: Practice and Experience, 9(6):579–619, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13878;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13878&PLACEBO=IE.pdf.

Bik:1997:NNL

[BG97b] Aart J. C. Bik and Dennis B.

Gannon. A note on native level 1 BLAS in Java. Concurrency: Practice and Experience, 9(11):1091–1099, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13826;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13826&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Bik:1998:PBP

[BG98] Aart J. C. Bik and Dennis B.

Gannon. A prototype bytecode parallelization tool. Concurrency: Practice and Experience, 10(11–13):879–885, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050417;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050417&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Brune:1999:MCG

[BGKR99] Matthias Brune, Jörn Gehring, Axel Keller, and Alexander Reinefeld. Managing clusters of geographically distributed highperformance computers. Concurrency: Practice and Experience, 11(15):887–911, December 25, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/71005733/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=71005733&PLACEBO=IE.pdf.

Barak:1992:MTS

[BGL92] A. Barak, S. Guday, and R. Laor. The MPE toolkit for supporting distributed applications. Concurrency: Practice and Experience, 4(6):459– 480, September 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Bednorz:1997:CDA

[BGZ97] M. Bednorz, A. Gwozdowski, and K. Zieliński. Contextual debugging and analysis of multithreaded applications. Concurrency: Practice and Experience, 9(2):123–139, February 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13852;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13852&PLACEBO=IE.pdf.

Berrendorf:1992:EBP

[BH92] R. Berrendorf and J. Helin.

Evaluating the basic performance of the Intel iPSC/860 parallel computer. Concurrency: Practice and Experience, 4(3):223–240, May 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Bernaschi:1998:CCO

[BI98] Massimo Bernaschi and Giulio Iannello. Collective communication operations: experimental results vs. theory. Concurrency: Practice and Experience, 10(5):359–386, April 25, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=5383;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5383&PLACEBO=IE.pdf.

Budimlic:1997:OJT

[BK97] Zoran Budimlic and Ken Kennedy. Optimizing Java: theory and practice. Concurrency: Practice and Experience, 9(6):445–463, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13870;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13870&PLACEBO=IE.pdf.

Baratloo:1998:INC

[BKKK98] Arash Baratloo, Mehmet Karaul, Holger Karl, and Zvi M. Kedem. An infrastructure for network computing with Java applets. Concurrency: Practice and Experience, 10 (11–13):1029–1041, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050421;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050421&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Bal:1992:RTS

[BKTJ92] H. E. Bal, M. F. Kaashoek, A. S. Tanenbaum, and J. Jansen. Replication techniques for speeding up parallel applications on distributed systems. Concurrency: Practice and Experience, 4(5):337–355, August 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Beton:1991:TRE

[BKU91] R. Beton, J. Kingdon, and C. Upstill. A transputer radar ESM data processor. Concurrency: Practice and Experience, 3(4):303–313, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Bozyigit:1997:PJI

[BMAT97] M. Bozyigit, S. A. Mohammed, and M. Al-Tayyeb. Parallel join for IBGF partitioned relational databases. Concurrency: Practice and Experience, 9(8):821–836, August 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13891;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13891&PLACEBO=IE.pdf.

Boloni:2000:ABS

[BMR^{+}00] L. Bölöni, D. C. Marinescu, J. R. Rice, P. Tsompanopoulou, and E. A. Vavalis. Agent based scientific simulation and modeling. Concurrency: Practice and Experience, 12(9): 845–861, August 10, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516484/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516484&PLACEBO=IE.pdf.

Baxter:1996:EPF

[BMT96] Robert M. Baxter, Killian D. Murphy, and Shari M. Trewin. Experiences in parallelising FLITE3D on the Cray T3D. Concurrency: Practice and Experience, 8(10):741–755, December 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23263.

Berrendorf:2000:PCO

[BN00] Rudolf Berrendorf and Guido Nieken. Performance characteristics for OpenMP constructs on different parallel computer architectures. Concurrency: Practice and Experience, 12(12):1261–1273, October 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/76500355/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76500355&PLACEBO=IE.pdf.

Boillat:1990:LBP

[Boi90] J. E. Boillat. Load balancing and Poisson equation in a graph. Concurrency: Practice and Experience, 2(4):289– 314, December 1990. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Beavis:1992:PDD

[BP92] Andy Beavis and Chris Phillips. Porting a dusty deck Fortran program to a shared-memory multiprocessor. Concurrency: Practice and Experience, 4(8): 575–587, December 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Bomans:1989:BIH

[BR89] L. Bomans and D. Roose.

Benchmarking the iPSC/2 hypercube multiprocessor. Concurrency: Practice and Experience, 1(1):3–18, September 1989. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Binu:2000:ITP

[BR00] K. J. Binu and D. Janaki Ram. Integrating task parallelism in data parallel languages for parallel programming on NOWs. Concurrency: Practice and Experience, 12(13):1291–1315, November 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76500625/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500625&PLACEBO=IE.pdf.

Brieger:2000:HOO

[Bri00] Leesa Brieger. HPF to OpenMP on the Origin2000: a case study. Concurrency: Practice and Experience, 12(12):1147–1154, October 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.

                                        com/cgi-bin/abstract/76500351/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76500351&PLACEBO=IE.pdf.

Betello:1993:LBM

[BRSR93] G. Betello, G. Richelli, S. Succi, and F. Ruello. Lattice Boltzmann method on a cluster of IBM RISC System/6000 workstations. Concurrency: Practice and Experience, 5(4):359–366, June 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Biswas:1994:PBN

[BRTK94] P. Biswas, K. K. Ramakrishnan, D. Towsley, and C. M. Krishna. Performance benefits of non-volatile caches in distributed file systems. Concurrency: Practice and Experience, 6(4):289–323, June 1994. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Brugnano:1992:TTP

[Bru92] L. Brugnano. Two-level twisted preconditionings for parallel computers. Concurrency: Practice and Experience, 4(2):107– 117, April 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Barnard:1994:FMI

[BS94a] S. T. Barnard and H. D. Simon. Fast multilevel implementation of recursive spectral bisection for partitioning unstructured problems. Concurrency: Practice and Experience, 6(2):101–117, April 1994. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Braun:1994:PTS

[BS94b] T. Braun and C. Schmidt. Parallel transport subsystem implementation for high-performance communication. Concurrency: Practice and Experience, 6(4): 375–391, June 1994. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Barbier:1994:PDP

[BSD94] C. Barbier, P. Sen, and M. Downie. Parallel dynamic programming and ship voyage management. Concurrency: Practice and Experience, 6(8): 673–696, December 1994. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Bastide:2000:FNT

[BSP00] Rémi Bastide, Ousmane Sy, and Philippe Palanque. A formal notation and tool for the engineering of CORBA systems. Concurrency: Practice and Experience, 12(14): 1379–1403, December 10, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76500720/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500720&PLACEBO=IE.pdf.

Berryman:1991:ETS

[BSS91] H. Berryman, J. Saltz, and J. Scroggs. Execution time support for adaptive scientific algorithms on distributed memory machines. Concurrency: Practice and Experience, 3(3): 159–178, June 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Bizarro:1998:JJL

[BSS98] Pedro Bizarro, Luís M. Silva, and João Gabriel Silva. JWarp: a Java library for parallel discrete-event simulations. Concurrency: Practice and Experience, 10(11–13):999– 1005, September 1998. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (elec-

tronic). URL http://www3.
interscience.wiley.com/cgi-
bin/abstract?ID=10050415;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050415&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Bull:2000:BSH

[BSW^{+}00] J. M. Bull, L. A. Smith, M. D. Westhead, D. S. Henty, and R. A. Davey. A benchmark suite for high performance Java. Concurrency: Practice and Experience, 12(6):375–388, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72515724/

START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72515724&PLACEBO=IE.pdf.

Bishop:1993:EMA

[BT93] J. M. Bishop and K. S.

Thomas. Experience with multi-transputer Ada. Concurrency: Practice and Experience, 5(2):133–151, April 1993. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Baritompa:1997:PMB

[BV97] William Baritompa and Sami Viitanen. Parallel multidimensional bisection. Concurrency: Practice and Experience, 9(2):113–121, February 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13851;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13851&PLACEBO=IE.pdf.

Badia:2000:ITE

[BV00] J. M. Badía and A. M. Vidal. Inverse Toeplitz eigenproblem on personal computer networks. Concurrency: Practice and Experience, 12(13): 1275–1290, November 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76500624/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500624&PLACEBO=IE.pdf.

Bik:1997:JPJ

[BVG97] Aart J. C. Bik, Juan E. Villacis, and Dennis B. Gannon. javar: A prototype Java restructuring compiler. Concurrency: Practice and Experience, 9(11):1181–1191, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13819;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13819&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Butcher:1994:GSL

[BWA94] P. Butcher, A. Wood, and M. Atkins. Global synchronisation in Linda. Concurrency: Practice and Experience, 6(6): 505–516, September 1994. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Brehm:1998:PMS

[BWM98] Jürgen Brehm, Patrick H. Worley, and Manish Madhukar. Performance modeling for SPMD messagepassing programs. Concurrency: Practice and Experience, 10(5):333–357, April 25, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=5382;
                                         http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=5382&PLACEBO=IE.pdf.

Coetzee:1996:PNN

[CB96] Louis Coetzee and Elizabeth C.

Botha. Parallel neural net training on the KSR1. Concurrency: Practice and Experience, 8(8):617–638, October 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23300.

Coetzee:1998:PDS

[CB98] Louis Coetzee and Elizabeth C.

Botha. The parallel downhill simplex algorithm for unconstrained optimisation. Concurrency: Practice and Experience, 10(2):121–137, February 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=5369;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5369&PLACEBO=IE.pdf.

Carmona:1997:PPV

[CC97a] Edward A. Carmona and Leon J. Chandler. On parallel PIC versatility and the structure of parallel PIC approaches. Concurrency: Practice and Experience, 9(12): 1377–1405, December 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.

                                        com/cgi-bin/abstract?ID=13848;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13848&PLACEBO=IE.pdf.

Chen:1997:JRD

[CC97b] Marina Chen and James Cowie. Java’s role in distributed collaboration. Concurrency: Practice and Experience, 9(6):509–519, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13874;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13874&PLACEBO=IE.pdf.

Chang:1997:TAS

[CCCL97] Rong-Guey Chang, Cheng-Wei Chen, Tyng-Ruey Chuang, and Jenq Kuen Lee. Towards automatic support of parallel sparse computation in Java with continuous compilation. Concurrency: Practice and Experience, 9(11):1101–1111, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13827;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13827&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Carpenter:1997:EHJ

[CCF^{+}97] Bryan Carpenter, Yuh-Jye Chang, Geoffrey Fox, Donald Leskiw, and Xiaoming Li. Experiments with ‘HP Java’. Concurrency: Practice and Experience, 9(6):633–648, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13880;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13880&PLACEBO=IE.pdf.

Christiansen:1997:JIB

[CCI^{+}97] Bernd O. Christiansen, Peter Cappello, Mihai F. Ionescu, Michael O. Neary, Klaus E. Schauser, and Daniel Wu. Javelin: Internet-based parallel computing using Java. Concurrency: Practice and Experience, 9(11):1139–1160, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13833;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13833&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Clarke:1999:UAL

[CD99] Sheldon Clarke and Sivarama P.

Dandamudi. Usefulness of adaptive load sharing for parallel processing on networks of workstations. Concurrency: Practice and Experience, 11(8):387–405, July 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=63000995;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=63000995&PLACEBO=IE.pdf.

Casanova:1997:JAN

[CDD97] Henri Casanova, Jack Dongarra, and David M. Doolin. Java access to numerical libraries. Concurrency: Practice and Experience, 9(11): 1279–1291, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13823;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13823&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Chaves:1993:KCS

[CDL^{+}93] E. M. Chaves, Jr., P. C. Das, T. J. Leblanc, B. D. Marsh, and M. L. Scott. Kernel-kernel communication in a shared-memory multiprocessor. Concurrency: Practice and Experience, 5(3): 171–191, May 1993. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Clayton:1994:DUL

[CDMW94] S. Clayton, R. J. Duckworth, W. Michalson, and A. Wilson. Determining update latency bounds in Galactica Net.

Concurrency: Practice and Experience, 6(7):595–611, October 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Calland:1999:TSC

[CDR99] Pierre-Yves Calland, Jack Dongarra, and Yves Robert. Tiling on systems with communication/computation overlap. Concurrency: Practice and Experience, 11(3):139–153, March 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=61003665;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=61003665&PLACEBO=IE.pdf.

Choi:1994:PPU

[CDW94] Jaeyoung Choi, Jack J. Dongarra, and David W. Walker. PUMMA: Parallel Universal Matrix Multiplication Algorithms on distributed memory concurrent computers. Concurrency: Practice and Experience, 6(7):543–570, October 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http://

www.netlib.org/utk/people/
JackDongarra/PAPERS/PUMMA-
Parallel-Universal-Matrix-
Multiplication-Algorithms.
pdf.

Choi:1996:PBS

[CDW96] Jaeyoung Choi, Jack J. Dongarra, and David W. Walker.

PB-BLAS: a set of parallel block basic linear algebra subprograms. Concurrency: Practice and Experience, 8(7):517–535, September 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23295.

Chen:1995:FRM

[CE95] S. Chen and M. M. Eshaghian. A fast recursive mapping algorithm. Concurrency: Practice and Experience, 7(5):391– 409, August 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Czarnecki:2000:SO

[CE00] Krzysztof Czarnecki and Ulrich W. Eisenecker. Synthesizing objects. Concurrency: Practice and Experience, 12(14):1347–1377, December 10, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/76500719/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500719&PLACEBO=IE.pdf.

Cheng:1996:IMP

[CF96] Gang Cheng and Geoffrey C. Fox. Integrating multiple parallel programming paradigms in a dataflow-based software environment. Concurrency: Practice and Experience, 8 (9):667–684, November 1996.

CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=23303.

Carpenter:2000:OSM

[CFKL00] Bryan Carpenter, Geoffrey Fox, Sung Hoon Ko, and Sang Lim. Object serialization for marshaling data in a Java interface to MPI. Concurrency: Practice and Experience, 12(7):539–553, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72516217/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516217&PLACEBO=IE.pdf.

Cheng:1999:CFI

[CFLH99] Gang Cheng, Geoffrey C. Fox, Tseng-Hui Lin, and Tomasz Haupt. A computing framework for integrating interactive visualization in HPCC applications. Concurrency: Practice and Experience, 11(2):71–92, February 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=61004056;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=61004056&PLACEBO=IE.pdf.

Chong:1991:BRR

[CFMP91] M. Chong, F. Fallside, T. Marsland, and R. Prager. ‘Blade Runner’: a real-time speech recognizer. Concurrency: Practice and Experience, 3(3):145–157, June 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Carpenter:2000:MML

[CGJ^{+}00] Bryan Carpenter, Vladimir Getov, Glenn Judd, Anthony Skjellum, and Geoffrey Fox. MPJ: MPI-like message passing for Java. Concurrency: Practice and Experience, 12(11): 1019–1038, September 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76000188/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76000188&PLACEBO=IE.pdf.

Chtchelkanova:1997:PIB

[CGM^{+}97] Almadena Chtchelkanova, John Gunnels, Greg Morrow, James Overfelt, and Robert A. van de Geijn. Parallel implementation of BLAS: general techniques for Level 3 BLAS. Concurrency: Practice and Experience, 9(9): 837–857, September 1997. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (elec-

tronic). URL http://www.
cs.utexas.edu/users/rvdg/
papers/sB_BLAS.ps;   http:
//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13893;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13893&PLACEBO=IE.pdf.

Claver:1999:PAW

[CH99] Jose M. Claver and Vicente Hernandez. Parallel adaptive wavefront algorithms solving Lyapunov equations for the Cholesky factor on message passing multiprocessors. Concurrency: Practice and Experience, 11(14):849–862, December 10, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/67501483/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=67501483&PLACEBO=IE.pdf.

Cai:1998:DEB

[CHB98] F. F. Cai, M. E. C. Hull, and D. A. Bell. The design and evaluation of a buffer algorithm for database machines. Concurrency: Practice and Experience, 10(15):1291–1316, December 25, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=40005938;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=40005938&PLACEBO=IE.pdf.

Cornea-Hasegan:1994:DMC

[CHMZ94] M. C. Cornea-Hasegan, D. C.

Marinescu, and Z. Zhang. Data management for a class of iterative computations on distributed-memory MIMD systems. Concurrency: Practice and Experience, 6(3):205–229, May 1994. CODEN CPEXEI.

ISSN 1040-3108 (print), 1096-

9128 (electronic).

Choi:1998:NPM

[Cho98] Jaeyoung Choi. A new parallel matrix multiplication algorithm on distributed-memory concurrent computers. Concurrency: Practice and Experience, 10(8):655–670, July 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10008698;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10008698&PLACEBO=IE.pdf.

Christian:1991:VDS

[Chr91] S. Christian. VLSI design stations using transputers. Concurrency: Practice and Experience, 3(4):435–443, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Chatterjee:1991:EOD

[CKH91] A. Chatterjee, A. Khanna, and Y. Hung. ES-Kit: an objectoriented distributed system. Concurrency: Practice and Experience, 3(6):525–539, December 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Caromel:1998:TSC

[CKV98] Denis Caromel, Wilfried Klauser, and Julien Vayssière. Towards seamless computing and metacomputing in Java. Concurrency: Practice and Experience, 10(11–13):1043–1061, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050422;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050422&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Cierniak:1997:JTO

[CL97a] Michał Cierniak and Wei Li.

Just-in-time optimizations for high-performance Java programs. Concurrency: Practice and Experience, 9(11): 1063–1073, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13828;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13828&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Cierniak:1997:OJB

[CL97b] Michał Cierniak and Wei Li.

Optimizing Java bytecodes. Concurrency: Practice and Experience, 9(6):427–444, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13869;

                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13869&PLACEBO=IE.pdf.

Clausen:1997:JBO

[Cla97] Lars R. Clausen. A Java bytecode optimizer using sideeffect analysis. Concurrency: Practice and Experience, 9(11): 1031–1045, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13838;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13838&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Cignoni:1995:EPS

[CLMP95] P. Cignoni, D. Laforenza, C. Montani, and R. Perego. Evaluation of parallelization strategies for an incremental Delaunay Triangulator in E^{3}. Concurrency: Practice and Experience, 7(1):61–80, February 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Chrisochoides:1997:COH

[CMF97] Nikos Chrisochoides, Nashat Mansour, and Geoffrey Fox. A comparison of optimization heuristics for the data mapping problem. Concurrency: Practice and Experience, 9(5):319–343, May 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13864;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13864&PLACEBO=IE.pdf.

Chen:1997:WBF

[CMM^{+}97] Zhikai Chen, Kurt Maly, Piyush Mehrotra, Praveen K. Vangala, and Mohammad Zubair. Web-based framework for distributed computing. Concurrency: Practice and Experience, 9(11):1175–1180, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13839;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13839&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Corana:1998:CCD

[Cor98] Angelo Corana. Computing the correlation dimension on a network of workstations. Concurrency: Practice and Experience, 10(10):737–762, August 25, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10006461;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10006461&PLACEBO=IE.pdf.

Cowie:1998:JPD

[Cow98] James Cowie. JTED: parallel discrete-event simulation in Java. Concurrency: Practice and Experience, 10(11– 13):993–997, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050414;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050414&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Choudhary:1992:RDD

[CP92] A. N. Choudhary and R. Ponnusamy. Run-time data decomposition for parallel implementation of image processing and computer vision tasks. Concurrency: Practice and Experience, 4(4):313–334, June 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Candlin:1995:DBP

[CP95] R. Candlin and J. Phillips. The dynamic behaviour of parallel programs under process migration. Concurrency: Practice and Experience, 7(7):591– 613, October 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Crooks:1998:ADD

[CP98] P. Crooks and R. H. Perrott. An automatic data distribution generator for distributed memory machines. Concurrency: Practice and Experience, 10(8):607–629, July 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10008696;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10008696&PLACEBO=IE.pdf.

Coen-Porisini:2000:UTD

[CPM00] Alberto Coen-Porisini and Dino Mandrioli. Using TRIO for designing a CORBA-based application. Concurrency: Practice and Experience, 12(10): 981–1015, August 25, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/73000574/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73000574&PLACEBO=IE.pdf.

Ceccarelli:1993:CNN

[CPV93] M. Ceccarelli, A. Petrosino, and R. Vaccaro. Competitive neural networks on message-passing parallel computers. Concurrency: Practice and Experience, 5(6):449–470, September 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Crandall:1995:PAS

[CQ95] P. E. Crandall and M. J. Quinn.

A partitioning advisory system for networked data-parallel processing. Concurrency: Practice and Experience, 7(5):479– 495, August 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Chuang:1994:EPU

[CRM94] L.-Y. Chuang, V. Rego, and A. Mathur. Experiments with program unification on the Cray Y-MP. Concurrency: Practice and Experience, 6(1):33– 53, February 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Chandy:1998:UAL

[CRS98] K. Mani Chandy, Adam Rifkin, and Eve Schooler. Using announce-listen with global events to develop distributed control systems. Concurrency: Practice and Experience, 10(11–13):1021–1027, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050420;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050420&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Chang:1995:CSM

[CS95] C.-Y. Chang and J.-P. Sheu. Compile-time scheduling of multithread with data localities on multiple vector processors. Concurrency: Practice and Experience, 7(5):349–369, August 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Cierniak:1997:PBP

[CS97] Michał Cierniak and Suresh Srinivas. A portable browser for performance programming. Concurrency: Practice and Experience, 9(11):1243–1248, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

URL http://www3.interscience.
wiley.com/cgi-bin/abstract?
ID=13815;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13815&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Chrobot:1998:AIP

[CSS98] Stanislaw Chrobot, Agata Stras, and Robert Stras. ARS: an integrated package for shared-variable and messagepassing computations in objectbased systems. Concurrency: Practice and Experience, 10 (15):1271–1290, December 25, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=40005937;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=40005937&PLACEBO=IE.pdf.

Cannataro:1992:DIE

[CSTG92] M. Cannataro, G. Spezzano, D. Talia, and E. Gallizzi. Design, implementation and evaluation of a deadlock-free routing algorithm for concurrent computers. Concurrency: Practice and Experience, 4(2):143–161, April 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Chen:2000:PPM

[CT00] Jian Chen and Valerie E. Taylor. ParaPART: parallel mesh partitioning tool for distributed systems. Concurrency: Practice and Experience, 12 (2-3):111–123, February–March 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72504934/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72504934&PLACEBO=IE.pdf.

Cunha:1997:BSB

[Cun97] Rudnei Dias Da Cunha. A benchmark study based on the parallel computation of the vector outer-product A = uv^{T} operation. Concurrency: Practice and Experience, 9(8):803–819, August 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13890;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13890&PLACEBO=IE.pdf.

Chang:2000:JJI

[CvE00] Chi-Chao Chang and Thorsten von Eicken. Javia: A Java interface to the virtual interface architecture. Concurrency: Practice and Experience, 12(7):573–593, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516214/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516214&PLACEBO=IE.pdf.

Clarke:1991:TER

[CW91] L. Clarke and G. Wilson. Tiny:

an efficient routing harness for the Inmos transputer. Concurrency: Practice and Experience, 3(3):221–245, June 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Clint:1994:CPF

[CWB94] M. Clint, J. S. Weston, and C. W. Bleakney. Comparison of parallel Fortran environments on the AMT DAP510 for a linear algebra application. Concurrency: Practice and Experience, 6(3):193–204, May 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Chung:1995:PDL

[CYL95] Y.-C. Chung, Y.-J. Yeh, and J.-S Liu. A parallel dynamic load-balancing algorithm for solution-adaptive finite element meshes on 2D tori. Concurrency: Practice and Experience, 7(7):615–631, October 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Carpenter:1998:HDP

[CZF^{+}98] Bryan Carpenter, Guansong Zhang, Geoffrey Fox, Xinying Li, and Yuhong Wen. HPJava: data parallel extensions to Java. Concurrency: Practice and Experience, 10(11–13):873–877, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050411;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050411&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Driggers:1997:DCE

[DAB97] Brent Driggers, Jay Alameda, and Ken Bishop. Distributed collaboration for engineering and scientific applications implemented in Habanero, a Javabased environment. Concurrency: Practice and Experience, 9(11):1269–1277, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13841;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?

ID=13841&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Dagum:1992:DPS

[Dag92] L. Dagum. Data parallel sorting for particle simulation. Concurrency: Practice and Experience, 4(3):241–255, May 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

David:1997:KAI

[Dav97] Olaf David. A kernel approach for interactive-oriented model construction in Java. Concurrency: Practice and Experience, 9(11):1319–1326, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13836;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13836&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Dincer:1996:BCC

[DBRF96] Kivanc Dincer, Zeki Bozkus, Sanjay Ranka, and Geoffrey Fox. Benchmarking the computation and communication performance of the CM-5. Concurrency: Practice and Experience, 8(1):47–69, January 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23258.

Davis:1992:PAS

[DCGC92] N. J. Davis IV, B. A. Carpenter, C. W. Glover, and J. C. Culioli. Parallel approaches to the solution of the assignment problem. Concurrency: Practice and Experience, 4(2):163–184, April 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

DApuzzo:1994:PBR

[DD94] M. D’Apuzzo and M. A. De Rosa. A parallel block rowaction method for solving large sparse linear systems on distributed memory multiprocessors. Concurrency: Practice and Experience, 6(1):69– 84, February 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Dongarra:1997:MPP

[DD97] Jack J. Dongarra and Tom Dunigan. Message-passing performance of various computers. Concurrency: Practice and Experience, 9(10):915–926, October 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13808;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13808&PLACEBO=IE.pdf.

DAzevedo:2000:DIP

[DD00] Eduardo D’Azevedo and Jack Dongarra. The design and implementation of the parallel out-of-core ScaLAPACK LU, QR, and Cholesky factorization routines. Concurrency: Practice and Experience, 12(15):1481–1493, December 25, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/76505648/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76505648&PLACEBO=IE.pdf.

Demmel:1991:LPL

[Dem91]  J. Demmel.  LAPACK: A

portable linear algebra library for high-performance computers. Concurrency: Practice and Experience, 3(6):655– 666, December 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Demaine:1998:CJC

[Dem98] Erik D. Demaine. C to Java:

converting pointers into references. Concurrency: Practice and Experience, 10(11– 13):851–861, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050394;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050394&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Dincer:1997:UJJ

[DF97] Kivanc Dincer and Geoffrey C. Fox. Using Java and JavaScript in the Virtual Programming Laboratory: a Web-based parallel programming environment. Concurrency: Practice and Experience, 9(6):485–508, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13873;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13873&PLACEBO=IE.pdf.

Debbage:1993:GCL

[DHN93] M. Debbage, M. B. Hill, and D. A. Nicole. Global communications on locally connected message-passing parallel computers. Concurrency: Practice and Experience, 5(6):491– 509, September 1993. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Donaldson:1999:PCU

[DHS99] Stephen R. Donaldson, Jonathan M. D. Hill, and David B. Skillicorn. Predictable communication on unpredictable networks: implementing BSP over TCP/IP and UDP/IP. Concurrency: Practice and Experience, 11(11):687–700, September 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=64500095;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=64500095&PLACEBO=IE.pdf.

Daniel:1996:EAP

[DM96] J. W. H. Daniel and M. R.

Moulding. Exploiting application parallelism in knowledgebased systems: an experimental method. Concurrency: Practice and Experience, 8(1):1–18, January 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23256.

Dowsing:1998:PSS

[DM98] R. D. Dowsing and W. S.

Martins. The performance of a selection of sorting algorithms on a general purpose parallel computer. Concurrency: Practice and Experience, 10(4):315–332, April 10, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=5380;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5380&PLACEBO=IE.pdf.

Diekmann:1998:PDU

[DMM98] Ralf Diekmann, Derk Meyer, and Burkhard Monien. Parallel decomposition of unstructured FEM-meshes. Concurrency: Practice and Experience, 10(1):53–72, January 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=5365;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5365&PLACEBO=IE.pdf.

diSerafino:1992:FPS

[dMP92] D. di Serafino, A. Murli, and F. Perla. A fast Poisson solver for distributed memory multiprocessors. Concurrency: Practice and Experience, 4(7): 499–508, October 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

deLeone:1993:MPS

[dR93] R. de Leone and M. A. Tork Roth. Massively parallel solution of quadratic programs via successive overrelaxation. Concurrency: Practice and Experience, 5(8):623–634, December 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Diniz:1999:STP

[DR99] Pedro C. Diniz and Martin C. Rinard. Synchronization transformations for parallel computing. Concurrency: Practice and Experience, 11(13):773–802, November 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/67500119/

START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=67500119&PLACEBO=IE.pdf.

Darlington:1990:DLP

[DRW90] J. Darlington, M. Reeve, and S. Wright. Declarative languages and program transformation for programming parallel systems. Concurrency: Practice and Experience, 2(3):149– 170, September 1990. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Day:1998:PID

[DT98] A. M. Day and D. Tracey.

Parallel implementations for determining the 2D convex hull. Concurrency: Practice and Experience, 10(6):449–466, May 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=5388;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5388&PLACEBO=IE.pdf.

Dunford:1991:TIE

[DTY91] C. M. Dunford, J. A. Thompson, and K. H. Yearby. A transputer-based instrument for the ESA/NASA cluster mission. Concurrency: Practice and Experience, 3(4):293–302, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Dunigan:1992:HCS

[Dun92] T. H. Dunigan. Hypercube clock synchronization. Concurrency: Practice and Experience, 4(3):257–268, May 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

DePietro:1994:CSA

[DV94] G. De Pietro and U. Villano. A clock synchronization algorithm for the performance analysis of multicomputer systems. Concurrency: Practice and Experience, 6(8):653–671, December 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Dwyer:1997:FPA

[DW97] Matthew B. Dwyer and Virgil Wallentine. A framework for parallel adaptive grid simulations. Concurrency: Practice and Experience, 9(11):1293–1310, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13824;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13824&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Everaars:2000:DPC

[EAK00] C. T. H. Everaars, F. Arbab, and B. Koren. Dynamic process composition and communication patterns in irregularly structured applications. Concurrency: Practice and Experience, 12(2-3): 157–174, February–March 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72504935/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72504935&PLACEBO=IE.pdf.

Eidson:1995:IFB

[EE95] T. M. Eidson and G. Erlebacher. Implementation of a fully balanced periodic tridiagonal solver on a parallel distributed memory architecture. Concurrency: Practice and Experience, 7(4):273–302, June 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Eigenmann:1993:RFP

[EHJ^{+}93] R. Eigenmann, J. Hoeflinger, G. Jaxon, Z. Li, and D. Padua. Restructuring Fortran programs for Cedar. Concurrency: Practice and Experience, 5(7): 553–573, October 1993. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Egan:1997:JST

[EKR97] M. A. Egan, M. Krishnamoorthy, and K. Rajan. A Java simulation tool for fuzzy clustering. Concurrency: Practice and Experience, 9(11): 1327–1332, November 1997.

CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13837;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13837&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Eppinger:1991:OSM

[Epp91] A. Eppinger. Object-oriented simulations in mechatronics. Concurrency: Practice and Experience, 3(4):387–397, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Erhard:1992:NST

[ES92] W. Erhard and M. Schafer. Numerical simulation of threedimensional thermal convection on the array processor DAP 510. Concurrency: Practice and Experience, 4(1):19–35, February 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Fahringer:1996:EUW

[Fah96] Thomas Fahringer. On estimating the useful work distribution of parallel programs under P^{3} T: a static performance estimator. Concurrency: Practice and Experience, 8(4):261–282, May 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23278.

Fleury:1999:CGD

[FDC99] M. Fleury, A. C. Downton, and A. F. Clark. Constructing generic data-farm templates. Concurrency: Practice and Experience, 11(9):509–528, August 10, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=63001743;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=63001743&PLACEBO=IE.pdf.

Ferrari:1998:JNP

[Fer98] Adam Ferrari. JPVM: network parallel computing in Java. Concurrency: Practice and Experience, 10(11– 13):985–992, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050413;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050413&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Fox:1997:JPC

[FF97] Geoffrey C. Fox and Wojtek Furmanski. Java for parallel computing and as a general language for scientific and engineering simulation and modeling. Concurrency: Practice and Experience, 9(6):415–425, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13868;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13868&PLACEBO=IE.pdf.

Furmanski:1993:MMO

[FFH^{+}93] W. Furmanski, C. Faigle, T. Haupt, J. Niemiec, M. Podgorny, and D. Simoni. MOVIE model for open systems based high-performance distributed computing. Concurrency: Practice and Experience, 5(4):287– 308, June 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Foster:1998:SIW

[FGN^{+}98] Ian Foster, Jonathan Geisler, Bill Nickless, Warren Smith, and Steven Tuecke. Software infrastructure for the I-WAY metacomputing experiment. Concurrency: Practice and Experience, 10(7):567–581, June 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10008693;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10008693&PLACEBO=IE.pdf.

Fleury:1998:PRI

[FHC98] M. Fleury, L. Hayat, and A. F. Clark. Parallel reconfiguration in an image-processing context.

Concurrency: Practice and Experience, 10(4):249–263, April 10, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=5377;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5377&PLACEBO=IE.pdf.

Fidanova:1997:LAS

[Fid97] Stefka Fidanova. Linear array for spelling correction. Concurrency: Practice and Experience, 9(10):967–973, October 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13811;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13811&PLACEBO=IE.pdf.

Fox:1997:PFJ

[FLQZ97] Geoffrey Fox, Xiaoming Li, Zheng Qiang, and Wu Zhigang. A prototype of Fortranto-Java converter. Concurrency: Practice and Experience, 9(11):1047–1061, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13832;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13832&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Fox:1989:PCC

[Fox89] G. C. Fox. Parallel computing comes of age: Supercomputer level parallel computations at Caltech. Concurrency: Practice and Experience, 1(1): 63–103, September 1989. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Fox:1991:APP

[Fox91a]  G. C. Fox.  Achievements

and prospects for parallel computing. Concurrency: Practice and Experience, 3(6):725– 739, December 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Fox:1991:PC

[Fox91b] G. C. Fox. Physical computation. Concurrency: Practice and Experience, 3(6):627– 653, December 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Fox:1997:EJC

[Fox97a]  Geoffrey Fox.  Editorial:

Java for computational science and engineering — simulation and modeling. Concurrency: Practice and Experience, 9(6):413–414, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13882;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13882&PLACEBO=IE.pdf.

Fox:1997:JCS

[Fox97b] Geoffrey Fox. Java for computational science and engineering — simulation and modeling II. Concurrency: Practice and Experience, 9(11): 1001–1002, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13845;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13845&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Fox:1998:EJH

[Fox98] Geoffrey Fox. Editorial: Java for high-performance network computing. Concurrency: Practice and Experience, 10(11– 13):821–824, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050427;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050427&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Fox:2000:ESIa

[Fox00a] Geoffrey Fox. Editorial: Special issue: ACM 1999 Java Grande Conference. Concurrency: Practice and Experience, 12(6):341–342, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72515722/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72515722&PLACEBO=IE.pdf.

Fox:2000:ESIb

[Fox00b] Geoffrey Fox. Editorial: Special issue: ACM 1999 Java Grande Conference. Concurrency: Practice and Experience, 12(7):493–494, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516216/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516216&PLACEBO=IE.pdf.

Fox:2000:ESIc

[Fox00c] Geoffrey Fox. Editorial: Special issue: ACM 1999 Java Grande Conference. Concurrency: Practice and Experience, 12(8):611–612, July 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516225/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516225&PLACEBO=IE.pdf.

Ferrari:1998:MDC

[FS98] Adam Ferrari and V. S. Sunderam. Multiparadigm distributed computing with TPVM. Concurrency: Practice and Experience, 10(3):199–228, March 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=5374;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5374&PLACEBO=IE.pdf.

Fleury:1999:MTS

[FSDC99] M. Fleury, N. Sarvan, A. C.

Downton, and A. F. Clark. Methodology and tools for system analysis of parallel pipelines. Concurrency: Practice and Experience, 11(11): 655–670, September 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=64500090;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=64500090&PLACEBO=IE.pdf.

Foster:1997:TUS

[FTT97] Ian Foster, George K. Thiruvathukal, and Steven Tuecke. Technologies for ubiquitous supercomputing: a Java interface to the Nexus communication system. Concurrency: Practice and Experience, 9(6):465–475, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13871;
http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=13871&PLACEBO=IE.pdf.

Fur:1996:SPP

[Fur96] Marc Le Fur. Scanning parameterized polyhedron using Fourier-Motzkin elimination. Concurrency: Practice and Experience, 8(6):445–460, July 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23287.

Gonzalez:2000:NSF

[GAM^{+}00] Marc Gonzàlez, Eduard Ayguadé, Xavier Martorell, Jesús Labarta, Nacho Navarro, and José Oliver. NanosCompiler: supporting flexible multilevel parallelism exploitation in OpenMP. Concurrency: Practice and Experience, 12(12):1205–1218, October 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/76500358/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500358&PLACEBO=IE.pdf.

Gonzalez:2000:TTP

[GARR00] D. González, F. Almeida, J. Roda, and C. Rodríguez. From the theory to the tools: parallel dynamic programming. Concurrency: Practice and Experience, 12(1):21–34, January 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72504925/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72504925&PLACEBO=IE.pdf.

Gerndt:1990:UDV

[Ger90] M. Gerndt. Updating distributed variables in local computations. Concurrency: Practice and Experience, 2(3):171– 193, September 1990. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Gu:1998:FLM

[GESV98] Weiming Gu, Greg Eisenhauer, Karsten Schwan, and Jeffrey Vetter. Falcon: Online monitoring for steering parallel programs. Concurrency: Practice and Experience, 10(9):699–736, August 10, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10008704;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10008704&PLACEBO=IE.pdf.

Ghosh:1999:CHP

[GFS99] Kaushik Ghosh, Richard M. Fujimoto, and Karsten Schwan. Composing high-performance schedulers: a case study from real-time simulation. Concurrency: Practice and Experience, 11(5):221–245, April 25, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=61004077;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=61004077&PLACEBO=IE.pdf.

Getov:2000:APD

[GGS00] Vladimir S. Getov, Paul A. Gray, and Vaidy S. Sunderam. Aspects of portability and distributed execution for JNIwrapped message passing libraries. Concurrency: Practice and Experience, 12(11): 1039–1050, September 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/76000191/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76000191&PLACEBO=IE.pdf.

Grimstead:1995:ART

[GH95] I. J. Grimstead and S. Hurley. Accelerated ray tracing using an nCUBE2 multicomputer. Concurrency: Practice and Experience, 7(6):571–586, September 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Getov:1998:HPP

[GHM98] Vladimir Getov, Susan Flynn Hummel, and Sava Mintchev. High-performance parallel programming in Java: exploiting native libraries. Concurrency: Practice and Experience, 10(11–13):863–872, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050396;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050396&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Gupta:1998:TOC

[GHSJ98] S. K. S. Gupta, C.-H. Huang, P. Sadayappan, and R. W. Johnson. A technique for overlapping computation and communication for block recursive algorithms. Concurrency: Practice and Experience, 10(2):73–90, February 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=5367;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5367&PLACEBO=IE.pdf.

Gimenez:1997:BJM

[GHvdGV97] D. Giménez, V. Hernández, R. van de Geijn, and A. M. Vidal. A block Jacobi method on a mesh of processors. Concurrency: Practice and Experience, 9(5):391–411, May 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13866;
 http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?
                                        ID=13866&PLACEBO=IE.pdf.

Gorton:1996:RPS

[GJGC96] Ian Gorton, Innes Jelly, Jonathan Gray, and Toong Shoon Chan. Reliable parallel software construction using PARSE. Concurrency: Practice and Experience, 8(2):125–146, March 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23269.

Germain:1996:VAM

[GLV96] Daniel Germain, Denis Laurendeau, and Guy Vézina. Visibility analysis on a massively data-parallel computer. Concurrency: Practice and Experience, 8(6):475–487, July 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23288.

Gebremedhin:2000:SPG

[GM00] Assefaw Hadish Gebremedhin and Fredrik Manne. Scalable parallel graph coloring algorithms. Concurrency: Practice and Experience, 12 (12):1131–1146, October 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/76500356/
                                        START;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=76500356&PLACEBO=IE.pdf.

Gerogiannis:1989:HCD

[GOJ89] D. C. Gerogiannis, S. C. Orphanoudakis, and S. L. Johnsson. Histogram computation on distributed memory architectures. Concurrency: Practice and Experience, 1(2):219– 238, December 1989. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Gil:2000:PVT

[GOM00] C. Gil, J. Ortega, and M. G.

Montoya. Parallel VLSI test in a shared-memory multiprocessor. Concurrency: Practice and Experience, 12(5): 311–326, April 25, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72509236/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72509236&PLACEBO=IE.pdf.

Gorton:1993:PPD

[Gor93] I. Gorton. Parallel program design using high-level Petri nets. Concurrency: Practice and Experience, 5(2):87–107, April 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Gladychev:1998:CRJ

[GPO98] Pavel Gladychev, Ahmed Patel, and Donal O’Mahony. Cracking RC5 with Java applets. Concurrency: Practice and Experience, 10(11–13):1165–1171, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050408;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050408&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Grabienski:1991:DPM

[Gra91]  P. Grabienski.  DACAPO-

III: Parallel multilevel hardware simulation on transputers. Concurrency: Practice and Experience, 3(4):381–386, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

George:1997:ASS

[GRS^{+}97] Felicity George, Nicholas Radcliffe, Mark Smith, Mark Birkin, and Martin Clarke. Algorithms for solving a spatial optimisation problem on a parallel computer. Concurrency: Practice and Experience, 9(8):753–780, August 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13888;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13888&PLACEBO=IE.pdf.

Geist:1992:NBC

[GS92] G. A. Geist and V. S. Sunderam. Network-based concurrent computing on the PVM system. Concurrency: Practice and Experience, 4(4):293–311, June 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Garg:1997:PMD

[GS97a] V. Garg and D. E. Schimmel. Performance modeling of dense Cholesky factorization on the MasPar MP-2. Concurrency: Practice and Experience, 9(7):697–719, July 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13885;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13885&PLACEBO=IE.pdf.

Gray:1997:IDC

[GS97b] Paul A. Gray and Vaidy S.

Sunderam. IceT: distributed computing and Java. Concurrency: Practice and Experience, 9(11):1161–1167, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13820;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13820&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Gray:1998:NLB

[GS98] Paul A. Gray and Vaidy S. Sunderam. Native-language-based distributed computing across network and filesystem boundaries. Concurrency: Practice and Experience, 10(11–13):901–909, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050397;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050397&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Glossner:1998:JSP

[GTV98] John Glossner, Jesse Thilo, and Stamatis Vassiliadis. Java signal processing: FFTs with bytecodes. Concurrency: Practice and Experience, 10(11–13): 1173–1178, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050412;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050412&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Guerraoui:2000:SIE

[Gue00]  Rachid Guerraoui.  Special

issue: European Conference on Object-oriented Programming 1999. Concurrency: Practice and Experience, 12 (14):1345, December 10, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76500718/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500718&PLACEBO=IE.pdf.

Grimshaw:1993:NPG

[GWP93] A. S. Grimshaw, E. A. West, and W. R. Pearson. No pain and gain! — experiences with Mentat on a biological application. Concurrency: Practice and Experience, 5(4):309–328, June 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Haupt:2000:GSU

[HAFY00] Tomasz Haupt, Erol Akarsu, Geoffrey Fox, and Choon-Han Youn. The Gateway system: uniform Web based access to remote resources. Concurrency: Practice and Experience, 12(8):629–642, July 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516219/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516219&PLACEBO=IE.pdf.

Hummel:1997:AJB

[HAKN97] Joseph Hummel, Ana Azevedo, David Kolson, and Alexandru Nicolau. Annotating the Java bytecodes in support of optimization. Concurrency: Practice and Experience, 9(11):1003–1016, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13830;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13830&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Hansen:1993:MPC

[Han93a] Per Brinch Hansen. Model programs for computational science: A programming methodology for multicomputers. Concurrency: Practice and Experience, 5(5):407–423, August 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hansen:1993:PCA

[Han93b] Per Brinch Hansen. Parallel cellular automata: A model program for computational science. Concurrency: Practice and Experience, 5(5):425–448, August 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hansen:1994:DHS

[Han94a] P. Brinch Hansen. Do hypercubes sort faster than tree machines? Concurrency: Practice and Experience, 6(2):143–151, April 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hansen:1994:SPL

[Han94b] P. Brinch Hansen. SuperPascal — a publication language for parallel scientific computing. Concurrency: Practice and Experience, 6(5):461–483, August 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hansen:1998:CGS

[Han98]  Per Brinch Hansen.  Con-

jugate gradient solution of linear equations. Concurrency: Practice and Experience, 10(2):139–156, February 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=5370;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5370&PLACEBO=IE.pdf.

Hansen:2000:LEP

[Han00] Per Brinch Hansen. Letter to the Editor: ‘Parallelism and recursion in message passing libraries: An efficient methodology’, C. Rodríguez, F. de Sande, C. León and L. García, Concurrency: Practice and Experience 1999; 11(7):355–365. Concurrency: Practice and Experience, 12(15):1511–1513, December 25, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/76505649/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76505649&PLACEBO=IE.pdf.

See [RDLG99, Leó00].

Hisley:2000:PPE

[HASnP00] Dixie Hisley, Gagan Agrawal, Punyam Satya-narayana, and Lori Pollock. Porting and performance evaluation of irregular codes using OpenMP. Concurrency: Practice and Experience, 12(12):1241–1259, October 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/76500349/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76500349&PLACEBO=IE.pdf.

Hu:1998:OMA

[HBE98] Y. F. Hu, R. J. Blake, and D. R. Emerson. An optimal migration algorithm for dynamic load balancing. Concurrency: Practice and Experience, 10(6):467–483, May 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=5389;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=5389&PLACEBO=IE.pdf.

Heber:2000:SAW

[HBG00] Gerd Heber, Rupak Biswas, and Guang R. Gao. Selfavoiding walks over adaptive unstructured grids. Concurrency: Practice and Experience, 12(2-3):85–109, February– March 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72504929/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72504929&PLACEBO=IE.pdf.

Hartenstein:1996:HPC

[HBKR96] Reiner W. Hartenstein, Jürgen Becker, Rainer Kress, and Helmut Reinig. High-performance computing using a reconfigurable accelerator. Concurrency: Practice and Experience, 8(6):429–443, July 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=23286.

Heiss:1996:PMP

[HD96] Hans-Ulrich Heiss and Marcus Dormanns. Partitioning and mapping of parallel programs by self-organization. Concurrency: Practice and Experience, 8(9):685–706, November 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23304.

Hodgson:2000:ESO

[HDF00] S. Hodgson, N. Dunstan, and I. Fris. Extended semaphore operations. Concurrency: Practice and Experience, 12(15): 1495–1509, December 25, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/76505646/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76505646&PLACEBO=IE.pdf.

Helin:1993:PAC

[Hel93] Jukka Helin. Performance analysis of the CM-2, a massively parallel SIMD computer. Concurrency: Practice and Experience, 5(1):71–85, February 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hobbs:1999:CPC

[HG99] M. Hobbs and A. Goscinski. A concurrent process creation service to support SPMD based parallel processing on COWs. Concurrency: Practice and Experience, 11(13):803–821, November 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/67500118/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=67500118&PLACEBO=IE.pdf.

Hui:1996:SPP

[HHA96] Chi-Chung Hui, Mounir Hamdi, and Ishfaq Ahmad. SPEED: A parallel platform for solving and predicting the performance of PDEs on distributed systems. Concurrency: Practice and Experience, 8(7):537–568, September 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23294.

Hey:1995:GDM

[HHGW95] A. J. G. Hey, R. Hockney, V. Getov, and I. Wolton. The Genesis distributed memory benchmarks. Part 2: COMMS1, TRANS1, FFT1 and QCD2 benchmarks on the SUPRENUM and IPSC/860 computers. Concurrency: Practice and Experience, 7(6):543–570, September 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hameed:1997:CRC

[HHKP97] F. Hameed, S. E. Hambrusch, A. A. Khokhar, and J. N. Patel. Contour ranking on coarse grained machines: a case study for low-level vision computations. Concurrency: Practice and Experience, 9(3):203–221, March 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13857;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13857&PLACEBO=IE.pdf.

Hartel:1995:TPF

[HHLM95] P. H. Hartel, R. F. H. Hofman, K. G. Langendoen, and H. L. Muller. A toolkit for parallel functional programming. Concurrency: Practice and Experience, 7(8):765–793, December 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hinrichs:1997:CRM

[Hin97] S. Hinrichs. Connection resource management for compilergenerated communication. Concurrency: Practice and Experience, 9(2):85–112, February 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13850;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13850&PLACEBO=IE.pdf.

Hamalainen:1997:PRK

[HKSK97] Timo Hämäläinen, Harri Klapuri, Jukka Saarinen, and Kimmo Kaski. Parallel realizations of Kanerva’s sparse distributed memory on a treeshaped computer. Concurrency: Practice and Experience, 9(9):877–896, September 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13895;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13895&PLACEBO=IE.pdf.

Hamidzadeh:1995:DST

[HLA95] B. Hamidzadeh, D. J. Lilja, and Y. Atif. Dynamic scheduling techniques for heterogeneous computing systems. Concurrency: Practice and Experience, 7(7):633–652, October 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Huss-Lederman:1994:MMI

[HLJTZ94] S. Huss-Lederman, E. M. Jacobson, A. Tsao, and G. Zhang. Matrix multiplication on the Intel Touchstone Delta. Concurrency: Practice and Experience, 6(7):571–594, October 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hoogerbrugge:1995:ENP

[HM95] P. Hoogerbrugge and R. Mirchandaney. Experiences with networked parallel computing. Concurrency: Practice and Experience, 7(1):1–16, February 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hall:1998:APC

[HM98] Mary W. Hall and Margaret Martonosi. Adaptive parallelism in compiler-parallelized code. Concurrency: Practice and Experience, 10(14): 1235–1250, December 10, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=40005934;
 http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?
                                        ID=40005934&PLACEBO=IE.pdf.

Heydon:2000:PLJ

[HN00] Allan Heydon and Marc Najork. Performance limitations of the Java core libraries. Concurrency: Practice and Experience, 12(6):363–373, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/72515723/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72515723&PLACEBO=IE.pdf.

Hardwick:1997:ISW

[HNS97a] Jonathan C. Hardwick, Girija Narlikar, and Jay Sipelstein. Interactive simulations on the Web: compiling N ESL into Java. Concurrency: Practice and Experience, 9(11):1075–1089, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13829;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13829&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Hummel:1997:SPJ

[HNS97b] Susan Flynn Hummel, Ton Ngo, and Harini Srinivasan. SPMD programming in Java. Concurrency: Practice and Experience, 9(6):621–631, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13879;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13879&PLACEBO=IE.pdf.

Hoare:1991:TOP

[Hoa91] C. A. R. Hoare. The transputer and occam: A personal story. Concurrency: Practice and Experience, 3(4):249–264, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hockney:1995:CS

[Hoc95] R. W. Hockney. Computational similarity. Concurrency: Practice and Experience, 7(2): 147–166, April 1995. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hariri:1994:CSH

[HPPF94] S. Hariri, J.-B. Park, M. Parashar, and G. C. Fox. Communication system for high-performance distributed computing. Concurrency: Practice and Experience, 6(4):251–270, June 1994. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Ho:1992:ECP

[HR92] C.-T. Ho and M. T. Raghunath.

Efficient communication primitives on hypercubes. Concurrency: Practice and Experience, 4(6):427–457, September 1992.

CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hanson:1990:SPP

[HS90] F. B. Hanson and D. C. Sorensen. The SCHEDULE parallel programming package with recycling job queues and iterated dependency graphs. Concurrency: Practice and Experience, 2(1):33–54, March 1990. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Homer:1994:USS

[HS94] P. T. Homer and R. D. Schlichting. Using Schooner to support distribution and heterogeneity in the Numerical Propulsion System Simulation project. Concurrency: Practice and Experience, 6(4):271–287, June 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Sun:1997:PCS

[hSM97] Xian he Sun and Stuti Moitra. Performance comparison of a set of periodic and non-periodic tridiagonal solvers on SP2 and Paragon parallel computers. Concurrency: Practice and Experience, 9(8):781–801, August 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13889;
 http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=13889&PLACEBO=IE.pdf.

Hoekstra:1996:NGP

[HSvdL^{+}96] A. G. Hoekstra, P. M. A. Sloot, F. van der Linden, M. van Muiswinkel, J. J. J. Vesseur, and L. O. Hertzberger. Native and generic parallel programming environments on a transputer and a PowerPC platform. Concurrency: Practice and Experience, 8(1):19– 46, January 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

URL http://www.fwi.uva.
nl/fwi/research/vg4/pwrs/
papers/archive/Hoekstra95_
1.ps.gz; http://www3.interscience.
wiley.com/cgi-bin/abstract?
ID=23257.

Hall:1991:RTD

[HTM91] G. Hall, T. J. Terrell, and L. M. Murphy. The Radon transform on a dynamically switched transputer network. Concurrency: Practice and Experience, 3(4):315–323, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hariri:1993:HDC

[HV93] S. Hariri and A. Varma. Highperformance distributed computing: Promises and challenges. Concurrency: Practice and Experience, 5(4):233–238, June 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hopkins:1996:TDF

[HW96] Tim Hopkins and Peter Welch. Transputer data-flow solution for systems of linear equations. Concurrency: Practice and Experience, 8(8):569–580, October 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23298.

Herlihy:2000:TTD

[HW00] Maurice Herlihy and Michael P. Warres. A tale of two directories: implementing distributed shared objects in Java. Concurrency: Practice and Experience, 12(7):555–572, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/72516215/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72516215&PLACEBO=IE.pdf.

Hauser:1992:PCA

[HWPG92] J. Häuser, H. Wong, H. G. Paap, and W. Gentzsch. Parallel computing in aerospace using multi-block grids. Part 1: Application to grid generation. Concurrency: Practice and Experience, 4(5):357–376, August 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Hirano:1998:PEP

[HYI98] Satoshi Hirano, Yoshiji Yasu, and Hirotaka Igarashi. Performance evaluation of popular distributed object technologies for Java. Concurrency: Practice and Experience, 10(11–13):927–940, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050399;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050399&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Izatt:2000:ATE

[ICB00] Matthew Izatt, Patrick Chan, and Tim Brecht. Ajents: towards an environment for parallel, distributed and mobile Java applications. Concurrency: Practice and Experience, 12(8):667–685, July 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516221/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516221&PLACEBO=IE.pdf.

Ishizaki:2000:DIE

[IKY^{+}00] Kazuaki Ishizaki, Motohiro Kawahito, Toshiaki Yasue, Mikio Takeuchi, Takeshi Ogasawara, Toshio Suganuma, Tamiya Onodera, Hideaki Komatsu, and Toshio Nakatani. Design, implementation, and evaluation of optimizations in a Java^{T M} Just-In-Time compiler. Concurrency: Practice and Experience, 12(6):457–475, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/72515728/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72515728&PLACEBO=IE.pdf.

Jain:1998:DDO

[JA98] Lalit Kumar Jain and Mark Abbott. DOVE: distributed objects based scientific visualization environment. Concurrency: Practice and Experience, 10(11–13):1087–1095, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050410;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050410&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Judd:1998:DDO

[JCS98] Glenn Judd, Mark Clement, and Quinn Snell. DOGMA: distributed object group metacomputing architecture. Concurrency: Practice and Experience, 10(11–13):977–983, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050409;

http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050409&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Jagannathan:1997:GHL

[JDA97] R. Jagannathan, C. Dodd, and I. Agi. GLU: A highlevel system for granular dataparallel programming. Concurrency: Practice and Experience, 9(1):63–83, January 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13806;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13806&PLACEBO=IE.pdf.

Justo:1999:APP

[JDZW99] G. R. Ribeiro Justo, T. Delaitre, M. J. Zemerly, and S. C. Winter. Accurate performance prediction using visual prototypes. Concurrency: Practice and Experience, 11(11):615–634, September 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=64500092;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=64500092&PLACEBO=IE.pdf.

Jin:2000:OSR

[JH00] Hai Jin and Kai Hwang. Optimal striping in RAID architecture. Concurrency: Practice and Experience, 12(10): 909–916, August 25, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/73000569/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=73000569&PLACEBO=IE.pdf.

Jard:1992:EEC

[JJ92] C. Jard and J.-M. Jézéquel. Echidna, an Estelle compiler to prototype protocols on distributed computers. Concurrency: Practice and Experience, 4(5):377–397, August 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Jezequel:1996:BGC

[JJ96] Jean-Marc Jézéquel and Claude Jard. Building a global clock for observing computations in distributed memory parallel computers. Concurrency: Practice and Experience, 8(1):71–89, January 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=23259.

Jones:2000:AJC

[JK00] Joel Jones and Samuel Kamin. Annotating Java class files with virtual registers for performance. Concurrency: Practice and Experience, 12(6):389–406, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72515727/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72515727&PLACEBO=IE.pdf.

Jia:1995:DIC

[JL95] W. Jia and G. Libert. Distributed implementations of communicating objects. Concurrency: Practice and Experience, 7(6):515–541, September 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Jones:1989:PER

[JM89] P. Jones and A. Murta. Practical experience of run-time link reconfiguration in a multitransputer machine. Concurrency: Practice and Experience, 1(2):171–190, December 1989. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Joe:1990:CLN

[JMM90] K. Joe, Y. Mori, and S. Miyake.

Construction of a large-scale neural network: Simulation of handwritten Japanese character recognition on NCUBE. Concurrency: Practice and Experience, 2(2):79–108, June 1990. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Jacob:1998:LSP

[JPK98] Matthias Jacob, Michael Philippsen, and Martin Karrenbach. Largescale parallel geophysical algorithms in Java: a feasibility study. Concurrency: Practice and Experience, 10 (11–13):1143–1153, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050403;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050403&PLACEBO=IE.pdf;
http://wwwipd.ira.uka.de/~
phlipp/veltran.ps.gz. Spe-

cial Issue: Java for Highperformance Network Computing.

Jensen:1992:PAE

[JR92] D. W. Jensen and D. A. Reed. A performance analysis exemplar: Parallel ray tracing. Concurrency: Practice and Experience, 4(2):119–141, April 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

James:2000:DKD

[JS00] Jerry James and Ambuj Singh. Design of the Kan distributed object system. Concurrency: Practice and Experience, 12(8):755–797, July 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516222/
START;   http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=72516222&PLACEBO=IE.pdf.

Jackson:1996:PAD

[JT96] David Jeff Jackson and Greg Scott Tinney. Performance analysis of distributed implementations of a fractal image compression algorithm. Concurrency: Practice and Experience, 8(5):357–386, June 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23284.

Khaddaj:1991:SRE

[KABG^{+}91] S. A. Khaddaj, H. Al-Bahadili, A. J. H. Goddard, C. R. E. de Oliveira, and J. Wood. The solution of radiation engineering problems on a transputerbased system. Concurrency: Practice and Experience, 3(4): 423–434, August 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Karl:1998:BGB

[Kar98] Holger Karl. Bridging the gap between distributed shared memory and message passing. Concurrency: Practice and Experience, 10(11– 13):887–900, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050425;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050425&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Karimi:1999:SDT

[Kar99] Jahangir Karimi. A software design technique for clientserver applications. Concurrency: Practice and Experience, 11(1):21–42, January 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=40006534;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=40006534&PLACEBO=IE.pdf.

Kohli:1998:IUL

[KAS98] Prince Kohli, Mustaque Ahamad, and Karsten Schwan. Indigo: user-level support for building distributed shared abstractions. Concurrency: Practice and Experience, 10(1):1– 29, January 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

                                        URL http://www.cc.gatech.
                                        edu/grads/k/Prince.Kohli/
                                        indigo_jour.ps.gz;   http:
                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=5363;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=5363&PLACEBO=IE.pdf.

Keren:1998:APP

[KB98] Arie Keren and Amnon Barak. Adaptive placement of parallel Java agents in a scalable computing cluster. Concurrency: Practice and Experience, 10(11–13):971–976, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050404;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050404&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Keane:1996:PSF

[Kea96] J. A. Keane. Parallel systems in financial information processing. Concurrency: Practice and Experience, 8(10):757–768, December 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23262.

Keane:1999:COF

[Kea99] J. A. Keane. A co-ordination framework for distributed financial applications. Concurrency: Practice and Experience, 11(4):205–219, April 10, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=61004081;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=61004081&PLACEBO=IE.pdf.

Krall:1997:CBJ

[KG97] Andreas Krall and Reinhard Grafl. CACAO — A 64bit JavaVM just-in-time compiler. Concurrency: Practice and Experience, 9(11): 1017–1030, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13831;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13831&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Kim:1990:SPA

[KK90] Y. Kim and M. Kim. A stepwise-overlapped parallel annealing algorithm on a messagepassing multiprocessor system. Concurrency: Practice and Experience, 2(2):123–148, June 1990. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Ki:1997:CSD

[KK97] Byeongseob Ki and Scott Klasky. Collaborative scientific data visualization. Concurrency: Practice and Experience, 9(11):1249–1259, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13822;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13822&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Ki:1998:S

[KK98] Byeongseob Ki and Scott Klasky. Scivis. Concurrency: Practice and Experience, 10(11–13):1107–1115, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050426;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050426&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Knobe:1993:OTS

[KLW93] K. Knobe, J. D. Lukas, and M. Weiss. Optimization techniques for SIMD Fortran compilers. Concurrency: Practice and Experience, 5(7):527– 552, October 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Kennedy:1993:ATI

[KMT93] K. Kennedy, K. S. McKinley, and C.-W. Tseng. Analysis and transformation in an interactive parallel programming tool. Concurrency: Practice and Experience, 5(7):575–602, October 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/suif.stanford.edu/~tseng/
papers/cpe93.ps.

Kontoghiorghes:1999:OLM

[Kon99] Erricos John Kontoghiorghes. Ordinary linear model estimation on a massively parallel SIMD computer. Concurrency: Practice and Experience, 11(7):323–341, June 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=62500300;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=62500300&PLACEBO=IE.pdf.

Kumar:1991:PTF

[KP91] S. P. Kumar and I. R. Philips. Portable tools for Fortran parallel programming. Concurrency: Practice and Experience, 3(6): 559–572, December 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Krall:1998:MEH

[KP98] Andreas Krall and Mark Probst. Monitors and exceptions: how to implement Java efficiently. Concurrency: Practice and Experience, 10(11– 13):837–850, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050393;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050393&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Kuhn:2000:OVT

[KPO00] Bob Kuhn, Paul Petersen, and Eamonn O’Toole. OpenMP versus threading in C/C++. Concurrency: Practice and Experience, 12(12):1165–1176, October 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/76500354/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500354&PLACEBO=IE.pdf.

Knop:1996:FSC

[KRS96] Felipe Knop, Vernon Rego, and Vaidy Sunderam. Fail-safe concurrency in the EcliPSe system. Concurrency: Practice and Experience, 8(4):283–312, May 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23279.

Kenyon:1995:PPC

[KSAC95] P. Kenyon, S. Seth, P. Agrawal, and A. Clematis. Programming pipelined CAD applications on message-passing architectures. Concurrency: Practice and Experience, 7(4):315–337, June 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Konstantopoulos:2000:EPA

[KSK00] Charalampos Konstantopoulos, Andreas Svolos, and Christos Kaklamanis. An efficient parallel algorithm for motion estimation in very low bit-rate video coding systems. Concurrency: Practice and Experience, 12(5):289–309, April 25, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72509235/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72509235&PLACEBO=IE.pdf.

Koide:1997:NMA

[KSN97] Hiroshi Koide, Mitsugu Suzuki, and Yasuichi Nakayama. A new memory allocation method for shared memory multiprocessors with large virtual address space. Concurrency: Practice and Experience, 9(9):897–914, September 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13896;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13896&PLACEBO=IE.pdf.

Kindler:1996:PSM

[KSS^{+}96] Thomas Kindler, Karsten Schwan, Dilma Silva, Mary Trauner, and Fred Alyea. A parallel spectral model for atmospheric transport processes. Concurrency: Practice and Experience, 8(9):639–666, November 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23302.

Krasteva:1999:DCP

[KWB^{+}99] Denitza T. Krasteva, Layne T.

Watson, Chuck Baker, Bernard Grossman, William H. Mason, and Raphael T. Haftka. Distributed control parallelism in multidisciplinary aircraft design. Concurrency: Practice and Experience, 11(8):435–459, July 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=63000993;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=63000993&PLACEBO=IE.pdf.

Kennedy:1993:E

[KZ93] Ken Kennedy and Hans Zima.

Editorial. Concurrency: Practice and Experience, 5(7):iii– iv, October 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Landry:1995:IFS

[LA95] K. Landry and J. D. Arthur. Instructional footprinting and semantic preservation in Linda. Concurrency: Practice and Experience, 7(3):191–207, May 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Lastovetsky:2000:PLP

[LAKL00] Alexey Lastovetsky, Dmitry Arapov, Alexey Kalinov, and Ilya Ledovskih. A parallel language and its programming system for heterogeneous networks. Concurrency: Practice and Experience, 12(13): 1317–1343, November 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76500626/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500626&PLACEBO=IE.pdf.

Lapegna:1992:GAQ

[Lap92] M. Lapegna. Global adaptive quadrature for the approximate computation of multidimensional integrals on a distributedmemory multiprocessor. Concurrency: Practice and Experience, 4(6):413–426, September 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Lu:1997:WBD

[LCLL97] Tain-Chi Lu, Chung-Wen Chiang, Chungnan Lee, and Tong-Yee Lee. A Web-based distributed and collaborative 3D animation environment. Concurrency: Practice and Experience, 9(11):1261–1268, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13840;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13840&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Leuze:1989:MDM

[LDP89] M. R. Leuze, L. W. Dowdy, and K. H. Park. Multiprogramming a distributed-memory multiprocessor. Concurrency: Practice and Experience, 1(1):19– 34, September 1989. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Leon:2000:ARP

[Leó00] Casiano Rodríguez León. Authors’ response: ‘Parallelism and recursion in message passing libraries: An efficient methodology’, C. Rodríguez, F. de Sande, C. León and L. García, Concurrency: Practice and Experience 1999; 11(7):355–365. Concurrency: Practice and Experience, 12 (15):1515–1516, December 25, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76505647/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76505647&PLACEBO=IE.pdf.

See [RDLG99, Han00].

Lin:1991:ATI

[LF91] W. Lin and D. A. Fraser. Algorithms and timing for identification of objects from 2-D images. Concurrency: Practice and Experience, 3(4):325–331, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Lou:1998:PAO

[LF98] John Z. Lou and John D. Farrara. Performance analysis and optimization on a parallel atmospheric general circulation model code. Concurrency: Practice and Experience, 10(7):549–565, June 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10008692;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10008692&PLACEBO=IE.pdf.

Lowenthal:1998:ESF

[LFA98] David K. Lowenthal, Vincent W. Freeh, and Gregory R. Andrews. Efficient support for fine-grain parallelism on shared-memory machines. Concurrency: Practice and Experience, 10(3):157–173, March 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=5372;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=5372&PLACEBO=IE.pdf.

Levine:1999:PCT

[LGFK99] David Levine, William Gropp, Kimmo Forsman, and Lauri Kettunen. Parallel computation of three-dimensional nonlinear magnetostatic problems.

Concurrency: Practice and Experience, 11(2):109–120, February 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=61004054;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=61004054&PLACEBO=IE.pdf.

Laure:2000:IOC

[LHMZ00] Erwin Laure, Matthew Haines, Piyush Mehrotra, and Hans Zima. On the implementation of the Opus coordination language. Concurrency: Practice and Experience, 12(4):227–249, April 10, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72504933/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72504933&PLACEBO=IE.pdf.

Lilja:1995:PTB

[Lil95] D. J. Lilja. Partitioning tasks between a pair of interconnected heterogeneous. Concurrency: Practice and Experience, 7(3):209–223, May 1995. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Lin:1997:TDP

[Lin97] Weisi Lin. Task division for parallel implementation of object identification system based on alternating hypothesizeverify-extend strategy. Concurrency: Practice and Experience, 9(9):859–876, September 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13894;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13894&PLACEBO=IE.pdf.

Liebrock:1996:PSC

[LK96] Lorie M. Liebrock and Ken Kennedy. Parallelization support for coupled grid applications with small meshes. Concurrency: Practice and Experience, 8(8):581–615, October 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23299.

Lu:1995:AAR

[LL95] C. Lu and S.-M. Lau. An adaptive algorithm for resolving processor thrashing in load distribution. Concurrency: Practice and Experience, 7(7):653– 670, October 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Lanteri:1996:LSS

[LL96] S. Lanteri and M. Loriot. Large-scale solutions of threedimensional compressible flows using the parallel N3S-MUSCL solver. Concurrency: Practice and Experience, 8(10):769– 798, December 1996. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (elec-

tronic). URL http://www3.
interscience.wiley.com/cgi-
bin/abstract?ID=23264.

Lu:1999:ALD

[LL99] Qin Lu and Sau-Ming Lau.

An adaptive load distribution algorithm for resolving bursty workload. Concurrency: Practice and Experience, 11(1):1–20, January 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=40006533;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=40006533&PLACEBO=IE.pdf.

Lu:1998:DLD

[LLL98] Qin Lu, Sau-Ming Lau, and Kwong-Sak Leung. Dynamic load distribution using antitasks and load state vectors. Concurrency: Practice and Experience, 10(14): 1251–1269, December 10, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=40005935;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=40005935&PLACEBO=IE.pdf.

Lin:2000:SFC

[LP00] Yuan Lin and David Padua.

A simple framework to calculate the reaching definition of array references and its use in subscript array analysis. Concurrency: Practice and Experience, 12(2-3): 125–135, February–March 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72504940/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72504940&PLACEBO=IE.pdf.

Leon:1997:POP

[LRGS97] C. León, C. Rodríguez, F. García, and F. De Sande. A PRAM oriented programming system. Concurrency: Practice and Experience, 9(3):163–179, March 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13855;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13855&PLACEBO=IE.pdf.

Lee:1997:PIR

[LRN97] Tong-Yee Lee, C. S. Raghavendra, and John B. Nicholas. Parallel implementation of a ray tracing algorithm for distributed memory parallel computers. Concurrency: Practice and Experience, 9 (10):947–965, October 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13810;
 http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=13810&PLACEBO=IE.pdf.

Li:1997:PAP

[LSF97] J. Li, A. Skjellum, and R. D.

Falgout. A poly-algorithm for parallel dense matrix multiplication on two-dimensional process grid topologies. Concurrency: Practice and Experience, 9(5):345–389, May 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13865;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13865&PLACEBO=IE.pdf.

Lecomber:2000:PPT

[LSS00] D. S. Lecomber, C. J. Siniolakis, and K. R. Sujithan. PRAM programming: in theory and in practice. Concurrency: Practice and Experience, 12(4):211–226, April 10, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72504942/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72504942&PLACEBO=IE.pdf.

Loidl:1999:EPS

[LTH^{+}99] Hans-Wolfgang Loidl, Philip W.

Trinder, Kevin Hammond, Sahalu B. Junaidu, Richard G. Morgan, and Simon L. Peyton Jones. Engineering parallel symbolic programs in G P H.

Concurrency: Practice and Experience, 11(12):701–752, October 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/66002640/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=66002640&PLACEBO=IE.pdf.

Li:1996:PDA

[LZM96] Xi Li, Sotirios G. Ziavras, and Constantine N. Manikopoulos. Parallel DSP algorithms on TurboNet: an experimental system with hybrid messagepassing/shared-memory architecture. Concurrency: Practice and Experience, 8(5):387–411, June 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23283.

Li:1997:PGA

[LZM97] Xi Li, Sotirios G. Ziavras, and Constantine N. Manikopoulos. Parallel generation of adaptive multiresolution structures for image processing. Concurrency: Practice and Experience, 9(4):241–254, April 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13860;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13860&PLACEBO=IE.pdf.

Mans:1998:PDP

[Man98] Bernard Mans. Portable distributed priority queues with MPI. Concurrency: Practice and Experience, 10(3):175–198, March 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=5373;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5373&PLACEBO=IE.pdf.

Murillo:1999:DSU

[MB99] Alejandro E. Murillo and Jean Bell. Distributed Seismic Unix: a tool for seismic data processing. Concurrency: Practice and Experience, 11(4):169–187, April 10, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=61004082;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=61004082&PLACEBO=IE.pdf.

Special Issue: Applications of Distributed Computing Environments.

Messina:1990:BAA

[MBF^{+}90] P. Messina, C. Baillie, E. Felten, P. Hipes, R. Williams, A. Alagar, A. Kamrath, R. Leary, W. Pfeiffer, J. Rogers, and D. Walker. Benchmarking advanced architecture computers. Concurrency: Practice and Experience, 2(3):195–255, September 1990. CODEN CPEXEI.

ISSN 1040-3108 (print), 1096-

9128 (electronic).

Marsland:1991:NME

[MBS91] T. A. Marsland, T. Breitkruetz, and S. Sutphen. A network multi-processor for experiments in parallelism. Concurrency: Practice and Experience, 3(3): 203–219, June 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Morrow:1999:EIP

[MCB^{+}99] P. J. Morrow, D. Crookes, J. Brown, G. McAleese, D. Roantree, and I. Spence. Efficient implementation of a portable parallel programming model for image processing. Concurrency: Practice and Experience, 11(11):671–685, September 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=64500089;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=64500089&PLACEBO=IE.pdf.

Messina:1991:PCP

[Mes91] P. C. Messina. Parallel computing in the 1980s — one person’s view. Concurrency: Practice and Experience, 3(6): 501–524, December 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Mansour:1992:ADM

[MF92] N. Mansour and G. C. Fox. Allocating data to multicomputer nodes by physical optimization algorithms for loosely synchronous computations. Concurrency: Practice and Experience, 4(7):557–574, October 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

McKerrow:1993:IPA

[MF93] P. J. McKerrow and J. A.

Fulcher. Implementation of parallel algorithms for controlling robot manipulators. Concurrency: Practice and Experience, 5(3):213–232, May 1993. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Mansour:1994:ADD

[MF94] N. Mansour and G. C. Fox.

Allocating data to distributedmemory multiprocessors by genetic algorithms. Concurrency: Practice and Experience, 6(6): 485–504, September 1994. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Montoya:2000:PTA

[MGG00] M. G. Montoya, C. Gil, and I. Garcia. Parallel thinning algorithms on multicomputers: experimental study on load balancing. Concurrency: Practice and Experience, 12(5):327–340, April 25, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72509234/
START;   http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?
                                        ID=72509234&PLACEBO=IE.pdf.

Mirmehdi:1991:PLI

[Mir91] M. Mirmehdi. Product label inspection using transputers. Concurrency: Practice and Experience, 3(4):265–273, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Mullin:1996:EDP

[MJ96] L. M. R. Mullin and M. A. Jenkins. Effective data parallel computation using the Psi calculus. Concurrency: Practice and Experience, 8(7):499–515, September 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23296.

Mcmanus:1999:CBM

[MJC99] K. Mcmanus, S. P. Johnson, and M. Cross. Converting best manual practice into generic automatable strategies for unstructured mesh parallelisation. Concurrency: Practice and Experience, 11 (11):593–614, September 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=64500093;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=64500093&PLACEBO=IE.pdf.

Mayes:1993:LAA

[MK93] K. R. Mayes and J. A. Keane.

Levels of atomic action in the Flagship parallel system. Concurrency: Practice and Experience, 5(3):193–212, May 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

McCormack:1999:SCW

[MKD99] Ryan P. McCormack, John E.

Koontz, and Judith Devaney. Seamless computing with Web-Submit. Concurrency: Practice and Experience, 11(15): 949–963, December 25, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/71005730/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=71005730&PLACEBO=IE.pdf.

McGrath:1997:JTA

[MLF97] Robert E. McGrath, Xinjian Lu, and Michael Folk. Java applications using NCSA HDF files. Concurrency: Practice and Experience, 9(11): 1113–1125, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13825;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13825&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Melhart:1995:CPA

[MMN95] B. E. Melhart, C. A. Morgenstern, and T. Nute. A compendium of processor allocation strategies for two-dimensional mesh connected systems. Concurrency: Practice and Experience, 7(5):497–514, August 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Mazzeo:1998:EMH

[MMV98] A. Mazzeo, N. Mazzocca, and U. Villano. Efficiency measurements in heterogeneous distributed computing systems: from theory to practice. Concurrency: Practice and Experience, 10(4):285–313, April 10, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=5379;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=5379&PLACEBO=IE.pdf.

Mills:1991:PFI

[MO91] E. Mills and B. C. O’Neill. Particle flow instrumentation. Concurrency: Practice and Experience, 3(4):275–281, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Martin:1995:PUM

[MO95] O. C. Martin and S. W. Otto.

Partitioning of unstructured meshes for load balancing. Concurrency: Practice and Experience, 7(4):303–314, June 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Minnich:1993:RHT

[MP93] R. G. Minnich and D. V.

Pryor. Radiative heat transfer simulation on a SPARCstation farm. Concurrency: Practice and Experience, 5(4):345–357, June 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Meyer:1996:FPQ

[MP96] Gerard G. L. Meyer and Mike Pascale. A family of parallel QR factorization algorithms. Concurrency: Practice and Experience, 8(6):461–473, July 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23290.

Maier:1995:PSL

[MPR95] R. S. Maier, L. R. Petzold, and W. Rath. Parallel solution of large-scale differential-algebraic systems. Concurrency: Practice and Experience, 7(8):795– 822, December 1995. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Montani:1993:PRV

[MPS93] C. Montani, R. Perego, and R. Scopigno. Parallel rendering of volumetric data set on distributed memory architectures. Concurrency: Practice and Experience, 5(2):153–167, April 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Marinescu:1991:SLI

[MR91] D. C. Marinescu and J. R. Rice. Synchronization and load imbalance effects in distributed memory multi-processor systems. Concurrency: Practice and Experience, 3(6):593– 625, December 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Malfait:1996:CCM

[MR96] Maurits Malfait and Dirk Roose. Convergence of concurrent Markov chain Monte-Carlo algorithms. Concurrency: Practice and Experience, 8(3):167–189, April 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=23272.

Marenzoni:1998:BKT

[MR98a] P. Marenzoni and P. Rossi. Benchmark kernels as a tool for performance evaluation of MPP’s. Concurrency: Practice and Experience, 10(1): 31–51, January 1998. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (elec-

tronic). URL http://www3.
interscience.wiley.com/cgi-
bin/abstract?ID=5364;   http:
//www3.interscience.wiley.
com/cgi-bin/fulltext?ID=5364&
PLACEBO=IE.pdf.

Mascarenhas:1998:MTP

[MR98b] Edward Mascarenhas and Vernon Rego. Migrant threads on process farms: parallel programming with Ariadne. Concurrency: Practice and Experience, 10(9):673–698, August 10, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10008703;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10008703&PLACEBO=IE.pdf.

Marinescu:1993:MED

[MRCH^{+}93] D. C. Marinescu, J. R. Rice, M. A. Cornea-Hasegan, R. E. Lynch, and M. G. Rossmann. Macromolecular electron density averaging on distributed memory MIMD systems. Concurrency: Practice and Experience, 5(8):635–657, December 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Morales:1996:PAI

[MRR^{+}96] D. Morales, J. L. Roda, C. Rodríguez, F. Almeida, and F. García. A parallel algorithm for the integer knapsack problem. Concurrency: Practice and Experience, 8(4):251–260, May 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23277.

Marcenaro:1991:AMI

[MT91] G. Marcenaro and M. Tistarelli. Analysis of multidimensional images on the Connection Machine system. Concurrency: Practice and Experience, 3(6): 699–713, December 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Mehta:1998:PAC

[MW98] Dinesh P. Mehta and Erica D. Wilson. Parallel algorithms for corner stitching. Concurrency: Practice and Experience, 10(15):1317–1341, December 25, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=40005939;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=40005939&PLACEBO=IE.pdf.

Narasimhan:1996:PMT

[NA96] V. Lakshmi Narasimhan and J. Armstrong. Performance modelling of three parallel sorting algorithms on a pipelined transputer network. Concurrency: Practice and Experience, 8(5):335–355, June 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=23282.

Nakano:1999:MLB

[Nak99] Aiichiro Nakano. Multiresolution load balancing in curved space: the wavelet representation. Concurrency: Practice and Experience, 11(7):343–353, June 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=62500297;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=62500297&PLACEBO=IE.pdf.

Nash:1999:SSF

[Nas99] Jonathan Nash. A scalable and starvation-free concurrent locking mechanism. Concurrency: Practice and Experience, 11(13):823–833, November 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/67500117/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=67500117&PLACEBO=IE.pdf.

Nicklas:1998:DIP

[NASW98] Lisa D. Nicklas, Robert W.

Atkins, Sanjeev K. Setia, and Pearl Y. Wang. Design and implementation of a parallel solution to the cutting stock problem. Concurrency: Practice and Experience, 10(10):783–805, August 25, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10006458;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10006458&PLACEBO=IE.pdf.

Neary:2000:JSI

[NBK^{+}00] Michael O. Neary, Sean P. Brydon, Paul Kmiec, Sami Rollins, and Peter Cappello. Javelin++: scalability issues in global computing. Concurrency: Practice and Experience, 12(8):727–753, July 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/72516224/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72516224&PLACEBO=IE.pdf.

Narayanan:1994:PSI

[ND94] P. J. Narayanan and L. S. Davis. Parallel search for the interpretation of aerial images. Concurrency: Practice and Experience, 6(6):517–541, September 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Norman:1991:UCE

[NHMW91] M. G. Norman, J. R. Henderson, I. G. Main, and D. J. Wallace. The use of the CAPE environment in the simulation of rock fracturing. Concurrency: Practice and Experience, 3(6): 687–698, December 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Naim:1996:DLS

[NHZ96] Oscar Naím, Tony Hey, and Ed Zaluska. Do-loop-surface: an abstract representation of parallel program performance. Concurrency: Practice and Experience, 8(3):205–234, April 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=23275.

Newhall:2000:PMD

[NM00] Tia Newhall and Barton P.

Miller. Performance measurement of dynamically compiled Java executions. Concurrency: Practice and Experience, 12(6):343–362, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72515726/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72515726&PLACEBO=IE.pdf.

Nedeljkovic:1993:DPN

[NQ93] N. Nedeljkovic and M. J. Quinn.

Data-parallel programming on a network of heterogeneous workstations. Concurrency: Practice and Experience, 5(4):257–268, June 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Neuman:1994:PRM

[NR94] B. C. Neuman and S. Rao. The Prospero Resource Manager: A scalable framework for processor allocation in distributed systems. Concurrency: Practice and Experience, 6(4):339–355, June 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Ng:2000:TSH

[NR00] Esmond G. Ng and Padma Raghavan. Towards a scalable hybrid sparse solver. Concurrency: Practice and Experience, 12(2-3):53–68, February– March 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/72504937/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72504937&PLACEBO=IE.pdf.

Olsson:1999:RES

[Ols99] Ronald A. Olsson. Reproducible execution of SR programs. Concurrency: Practice and Experience, 11(9): 479–507, August 10, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=63001744;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=63001744&PLACEBO=IE.pdf.

Orlando:1999:CRT

[OP99] Salvatore Orlando and Raffaele Perego. COLT_{H}P F, a run-time support for the high-level co-ordination of HPF tasks. Concurrency: Practice and Experience, 11(8):407–434, July 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=63000994;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=63000994&PLACEBO=IE.pdf.

Pande:1995:RIP

[PA95] S. Pande and D. P. Agrawal.

Run-time issues in program partitioning on distributed memory systems. Concurrency: Practice and Experience, 7(5): 429–454, August 1995. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Poggi:1999:MOO

[PD99] Agostino Poggi and Giulio Destri. MPOOL: an objectoriented library for task composition and co-ordination. Concurrency: Practice and Experience, 11(14):835–848, December 10, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/67501482/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=67501482&PLACEBO=IE.pdf.

Probert:1991:TAS

[PDH91] P. J. Probert, D. Djian, and H. Hu. Transputer architectures for sensing in a robot controller: Formal methods for design. Concurrency: Practice and Experience, 3(4):283– 292, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Pain:1999:SAT

[PdOG99] C. C. Pain, C. R. E. de Oliveira, and A. J. H. Goddard. Simulated annealing task to processor mapping for domain decomposition methods on distributed parallel computers. Concurrency: Practice and Experience, 11(3):155–165, March 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=61003666;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=61003666&PLACEBO=IE.pdf.

Perrott:1992:PLD

[Per92]  R. Perrott.  Parallel lan-

guage developments in European overview. Concurrency: Practice and Experience, 4(8): 589–617, December 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Pu:1993:PCD

[PFS^{+}93] C. Pu, D. Florissi, P. Soares, K. Wu, and P. S. Yu. Performance comparison of dynamic policies for remote caching. Concurrency: Practice and Experience, 5(4):239–256, June 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Philippsen:2000:CNJ

[PG00] Michael Philippsen and Edwin Günthner. Complex numbers for Java. Concurrency: Practice and Experience, 12(6):477–491, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72515730/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72515730&PLACEBO=IE.pdf.

Pandey:2000:PFG

[PH00a] Raju Pandey and Brant Hashii.

Providing fine-grained access control for Java programs via binary editing. Concurrency: Practice and Experience, 12 (14):1405–1430, December 10, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76500721/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500721&PLACEBO=IE.pdf.

Philippsen:2000:LOJ

[PH00b] Michael Philippsen and Bernhard Haumacher. Locality optimization in JavaParty by means of static type analysis. Concurrency: Practice and Experience, 12(8):613–628, July 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/72516226/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72516226&PLACEBO=IE.pdf.

Philippsen:2000:CDG

[Phi00a] Michael Philippsen. Cooperating distributed garbage collectors for clusters and beyond. Concurrency: Practice and Experience, 12(7):595–610, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/72516212/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72516212&PLACEBO=IE.pdf.

Philippsen:2000:SCO

[Phi00b] Michael Philippsen. A survey of concurrent object-oriented languages. Concurrency: Practice and Experience, 12(10): 917–980, August 25, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/73000573/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=73000573&PLACEBO=IE.pdf.

Philippsen:2000:MES

[PHN00] Michael Philippsen, Bernhard Haumacher, and Christian Nester. More efficient serialization and RMI for Java. Concurrency: Practice and Experience, 12(7):495–518, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516213/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516213&PLACEBO=IE.pdf.

Pirozzi:1997:NSF

[Pir97] Maria A. Pirozzi. Numerical simulation of fluid dynamic problems on distributed memory parallel computers. Concurrency: Practice and Experience, 9(10):989–998, October 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13813;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13813&PLACEBO=IE.pdf.

Prior:1990:WPR

[PNRC90] D. M. N. Prior, M. G. Norman, N. J. Radcliffe, and L. J. Clarke. What price regularity? Concurrency: Practice and Experience, 2(1):55–78, March 1990. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Perrott:1992:ECS

[PR92] R. H. Perrott and A. Ramasubbu. An experiment in concurrent software evaluation. Concurrency: Practice and Experience, 4(7):533–555, October 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Pellegrini:2000:HND

[PRA00] François Pellegrini, Jean Roman, and Patrick Amestoy. Hybridizing nested dissection and halo approximate minimum degree for efficient sparse matrix ordering. Concurrency: Practice and Experience, 12(2-3):69–84, February– March 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/72504939/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72504939&PLACEBO=IE.pdf.

Pritchard:1991:E

[Pri91] D. Pritchard. Editorial. Concurrency: Practice and Experience, 3(4):247–??, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Pond:1996:DSM

[PS96] B. Pond and I. Sharf. Dynamics simulation of multibody chains on a transputer system. Concurrency: Practice and Experience, 8(3):235–249, April 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23274.

Pauw:2000:VRP

[PS00a] Wim De Pauw and Gary Sevitsky. Visualizing reference patterns for solving memory leaks in Java. Concurrency: Practice and Experience, 12 (14):1431–1454, December 10, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76500722/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500722&PLACEBO=IE.pdf.

Protopopov:2000:SMC

[PS00b] Boris V. Protopopov and Anthony Skjellum. Sharedmemory communication approaches for an MPI messagepassing library. Concurrency: Practice and Experience, 12(9):799–820, August 10, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72516482/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516482&PLACEBO=IE.pdf.

Pedroso:1997:WBM

[PSS97] Hernâni Pedroso, Luis M. Silva, and João Gabriel Silva. Webbased metacomputing with JET. Concurrency: Practice and Experience, 9(11): 1169–1173, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13834;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13834&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Pugh:2000:JMM

[Pug00] William Pugh. The Java memory model is fatally flawed. Concurrency: Practice and Experience, 12(6):445–455, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/72515725/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72515725&PLACEBO=IE.pdf.

Philippsen:1997:JTR

[PZ97] Michael Philippsen and Matthias Zenger. JavaParty — transparent remote objects in Java. Concurrency: Practice and Experience, 9(11): 1225–1242, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13816;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13816&PLACEBO=IE.pdf;
                                        http://wwwipd.ira.uka.de/~
                                        phlipp/partyd.ps.gz. Special

Issue: Java for computational science and engineering — simulation and modeling II.

Raje:1998:CCE

[RBF98] Rajeev R. Raje, Michael Boyles, and Shiaofen Fang. CEV: collaborative environment for visualization using Java RMI. Concurrency: Practice and Experience, 10 (11–13):1079–1085, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050406;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050406&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Robertson:1991:MAT

[RCZM91] B. Robertson, M. Chopping, K. Zielinski, and D. Milway. The Metrobridge — an application of transputers in transparent bridging. Concurrency: Practice and Experience, 3(4): 445–453, August 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Rodriguez:1999:PRM

[RDLG99] C. Rodríguez, F. De Sande, C. León, and L. García. Parallelism and recursion in message passing libraries: an efficient methodology. Concurrency: Practice and Experience, 11(7):355–365, June 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=62500298;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=62500298&PLACEBO=IE.pdf.

See correspondence [Han00,

Leó00].

Rhodes:2000:SAP

[RG00] David L. Rhodes and Apostolos Gerasoulis. A scheduling approach to parallel harmonic balance simulation. Concurrency: Practice and Experience, 12 (2-3):175–187, February–March 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/72504930/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=72504930&PLACEBO=IE.pdf.

Richards:1994:ATS

[RGSB94] A. Richards, T. Ginige, A. Seneviratne, and T. Buczkowska. Adaptive transport service for high speed networks. Concurrency: Practice and Experience, 6(4):357–373, June 1994. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Rinard:1998:AEJ

[Rin98] Martin C. Rinard. Applications experience in Jade. Concurrency: Practice and Experience, 10(6):417–448, May 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=5387;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5387&PLACEBO=IE.pdf.

Rieffel:2000:CSN

[RIST00] Marc Rieffel, Mikhail Ivanov, Sadasivan Shankar, and Stephen Taylor. Concurrent simulation of neutral flow in the GEC reference cell. Concurrency: Practice and Experience, 12(1):1–19, January 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72504926/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72504926&PLACEBO=IE.pdf.

Rabhi:1991:ETP

[RM91] F. A. Rabhi and G. A. Manson.

Experiments with a transputerbased parallel graph reduction machine. Concurrency: Practice and Experience, 3(4):413– 422, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Rygol:1991:MTM

[RPB91] M. Rygol, S. Pollard, and C. Brown. MARVIN and TINA: A multiprocessor 3-D vision system. Concurrency: Practice and Experience, 3(4): 333–356, August 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Reed:1996:PPP

[RPL96] Joy Reed, Kevin Parrott, and Tim Lanfear. Portability, predictability and performance for parallel computing: BSP in practice. Concurrency: Practice and Experience, 8(10):799–812, December 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23265.

Reeve:1996:MAS

[RR96] J. S. Reeve and M. C. Rogers. MP: an application specific concurrent language. Concurrency: Practice and Experience, 8(4):313–333, May 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23280.

Rauber:1997:LBS

[RR97] Thomas Rauber and Gudula Rünger. Load balancing schemes for extrapolation methods. Concurrency: Practice and Experience, 9(3):181–202, March 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13856;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13856&PLACEBO=IE.pdf.

Rauber:1999:PEE

[RR99] Thomas Rauber and Gudula Rünger. Parallel execution of embedded and iterated Runge– Kutta methods. Concurrency: Practice and Experience, 11(7):367–385, June 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=62500299;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=62500299&PLACEBO=IE.pdf.

Russ:1999:UHR

[RRG^{+}99] Samuel H. Russ, Jonathan Robinson, Matt Gleeson, Brad Meyers, Laxman Rajagopalan, and Chun-Heong Tan. Using Hector to run MPI programs over networked workstations. Concurrency: Practice and Experience, 11(4):189–204, April 10, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=61004080;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=61004080&PLACEBO=IE.pdf.

Special Issue: Applications of Distributed Computing Environments.

Roda:1999:PET

[RRMA99] J. L. Roda, C. Rodríguez, D. G.

Morales, and F. Almeida. Predicting the execution time of message passing models. Concurrency: Practice and Experience, 11(9):461–477, August 10, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=63001745;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=63001745&PLACEBO=IE.pdf.

Russell:1997:PCP

[RSH97] Stuart Russell, Lewis Stiller, and Othar Hansson. PN-PACK: Computing with probabilities in Java. Concurrency: Practice and Experience, 9(11):1333–1339, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13843;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13843&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Raje:1997:ARM

[RWB97] Rajeev R. Raje, Joseph I. Williams, and Michael Boyles. Asynchronous Remote Method Invocation (ARMI) mechanism for Java. Concurrency: Practice and Experience, 9(11): 1207–1211, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13835;
 http://www3.interscience.

wiley.com/cgi-bin/fulltext?
ID=13835&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Selwood:1999:PUT

[SB99] P. M. Selwood and M. Berzins.

Parallel unstructured tetrahedral mesh adaptation: algorithms, implementation and scalability. Concurrency: Practice and Experience, 11(14): 863–884, December 10, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/67501484/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=67501484&PLACEBO=IE.pdf.

Schwan:1989:GDC

[SBBG89] K. Schwan, B. Blake, W. Bo, and J. Gawkowski. Global data and control in multicomputers: Operating system primitives and experimentation with a parallel branch-and-bound algorithm. Concurrency: Practice and Experience, 1(2):191– 218, December 1989. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Sathye:1996:EUH

[SBD^{+}96] A. Sathye, G. Bassett, K. Droegemeier, M. Xue, and K. Brewster. Experiences using high performance computing for operational storm scale weather prediction. Concurrency: Practice and Experience, 8(10):731–740, December 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23261.

Shao:1995:PRQ

[SBH95] J. Shao, D. A. Bell, and M. E. C. Hull. Processing recursive queries on transputers. Concurrency: Practice and Experience, 7(2):81–120, April 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Schmidt:1993:ADA

[SBS93] D. C. Schmidt, D. F. Box, and T. Suda. ADAPTIVE — A Dynamically Assembled Protocol Transformation, Integration and eValuation Environment. Concurrency: Practice and Experience, 5(4):269–286, June 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Saltz:1991:MRC

[SBW91] J. Saltz, H. Berryman, and J. Wu. Multiprocessors and run-time compilation. Concurrency: Practice and Experience, 3(6):573–592, December 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Schroeder:1999:SEA

[Sch99] Wayne Schroeder. The SDSC encryption/authentication (SEA) system. Concurrency: Practice and Experience, 11(15): 913–931, December 25, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/71005732/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=71005732&PLACEBO=IE.pdf.

Song:1997:ALL

[SCL97] Jianjian Song, Heng Kek Choo, and Kuok Ming Lee. Application-level load migration and its implementation on top of PVM. Concurrency: Practice and Experience, 9(1):1–19, January 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13803;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13803&PLACEBO=IE.pdf.

Shirazi:1995:CST

[SCM95] B. Shirazi, H.-B. Chen, and J. Marquis. Comparative study of task duplication static scheduling versus clustering and non-clustering techniques. Concurrency: Practice and Experience, 7(5):371–389, August 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Stephen:1991:VTS

[SDT91] G. A. Stephen, E. L. Dagless, and C. A. Taylor. A visual tracking system for the measurement of dynamic structural displacements. Concurrency: Practice and Experience, 3(4): 357–366, August 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Stein:1993:DAO

[SF93] J. Stein and G. C. Fox. Dependence analysis for outer loop parallelization of existing Fortran-77 programs. Concurrency: Practice and Experience, 5(8):659–674, December 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Smith:1997:MML

[SGB97] Taber H. Smith, Aaron E. Gower, and Duane S. Boning. A matrix math library for Java. Concurrency: Practice and Experience, 9(11):1127–1137, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13821;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13821&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering — simulation and modeling II.

Souder:1998:JCB

[SHGD98] Dan Souder, Morgan Herrington, Rajat P. Garg, and Dennis Deryke. JSPICE: a componentbased distributed Java frontend for SPICE. Concurrency: Practice and Experience, 10(11–13):1131–1141, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050401;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050401&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Shah:2000:FCS

[SHPT00] Sanjiv Shah, Grant Haab, Paul Petersen, and Joe Throop. Flexible control structures for parallelism in OpenMP. Concurrency: Practice and Experience, 12(12):1219–1239, October 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/76500348/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76500348&PLACEBO=IE.pdf.

Sarmenta:1998:TBB

[SHW98] Luis F. G. Sarmenta, Satoshi Hirano, and Stephen A. Ward. Towards Bayanihan: building an extensible framework for volunteer computing using Java. Concurrency: Practice and Experience, 10(11–13): 1015–1019, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050419;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050419&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Smith:2000:DPM

[SK00] Lorna Smith and Paul Kent. Development and performance of a mixed OpenMP/MPI quantum Monte Carlo code. Concurrency: Practice and Experience, 12(12):1121–1129, October 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract/76500350/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76500350&PLACEBO=IE.pdf.

Sturtevant:1992:PPP

[SMC92] J. E. Sturtevant, A. B. Maccabe, and P. M. Campbell. Performance of a particle-in-cell plasma simulation code on the BBN TC2000. Concurrency: Practice and Experience, 4(1): 1–18, February 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Silva:2000:HPC

[SMS00] Luís Moura Silva, Paulo Martins, and João Gabriel Silva. Heterogeneous parallel computing using Java and WMPI. Concurrency: Practice and Experience, 12(11): 1077–1091, September 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76000189/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76000189&PLACEBO=IE.pdf.

Shields:2000:JCB

[SRW^{+}00] Matthew S. Shields, Omer F.

Rana, David W. Walker, Maozhen Li, and David Golby. A Java/CORBA-based visual program composition environment for PSEs. Concurrency: Practice and Experience, 12(8):687–704, July 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516220/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516220&PLACEBO=IE.pdf.

Schmidt:1994:EAO

[SS94] B. K. Schmidt and V. S. Sunderam. Empirical analysis of overheads in cluster environments. Concurrency: Practice and Experience, 6(1):1– 32, February 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Shivaratri:1995:LIT

[SS95] N. G. Shivaratri and M. Singhal. A load index and a transfer policy for global scheduling tasks with deadlines. Concurrency: Practice and Experience, 7(7):671–688, October 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Szafron:1996:EMU

[SS96] Duane Szafron and Jonathan Schaeffer. An experiment to measure the usability of parallel programming systems. Concurrency: Practice and Experience, 8(2):147–166, March 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=23270.

Siegell:1997:ASL

[SS97] B. S. Siegell and P. A. Steenkiste. Automatic selection of load balancing parameters using compile-time and run-time information. Concurrency: Practice and Experience, 9(4):275–317, April 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13862;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13862&PLACEBO=IE.pdf.

Schwab:1998:AJC

[SS98] Matthias Schwab and Joel Schroeder. Algebraic Java classes for numerical optimization. Concurrency: Practice and Experience, 10(11–13):

1155–1164, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050405;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050405&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Skvoretz:1992:PAD

[SSB92] J. Skvoretz, S. A. Smith, and C. Baldwin. Parallel-processing applications for data analysis in the social sciences. Concurrency: Practice and Experience, 4(3):207–221, May 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Suzuoka:1997:PDT

[SSG97] Takashi Suzuoka, Jaspal Subhlok, and Thomas Gross. A performance debugging tool for high performance Fortran programs. Concurrency: Practice and Experience, 9(10):927–945, October 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13809;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13809&PLACEBO=IE.pdf.

Singh:1998:EPP

[SSS98] Ajit Singh, Jonathan Schaeffer, and Duane Szafron. Experience with parallel programming using code templates. Concurrency: Practice and Experience, 10(2):91–120, February 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=5368;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5368&PLACEBO=IE.pdf.

Scott:1989:TPO

[ST89] R. B. Scott and R. Trehan. Translating from PARLOG to occam2: A methodology. Concurrency: Practice and Experience, 1(1):105–134, September 1989. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Shadid:1992:SIA

[ST92] J. N. Shadid and R. S. Tuminaro. Sparse iterative algorithm software for large-scale MIMD machines: an initial discussion and implementation. Concurrency: Practice and Experience, 4(6):481–497, September 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Stankovic:2000:OJS

[Sta00] Nenad Stankovic. An open Java system for SPMD programming. Concurrency: Practice and Experience, 12(11): 1051–1076, September 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76000192/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76000192&PLACEBO=IE.pdf.

Succi:1996:TCF

[SU96] Giancarlo Succi and Carl Uhrik.

Towards a complete framework for parallel implementation of logic languages: the data parallel implementation of SEL. Concurrency: Practice and Experience, 8(3):191–204, April 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23273.

Sunderam:1990:PFP

[Sun90] V. S. Sunderam. PVM: A framework for parallel distributed computing. Concurrency: Practice and Experience, 2(4):315–339, December 1990. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Sipcic:1991:AAC

[SW91] S. R. Sipcic and T. Westbrook.

Aeroelastic applications of the Connection Machine. Concurrency: Practice and Experience, 3(6):715–724, December 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Tarnvik:1994:DPT

[Tär94] E. Tärnvik. Dynamo — a portable tool for dynamic load balancing on distributed memory multicomputers. Concurrency: Practice and Experience, 6(8):613–639, December 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Tanenbaum:1994:OAP

[TBBK94] A. S. Tanenbaum, H. E. Bal, S. Ben Hassen, and M. Frans Kaashoek. Object-based approach to programming distributed systems. Concurrency: Practice and Experience, 6(4): 235–249, June 1994. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Thiruvathukal:2000:JNW

[TDB00] George K. Thiruvathukal, Phillip M. Dickens, and Shahzad Bhatti. Java on networks of workstations (JavaNOW): a parallel computing framework inspired by Linda and the Message Passing Interface (MPI). Concurrency: Practice and Experience, 12(11):1093–1116, September 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/76000187/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76000187&PLACEBO=IE.pdf.

Teng:1993:RVA

[TDD93] Y. Ansel Teng, Daniel Dementhon, and Larry S. Davis. Region-to-region visibility analysis using data parallel machines. Concurrency: Practice and Experience, 5(5):379– 406, August 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Toomarian:1994:TSL

[TFB94] N. Toomarian, A. Fijany, and J. Barhen. Time-parallel solution of linear partial differential equations on the Intel Touchstone Delta supercomputer. Concurrency: Practice and Experience, 6(8):641– 652, December 1994. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Theobald:1996:WNL

[The96] Kevin B. Theobald. The Wnetwork: A low-cost faulttolerant multistage interconnection network for fine-grain multiprocessing. Concurrency: Practice and Experience, 8(6):415–428, July 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=23291.

Tripathi:1994:CMN

[TKKN94] A. Tripathi, N. M. Karnik, S. P. Koneru, and C. Nock. Configuration management in the Nexus distributed operating system. Concurrency: Practice and Experience, 6(4):325–338, June 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Takagi:1998:NMP

[TMN^{+}98] Hiromitsu Takagi, Satoshi Matsuoka, Hidemoto Nakada, Satoshi Sekiguchi, Mitsuhisa Satoh, and Umpei Nagashima. Ninflet: a migratable parallel objects framework using Java. Concurrency: Practice and Experience, 10(11–13): 1063–1078, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050423;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050423&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Togneri:1997:PPA

[Tog97]  R. Togneri.  Parallel pro-

gram analysis on workstation clusters: speedup profiling and latency hiding. Concurrency: Practice and Experience, 9(7):721–751, July 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13886;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13886&PLACEBO=IE.pdf.

Togneri:1998:PPA

[Tog98] Roberto Togneri. Parallel program analysis on workstation clusters: Memory utilisation and load balancing. Concurrency: Practice and Experience, 10(8):631–653, July 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10008697;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10008697&PLACEBO=IE.pdf.

Tuminaro:1998:PSM

[TSH98] Ray S. Tuminaro, John N. Shadid, and Scott A. Hutchinson. Parallel sparse matrix vector multiply software for matrices with data locality. Concurrency: Practice and Experience, 10(3):229–247, March 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=5375;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5375&PLACEBO=IE.pdf.

Topol:1998:PTV

[TSS98] Brad Topol, John T. Stasko, and Vaidy Sunderam. PVaniM: a tool for visualization in network computing environments. Concurrency: Practice and Experience, 10(14): 1197–1222, December 10, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=40005932;
 http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?
                                        ID=40005932&PLACEBO=IE.pdf.

Tollenaere:1994:MMN

[TSV94] T. Tollenaere, J. M. Saraiva, and M. M. Van Hulle. Massive MIMD neural network simulations: the connections dilemma. Concurrency: Practice and Experience, 6(3):153–191, May 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Thiruvathukal:1998:RRM

[TTK98] G. K. Thiruvathukal, L. S. Thomas, and A. T. Korczynski. Reflective remote method invocation. Concurrency: Practice and Experience, 10(11– 13):911–925, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050398;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050398&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Tahir:1990:RDA

[TV90] J. M. Tahir and G. S. Virk. A real-time distributed algorithm for an aircraft longitudinal optimal autopilot. Concurrency: Practice and Experience, 2(2):109–122, June 1990. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Urban:2000:CSM

[UTS00] Susan D. Urban, Michael Tjahjadi, and Jami J. Shah. A case study in mapping conceptual designs to objectrelational schemas. Concurrency: Practice and Experience, 12(9):863–907, August 10, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72516481/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516481&PLACEBO=IE.pdf.

vandeVelde:1990:EML

[van90] E. F. van de Velde. Experiments with multicomputer LU-decomposition. Concurrency: Practice and Experience, 2(1):1–26, March 1990. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Vaughan:1991:PPA

[VBFW91] J. R. Vaughan, G. R. Brookes, M. A. Fletcher, and D. P. M. Wills. Parallel processing architectures used in a real-time display application. Concurrency: Practice and Experience, 3(3):187–201, June 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

vanDongen:1996:E

[vD96] Vincent van Dongen. Editorial. Concurrency: Practice and Experience, 8(6):413, July 1996. CODEN CPEXEI.

ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23292.

vandeGeijn:1997:SSU

[vdGW97] R. A. van de Geijn and J. Watts. SUMMA: scalable universal matrix multiplication algorithm. Concurrency: Practice and Experience, 9(4):255–274, April 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13861;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13861&PLACEBO=IE.pdf.

vonLaszewski:1999:LCM

[vL99]  Gregor von Laszewski.  A

loosely coupled metacomputer: co-operating job submissions across multiple supercomputing sites. Concurrency: Practice and Experience, 11(15): 933–948, December 25, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/71005731/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=71005731&PLACEBO=IE.pdf.

Vanderwiel:1998:CAP

[VNL98] Steven P. Vanderwiel, Daphna Nathanson, and David J. Lilja. A comparative analysis of parallel programming language complexity and performance. Concurrency: Practice and Experience, 10(10):807–820, August 25, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=10006459;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10006459&PLACEBO=IE.pdf.

vanNieuwpoort:2000:WAP

[vNMB^{+}00] Rob van Nieuwpoort, Jason Maassen, Henri E. Bal, Thilo Kielmann, and Ronald Veldema. Wide-area parallel programming using the remote method invocation model. Concurrency: Practice and Experience, 12(8):643–666, July 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516223/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516223&PLACEBO=IE.pdf.

VanHalderen:1998:FWB

[VO98a] Berry A. W. Van Halderen and Benno J. Overeinder. Fornax: Web-based distributed discrete event simulation in Java. Concurrency: Practice and Experience, 10(11– 13):957–970, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050402;
 http://www3.interscience.

                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050402&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Vaughan:1998:EED

[VO98b] John G. Vaughan and Mary O’Donovan. Experimental evaluation of distributed load balancing implementations. Concurrency: Practice and Experience, 10(10):763–782, August 25, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10006460;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10006460&PLACEBO=IE.pdf.

vanRenterghem:1989:TIA

[vR89] P. van Renterghem. Transputers for industrial applications. Concurrency: Practice and Experience, 1(2):135–170, December 1989. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

vanReeuwijk:1997:SPL

[vRvGS97] Kees van Reeuwijk, Arjan J. C. van Gemund, and Henk J. Sips. Spar: A programming language for semi-automatic compilation of parallel programs. Concurrency: Practice and Experience, 9(11):1193–1205, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.

com/cgi-bin/abstract?ID=13818;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13818&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Volken:1996:PPM

[VSKN96] W. Volken, P. Schwab, P. G.

Kropf, and H. Neuenschwander. Parallelizing a powerful Monte Carlo method for electron beam dose calculations. Concurrency: Practice and Experience, 8(6):489–498, July 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23289.

Wadsworth:1999:EPS

[Wad99] Chris Wadsworth. Editorial:

Portable software tools for parallel architectures. Concurrency: Practice and Experience, 11(11):587–591, September 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=64500094;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=64500094&PLACEBO=IE.pdf.

Walker:1990:CPP

[Wal90] D. W. Walker. Characterizing the parallel performance of a large-scale, particle-in-cell plasma simulation code. Concurrency: Practice and Experience, 2(4):257–288, December 1990. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Wallcraft:2000:SOV

[Wal00] Alan J. Wallcraft. SPMD OpenMP versus MPI for ocean models. Concurrency: Practice and Experience, 12 (12):1155–1164, October 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract/76500353/
                                        START;   http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=76500353&PLACEBO=IE.pdf.

Walshaw:1995:DLP

[WB95a] C. Walshaw and M. Berzins. Dynamic load-balancing for PDE solvers on adaptive unstructured meshes. Concurrency: Practice and Experience, 7(1):17–28, February 1995. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Wang:1995:PHM

[WB95b] H. Wang and M. Brady. PARA-DOX — a heterogeneous machine for the 3D vision algorithm. Concurrency: Practice and Experience, 7(4):257–272, June 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Welsh:2000:JEE

[WC00] Matt Welsh and David Culler. Jaguar: enabling efficient communication and I/O in Java. Concurrency: Practice and Experience, 12(7):519–538, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516211/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516211&PLACEBO=IE.pdf.

Weston:1991:PCE

[WCB91] J. S. Weston, M. Clint, and C. W. Bleakney. The parallel computation of eigenvalues and eigenvectors of large Hermitian matrices using the AMT DAP

510. Concurrency: Practice and Experience, 3(3):179–185, June 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Wu:1999:CFF

[WCC99] Jan-Jan Wu, Marina Chen, and James Cowie. CRAFT: a framework for F90/HPF compiler optimizations. Concurrency: Practice and Experience, 11(10):529–569, August 25, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=63003393;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=63003393&PLACEBO=IE.pdf.

Warner:1997:JSP

[WCL97] Simeon Warner, Simon Catterall, and Edward Lipson. Java simulations for physics education. Concurrency: Practice and Experience, 9(6):477–484, June 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=13872;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=13872&PLACEBO=IE.pdf.

Worley:1992:PST

[WD92] P. H. Worley and J. B. Drake. Parallelizing the spectral transform method. Concurrency: Practice and Experience, 4(4): 269–291, June 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Williams:1999:ATP

[WDT^{+}99] M. H. Williams, E. W. Dempster, N. T. Tomov, C. S. Pua, H. Taylor, A. Burger, J. Lü, and P. Broughton. An analytical tool for predicting the performance of parallel relational databases. Concurrency: Practice and Experience, 11(11):635–653, September 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=64500091;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=64500091&PLACEBO=IE.pdf.

Wang:1991:IAT

[WDW91] H. Wang, P. M. Dew, and J. A. Webb. Implementation of apply on a transputer array. Concurrency: Practice and Experience, 3(1):43–54, February 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Weissman:1999:PAS

[Wei99] Jon B. Weissman. Prophet: automated scheduling of SPMD programs in workstation networks. Concurrency: Practice and Experience, 11(6):301–321, May 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=61004096;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=61004096&PLACEBO=IE.pdf.

Wang:1997:PCN

[WF97] P. Wang and R. D. Ferraro. Parallel computation for natural convection. Concurrency: Practice and Experience, 9(10):975–987, October 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13812;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13812&PLACEBO=IE.pdf.

Weissman:1995:FPP

[WG95] J. B. Weissman and A. S.

Grimshaw. A framework for partitioning parallel computations in heterogeneous environments. Concurrency: Practice and Experience, 7(5):455– 478, August 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Wilson:1993:ESD

[WGC93] L. F. Wilson, M. J. Gonzalez, and M. W. Cruess. Experiences with SLALOM on the distributed Pleiades/ESP system. Concurrency: Practice and Experience, 5(4):329–344, June 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Wheat:1990:PSA

[Whe90] M. Wheat. A parallel sorting algorithm for tightly coupled multiprocessors. Concurrency: Practice and Experience, 2(1):27–32, March 1990. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Wheat:1991:SNL

[Whe91] M. Wheat. Sorting with near linear speed-up on tightly coupled multiprocessors. Concurrency: Practice and Experience, 3(1):1–14, February 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Wieland:1989:PDC

[WHF^{+}89] F. Wieland, L. Hayley, A. Feinberg, M. DiLoreto, L. Blume, J. Ruffles, P. Reiher, B. Beckman, P. Hontalas, S. Bellenot, and D. Jefferson. The performance of a distributed combat simulation with the Time Warp operating system. Concurrency: Practice and Experience, 1(1):

35–50, September 1989. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Williams:1989:SFF

[Wil89] R. D. Williams. Supersonic fluid flow in parallel with an unstructured mesh. Concurrency: Practice and Experience, 1(1): 51–62, September 1989. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Williams:1991:PDL

[Wil91] R. D. Williams. Performance of dynamic load balancing algorithms for unstructured mesh calculations. Concurrency: Practice and Experience, 3(5): 457–481, October 1991. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Williams:1992:VDP

[Wil92]  Roy D. Williams.  Voxel

databases: A paradigm for parallelism with spatial structure. Concurrency: Practice and Experience, 4(8):619–636, December 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Wagner:1993:TPM

[WJC93] A. Wagner, J. Jiang, and S. Chanson. Tmon: A performance monitor for transputerbased multicomputers. Concurrency: Practice and Experience, 5(6):511–526, September 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Weerawarana:1997:NIN

[WJH^{+}97] Sanjiva Weerawarana, Anupam Joshi, Elias N. Houstis, John R. Rice, and Ann C. Catlin. Notebook interfaces for networked scientific computing: design and WWW implementation. Concurrency: Practice and Experience, 9(7):675–695, July 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13884;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13884&PLACEBO=IE.pdf.

Weber:1998:PCA

[WKE98] Andreas Weber, Wolfgang Küchlin, and Bernhard Eggers. Parallel computer algebra software as a Web component. Concurrency: Practice and Experience, 10(11–13): 1179–1188, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050416;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050416&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Walker:2000:SAD

[WLR^{+}00] D. W. Walker, M. Li, O. F. Rana, M. S. Shields, and Y. Huang. The software architecture of a distributed problem-solving environment. Concurrency: Practice and Experience, 12(15):1455–1480, December 25, 2000. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (elec-

tronic). URL http://www3.
interscience.wiley.com/cgi-
bin/abstract/76505645/START;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=76505645&PLACEBO=IE.pdf.

Wang:1998:DSP

[WMC98] Kuei Yu Wang, Dan C. Marinescu, and Octavian F. Carbunar. Dynamic scheduling of process groups. Concurrency: Practice and Experience, 10(4):265–283, April 10, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=5378;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5378&PLACEBO=IE.pdf.

Walker:1996:RBC

[WO96] D. W. Walker and S. W. Otto.

Redistribution of block-cyclic data distributions using MPI. Concurrency: Practice and Experience, 8(9):707–728, November 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=23305.

Wolfe:1993:EDD

[Wol93] M. Wolfe. Engineering a data dependence test. Concurrency: Practice and Experience, 5(7): 603–622, October 1993. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Wood:1991:DGD

[Woo91] K. R. Wood. Distributing gate-level digital timing simulation over arrays of transputers. Concurrency: Practice and Experience, 3(4):367–379, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Worley:1998:SAS

[WRMB98] Patrick H. Worley, Allen C. Robinson, David R. Mackay, and Edward J. Barragy. A study of application sensitivity to variation in messagepassing latency and bandwidth. Concurrency: Practice and Experience, 10(5):387–406, April 25, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=5384;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5384&PLACEBO=IE.pdf.

Wu:1995:SHS

[Wu95] M.-Y. Wu. Symmetrical Hopping: a scalable scheduling algorithm for irregular problems. Concurrency: Practice and Experience, 7(7):689–706, October 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Wu:1997:DEB

[Wu97] J. Wu. Dimension-exchange-based global load balancing on injured hypercubes. Concurrency: Practice and Experience, 9(1):41–61, January 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13805;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13805&PLACEBO=IE.pdf.

Walker:1992:PST

[WWD92] D. W. Walker, P. H. Worley, and J. B. Drake. Parallelizing the spectral transform method. part II. Concurrency: Practice and Experience, 4(7): 509–531, October 1992. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Xu:1997:DRD

[XLD97] Chengzhong Xu, Francis C. M.

Lau, and Ralf Diekmann. Decentralized remapping of data parallel applications in distributed memory multiprocessors. Concurrency: Practice and Experience, 9(12): 1351–1376, December 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=13847;
http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13847&PLACEBO=IE.pdf.

Xu:1995:NAL

[XMLL95] C. Xu, B. Monien, R. Lüling, and F. C. M. Lau. Nearestneighbor algorithms for loadbalancing in parallel computers. Concurrency: Practice and Experience, 7(7):707–736, October 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Xu:2000:MAB

[XW00] Cheng-Zhong Xu and Brian Wims. A mobile agent based push methodology for global parallel computing. Concurrency: Practice and Experience, 12(8):705–726, July 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72516227/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72516227&PLACEBO=IE.pdf.

Yang:2000:SIS

[Yan00] Tao Yang. Special issue: Solving irregularly structured problems. Concurrency: Practice and Experience, 12(2-3): i–ii, February–March 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract/72504936/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72504936&PLACEBO=IE.pdf.

Yu:1997:JDP

[YC97] Weimin Yu and Alan Cox.

Java/DSM: A platform for heterogeneous computing. Concurrency: Practice and Experience, 9(11):1213–1224, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=13817;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=13817&PLACEBO=IE.pdf. Spe-

cial Issue: Java for computational science and engineering

— simulation and modeling II.

Yalamanchilli:1998:CPJ

[YC98] Narendar Yalamanchilli and William Cohen. Communication performance of Javabased parallel virtual machines. Concurrency: Practice and Experience, 10(11–13): 1189–1196, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=10050407;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=10050407&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Yang:1995:FTE

[YQ95] C.-Q. Yang and Y. Qu. Fault tolerance in the execution of remote jobs on idling workstations. Concurrency: Practice and Experience, 7(1):43– 60, February 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Yelick:1998:THP

[YSP^{+}98] Kathy Yelick, Luigi Semenzato, Geoff Pike, Carleton Miyamoto, Ben Liblit, Arvind Krishnamurthy, Paul Hilfinger, Susan Graham, David Gay, Phil Colella, and Alex Aiken. Titanium: a high-performance Java dialect. Concurrency: Practice and Experience, 10 (11–13):825–836, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

                                        /www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050392;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050392&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Yalamanchili:1995:PME

[YTPS95] S. Yalamanchili, L. Te Winkel, D. Perschbacher, and B. Shenoy. Partitioning and mapping in embedded multiprocessor architectures in the presence of constraints. Concurrency: Practice and Experience, 7(3):167–189, May 1995. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Yutaka:2000:EJV

[YTY00] Oiwa Yutaka, Kenjiro Taura, and Akinori Yonezawa. Extending Java virtual machine with integer-reference conversion. Concurrency: Practice and Experience, 12(6):407–422, May 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract/72515729/
START;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=72515729&PLACEBO=IE.pdf.

Zhou:1998:LST

[ZG98] Honbo Zhou and Al Geist.

LPVM: a step towards multithread PVM. Concurrency: Practice and Experience, 10 (5):407–416, April 25, 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=5385;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=5385&PLACEBO=IE.pdf.

Zoraja:1999:SPD

[ZHS99] Ivan Zoraja, Hermann Hellwagner, and Vaidy Sunderam. SCIPVM: Parallel distributed computing on SCI workstation clusters. Concurrency: Practice and Experience, 11(3):121–138, March 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

//www3.interscience.wiley.
com/cgi-bin/abstract?ID=61003667;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=61003667&PLACEBO=IE.pdf.

Zimmermann:1991:VSM

[Zim91] P. Zimmermann. Visual simulation by means of a transputer network for a driving simulator. Concurrency: Practice and Experience, 3(4):399–412, August 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Ziavras:1999:ECC

[ZK99] Sotirios G. Ziavras and Sanjay Krishnamurthy. Evaluating the communications capabilities of the generalized hypercube interconnection network. Concurrency: Practice and Experience, 11(6):281–300, May 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=61004097;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=61004097&PLACEBO=IE.pdf.

Zomaya:1992:TAO

[ZM92] A. Y. H. Zomaya and A. S. Morris. Transputer arrays for the on-line computation of robot Jacobians. Concurrency: Practice and Experience, 4(5):399– 412, August 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Zomaya:1992:HET

[Zom92] A. Y. Zomaya. Highly efficient transputer arrays for the computation of robot dynamics. Concurrency: Practice and Experience, 4(2):185–205, April 1992. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Zhu:1999:PSA

[ZP99] Wenjie Zhu and Linda Petzold. Parallel sensitivity analysis for DAEs with many parameters. Concurrency: Practice and Experience, 11(10): 571–585, August 25, 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:/

/www3.interscience.wiley.
com/cgi-bin/abstract?ID=63003392;
 http://www3.interscience.
wiley.com/cgi-bin/fulltext?
ID=63003392&PLACEBO=IE.pdf.

Ziavras:1993:PMH

[ZS93] S. G. Ziavras and M. A. Siddiqui. Pyramid mappings onto hypercubes for computer vision: Connection Machine comparative study. Concurrency: Practice and Experience, 5(6):471– 489, September 1993. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Ziavras:1994:HEH

[ZS94] S. G. Ziavras and D. P. Shah.

High-performance emulation of hierarchical structures on hypercube supercomputers. Concurrency: Practice and Experience, 6(2):85–100, April 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Zhu:1998:SUP

[ZTLV98] Jinsong Zhu, Maria Törö, Victor Leung, and Son Vuong. Supporting universal personal computing on the Internet with Java and CORBA. Concurrency: Practice and Experience, 10(11–13):1007–1013, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http:

                                        //www3.interscience.wiley.
                                        com/cgi-bin/abstract?ID=10050418;
                                         http://www3.interscience.
                                        wiley.com/cgi-bin/fulltext?
                                        ID=10050418&PLACEBO=IE.pdf.

Special Issue: Java for Highperformance Network Computing.

Zhang:1993:MVM

[ZW93] X. Zhang and B. Wang. Monitoring and visualizing multiprocessor interconnection network contention. Concurrency: Practice and Experience, 5(8): 675–691, December 1993. CO-DEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).